### v1

# Bandit and Delayed Feedback in Online Structured Prediction

Abstract
Online structured prediction is a task of sequentially predicting outputs with complex structures based on inputs and past observations, encompassing online classification. Recent studies showed that in the full information setup, we can achieve finite bounds on the surrogate regret, i.e., the extra target loss relative to the best possible surrogate loss. In practice, however, full information feedback is often unrealistic as it requires immediate access to the whole structure of complex outputs. Motivated by this, we propose algorithms that work with less demanding feedback, bandit and delayed feedback. For the bandit setting, using a standard inverse-weighted gradient estimator, we achieve a surrogate regret bound of $O(\sqrt{KT})$ for the time horizon $T$ and the size of the output set $K$. However, $K$ can be extremely large when outputs are highly complex, making this result less desirable. To address this, we propose an algorithm that achieves a surrogate regret bound of $O(T^{2/3})$, which is independent of $K$. This is enabled with a carefully designed pseudo-inverse matrix estimator. Furthermore, for the delayed full information feedback setup, we obtain a surrogate regret bound of $O(D^{2/3}T^{1/3})$ for the delay time $D$. We also provide algorithms for the delayed bandit feedback setup. Finally, we numerically evaluate the performance of the proposed algorithms in online classification with bandit feedback.
1 Introduction
In many machine learning problems, given an input vector from a vector space $\mathcal{X}$, we aim to predict a vector in a finite output space $\mathcal{Y}$. Multiclass classification is one of the simplest examples, while in other cases output spaces may have more complex structures. Structured prediction refers to such a class of problems with structured output spaces, including multiclass classification, multilabel classification, ranking, and ordinal regression, and it has applications in various fields, ranging from natural language processing to bioinformatics (Tsochantaridis et al., 2005; BakIr et al., 2007). In structured prediction, training models that directly predict outputs in complex discrete output spaces is typically challenging. Therefore, we often adopt the surrogate loss framework (Bartlett et al., 2006)—define an intermediate space of score vectors and train models that estimate score vectors from inputs based on surrogate loss functions. Examples of surrogate losses include squared, logistic, and hinge losses, and a general framework encompassing them is the Fenchel–Young loss (Blondel et al., 2020), which we rely on in this study.
Structured prediction can be naturally extended to the online setting (Sakaue et al., 2024). In this online setting, at each round $t=1,\dots,T$, the environment selects an input-output pair $(\bm{x}_{t},\bm{y}_{t})\in\mathcal{X}\times\mathcal{Y}$. A learner then predicts $\widehat{\bm{y}}_{t}\in\mathcal{Y}$ based on the input $\bm{x}_{t}$ and incurs a loss $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$, where $L:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is the target loss function. As with Sakaue et al. (2024), we focus on the simple yet fundamental case where the learner’s model for estimating score vectors is linear.
Table 1: Surrogate regret upper and lower bounds in online multiclass classification and online structured prediction. Here, $T$ is the time horizon, $K$ is the size of the set, $\mathcal{Y}$, of output vectors, and $D$ is the fixed delayed time. “FI” is the abbreviation of full information. Delayed feedback is considered only when “Delayed” appears in the feedback column. In the target loss column, “SELF*” means SELF that satisfies Assumption 3.7. Note that the $O(T^{2/3})$ bounds for SELF* in lines 6 and 9 do not explicitly depend on $K$ but on $d$; in the case of multiclass classification with the 0-1 loss, the dependence on $K$ appears as $d=K$.
Problem setup Feedback Target loss Surrogate regret bound
Van der Hoeven et al. (2021, Cor. 1) Binary classification Bandit 0-1 loss $\Omega(\sqrt{T})$ ($d=2$)
Van der Hoeven (2020, Thm. 4) Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
Van der Hoeven et al. (2021, Thm. 1) Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
Sakaue et al. (2024, Thms. 7 and 8) Structured prediction FI SELF $O(1)$
This work (Theorems 3.5 and 3.6) Structured prediction Bandit SELF $O(\sqrt{KT})$
This work (Theorem 3.8) Structured prediction Bandit SELF* $O(T^{2/3})$
This work (Theorems 4.3 and 4.4) Structured prediction FI & Delayed SELF $O(D^{2/3}T^{1/3})$
This work (Theorem 5.1) Structured prediction Bandit & Delayed SELF $O(\sqrt{DKT})$
This work (Theorem 5.2) Structured prediction Bandit & Delayed SELF* $O(D^{1/3}T^{2/3})$
The goal of the learner is to minimize the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$. On the other hand, the best the learner can do in the surrogate loss framework is to minimize the cumulative surrogate loss, namely $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$, where $\bm{U}:\mathcal{X}\to\mathbb{R}^{d}$ is the best offline linear estimator and $S:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a surrogate loss, which measures the discrepancy between the score vector $\bm{U}\bm{x}_{t}\in\mathbb{R}^{d}$ and $\bm{y}_{t}\in\mathcal{Y}$. Given this, a natural performance measure of the learner’s predictions is the following surrogate regret $\mathcal{R}_{T}$:
$\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}.$ (1)
The surrogate regret was introduced in the seminal paper by Van der Hoeven (2020) in the context of online multiclass classification. Recently, Sakaue et al. (2024) showed that a finite surrogate regret bound can be achieved for online structured prediction under full information feedback, i.e., the learner can observe $\bm{y}_{t}$ at the end of each round $t$.
However, the assumption that full information feedback is available is often demanding, especially when outputs have complex structures. For example, in sequential ad assortment on an advertising platform, we may be able to observe only the click-through rate but not which ads were clicked, which boils down to the bandit feedback setting (Kakade et al., 2008; Gentile & Orabona, 2014a). Also, we may only have access to feedback from a while ago when designing an ad assortment for a new user—namely, delayed feedback (Weinberger & Ordentlich, 2002; Manwani & Agarwal, 2023). Similar situations have led to a plethora of studies in various online learning settings. In combinatorial bandits, algorithms under bandit feedback (referred to as full-bandit feedback in their context), instead of full information feedback, have been widely studied (Cesa-Bianchi & Lugosi, 2012; Combes et al., 2015; Rejwan & Mansour, 2020; Du et al., 2021). Delayed feedback is also explored in various settings, including full information and bandit feedback (Joulani et al., 2013; Cesa-Bianchi et al., 2016). Due to space limitations, we defer a further discussion of the background to Appendix B.
Our Contributions
To extend the applicability of online structured prediction, this study develops online structured prediction algorithms that can handle weaker feedback—bandit feedback and delayed feedback—instead of full information feedback. As with Sakaue et al. (2024), we consider the case where target loss functions belong to a class called the Structured Encoding Loss Function (SELF) (Ciliberto et al., 2016; Blondel, 2019), a general class including the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking (see Section 2.4 for the definition). Table 1 summarizes the surrogate regret bounds provided in this study and comparisons with existing results.
One of the challenges of bandit feedback is that the true output $\bm{y}_{t}$ is not observed, making it impossible to compute the true gradient of the surrogate loss. To deal with this, we use an inverse-weighted gradient estimator, a typical approach that assigns weights to gradients by the inverse of choosing each output, establishing an $O(\sqrt{KT})$ surrogate regret upper bounds (Theorems 3.5 and 3.6), where $K=\lvert\mathcal{Y}\rvert$ is the cardinality of $\mathcal{Y}$. The $O(\sqrt{KT})$ bound has an optimal dependence on $T$; it matches the $\sqrt{T}$ lower bound in the special case of online multiclass classification with bandit feedback (Van der Hoeven et al., 2021, Corollary 1). Furthermore, our bound is better than the existing $O(K\sqrt{T})$ bound of Van der Hoeven (2020) by a factor of $\sqrt{K}$, while it is not directly comparable to the latest $O(K\sqrt{T})$ bound in Van der Hoeven et al. (2021) due to differences in surrogate loss functions. See Appendix C for a more detailed discussion.
While the $O(\sqrt{KT})$ bound is satisfactory when $K=\lvert\mathcal{Y}\rvert$ is small, $K$ can be extremely large in some structured prediction problems: in multilabel classification with $m$ correct labels, we have $K=\binom{d}{m}$, and in ranking problems with $m$ items, we have $K=m!$. To address this issue, we consider a special case of SELF (denoted by SELF* in Table 1), which still includes the aforementioned examples: the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking. A technical challenge to resolve the issue lies in designing an appropriate gradient estimator used in online learning methods. To this end, we draw inspiration from pseudo-inverse estimators used in the adversarial linear/combinatorial bandit literature (Dani et al., 2007; Abernethy et al., 2008; Cesa-Bianchi & Lugosi, 2012). While we cannot naively apply the existing estimators, we design a new gradient estimator that applies to various specific structured prediction problems belonging to the special SELF framework. Armed with this gradient estimator, we achieve a surrogate regret upper bound of $O(T^{2/3})$, which does not explicitly depend on $K$ (Theorem 3.8).
For the delayed feedback setting with a known fixed delay time of $D$, it is actually not difficult to obtain a surrogate regret bound of $O(\sqrt{DT})$ with standard Online Convex Optimization (OCO) algorithms for delayed feedback. Our finding is that we can achieve a surrogate regret bound of $O(D^{2/3}T^{1/3})$ in online structured prediction under delayed full information feedback (Theorems 4.3 and 4.4) by leveraging ODAFTRL (Flaspohler et al., 2021), a Follow-the-Regularized-Leader-type algorithm that achieves an AdaGrad-type regret upper bound in OCO under delayed feedback. This bound is better than $O(\sqrt{DT})$ as $D\leq T$.
Given the contributions so far, it is natural to explore online structured prediction in environments where both delay and bandit feedback are present. We obtain algorithms for this setup by combining the theoretical developments for bandit feedback without delay and delayed full information feedback, offering surrogate regret bounds of $O(\sqrt{DKT})$ (Theorem 5.1) and $O(D^{1/3}T^{2/3})$ (Theorem 5.2).
We validate our algorithms through numerical experiments using both synthetic and real-world data. Specifically, we consider online multiclass classification with bandit feedback. We observe that, depending on the number of classes and the dataset, our algorithm, designed for general structured prediction, can achieve accuracy comparable to existing algorithms specialized for multiclass classification.
2 Preliminaries
We describe the detailed setup of online structured prediction and key tools used in this work: the Fenchel–Young loss, SELF, and randomized decoding.
Notation
[…]
2.1 Online Structured Prediction
Here, we describe the problem setting of online structured prediction. Let $\mathcal{X}$ be the input vector space and $\mathcal{Y}$ be the output vector space. Define $K\coloneqq|\mathcal{Y}|$. Following Blondel et al. (2020) and Sakaue et al. (2024), we assume that $\mathcal{Y}$ is embedded into $\mathbb{R}^{d}$ in a standard manner. For example, $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$ in $d$-class multiclass classification.
A linear estimator $\bm{W}\in\mathcal{W}$ for a convex domain $\mathcal{W}$ is used to transform the input vector $\bm{x}$ into the score vector $\bm{W}\bm{x}$. In online structured prediction, at each round $t=1,\dots,T$:
  1. 1.
    The environment selects an input $\bm{x}_{t}\in\mathcal{X}$ and the true output $\bm{y}_{t}\in\mathcal{Y}$;
  2. 2.
    The learner receives $\bm{x}_{t}$ and computes the score vector $\bm{\theta}_{t}=\bm{W}_{t}\bm{x}_{t}$ using the linear estimator $\bm{W}_{t}$;
  3. 3.
    The learner selects a predicted output $\widehat{\bm{y}}_{t}$ based on $\bm{\theta}_{t}$ and incurs a loss of $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$;
  4. 4.
    The learner receives feedback based on the problem setup and updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ using an online learning algorithm, $\mathsf{ALG}$.
The goal of the learner is to minimize the cumulative prediction loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$, which is equivalent to minimizing the surrogate regret $\mathcal{R}_{T}$ in (1). We assume that the input and output are generated in an oblivious manner. Note that when $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\dots,\bm{\mathrm{e}}_{d}\}$ and $L(\widehat{\bm{y}}_{t};\bm{y}_{t})=\mathds{1}[\widehat{\bm{y}}_{t}\neq\bm{y}_{t}]$, the above setting reduces to online multiclass classification, which was studied by Van der Hoeven (2020) and Van der Hoeven et al. (2021). We will use $B\coloneqq\mathrm{diam}(\mathcal{W})$, $C_{x}\coloneqq\mathrm{diam}(\mathcal{X})$, and $C_{y}\coloneqq\mathrm{diam}(\mathcal{Y})$ to denote the diameters of the sets $\mathcal{W}$, $\mathcal{X}$, and $\mathcal{Y}$, respectively.
The feedback observed by the learner depends on the problem setting. The most fundamental setting is the full information setup, where the true output $\bm{y}_{t}$ is observed as feedback at the end of each round $t$. This setup was extensively investigated in Sakaue et al. (2024). By contrast, our study investigates the following weaker feedback:
  • Bandit feedback: Only the value of the loss function is observed. Specifically, at the end of each round $t$, the learner observes the target loss value $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$ as feedback.
  • Delayed feedback: The feedback is observed with a certain delay. We consider a fixed $D$-round delay setting, i.e., no feedback is received for round $t\leq D$, and for $t>D$, the learner observes either full information feedback $\bm{y}_{t-D}$ or bandit feedback $L(\widehat{\bm{y}}_{t-D};\bm{y}_{t-D})$.
In this paper, we make the following assumptions:
Assumption 2.1.
(1) There exists $\nu>0$ such that for any distinct $\bm{y},\bm{y}^{\prime}\in\mathcal{Y}$, it holds that $\|\bm{y}-\bm{y}^{\prime}\|\geq\nu$. (2) For each $\bm{y}\in\mathcal{Y}$, the target loss function $L(\cdot;\bm{y})$ is defined on $\operatorname{conv}(\mathcal{Y})$, is non-negative, and is affine with respect to its first argument. (3) There exists $\gamma$ such that for any $\bm{y}^{\prime}\in\operatorname{conv}(\mathcal{Y})$ and $\bm{y}\in\mathcal{Y}$, it holds that $L(\bm{y}^{\prime};\bm{y})\leq\gamma\|\bm{y}^{\prime}-\bm{y}\|$ and $L(\bm{y}^{\prime};\bm{y})\leq 1$. (4) It holds that $L(\bm{y}^{\prime};\bm{y})=0$ only if $\bm{y}^{\prime}=\bm{y}$.
As discussed in Sakaue et al. (2024, Section 2.3), these assumptions are natural and hold for a broad range of problem settings and target loss functions, including SELF (see Section 2.4 for the formal definition).
2.2 Fenchel–Young Loss
We use the Fenchel–Young loss (Blondel et al., 2020) as the surrogate loss, which subsumes many representative surrogate losses, such as logistic loss, Conditional Random Field (CRF) loss (Lafferty et al., 2001), and SparseMAP (Niculae et al., 2018). See Blondel et al. (2020, Table 1) for more examples.
Definition 2.2 (Blondel et al. 2020, Fenchel–Young loss).
Let $\Omega:\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a regularization function with $\mathcal{Y}\subset\operatorname{dom}(\Omega)$. The Fenchel–Young loss generated by $\Omega$, denoted by $S_{\Omega}:\operatorname{dom}(\Omega^{\ast})\times\operatorname{dom}(\Omega)\rightarrow\mathbb{R}_{\geq 0}$, is defined as
$S_{\Omega}(\bm{\theta};\bm{y})\coloneqq\Omega^{\ast}(\bm{\theta})+\Omega(\bm{y})-\langle\bm{\theta},\bm{y}\rangle.$
The Fenchel–Young loss has the following properties, which will be useful in the subsequent discussion:
Proposition 2.3 (Blondel et al. 2020, Propositions 2 and 3 and Sakaue et al. 2024, Proposition 3).
[…]
$\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ $\coloneqq\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Omega(\bm{y})\>:\>\bm{y}\in\mathbb{R}^{d}\}$
$=\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Psi(\bm{y})\>:\>\bm{y}\in\operatorname{conv}(\mathcal{Y})\}.$
Then, for any $\bm{y}\in\mathcal{Y}$, $S_{\Omega}(\bm{\theta},\bm{y})$ is differentiable with respect to $\bm{\theta}$, and it satisfies $\nabla S_{\Omega}(\bm{\theta};\bm{y})=\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{y}.$ Furthermore, it holds that $S_{\Omega}(\bm{\theta};\bm{y})\geq\frac{\lambda}{2}\|\bm{y}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|^{2}.$
In what follows, let $S_{t}(\bm{W})\coloneqq S_{\Omega}(\bm{W}\bm{x}_{t};\bm{y}_{t})$ for simplicity. Importantly, from the properties of the Fenchel–Young loss, there exists some $b>0$ such that for any $\bm{W}\in\mathcal{W}$,
$\|\nabla S_{t}(\bm{W})\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}).$ (2)
Indeed, from Proposition 2.3 and Assumption 2.1, we have $\|\nabla S_{t}(\bm{W}_{t})\|_{\mathrm{F}}^{2}=\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|_{2}^{2}\|\bm{x}_{t}\|^{2}\leq C_{x}^{2}\kappa^{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|^{2}\leq\frac{2C_{x}^{2}\kappa^{2}}{\lambda}S_{t}(\bm{W}_{t})$, where we used $\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$ and $\|\cdot\|_{2}\leq\kappa\|\cdot\|$. Thus, (2) holds with $b=\frac{2C_{x}^{2}\kappa^{2}}{\lambda}$. Below, let $L_{t}(\bm{y})\coloneqq L(\bm{y};\bm{y}_{t})$, $\bm{G}_{t}\coloneqq\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$, and $L\coloneqq\max_{\bm{W}\in\mathcal{W}}\|\nabla S_{t}(\bm{W})\|_{\mathrm{F}}$.
2.3 Examples of Structured Prediction
We present several special cases of structured prediction along with specific parameter values introduced so far; see Sakaue et al. (2024, Section 2.3) for further details.
Multiclass classification
Let $\mathcal{Y}=\{\mathbf{e}_{1},\dots,\mathbf{e}_{d}\}$ and $\|\cdot\|=\|\cdot\|_{1}$. When using the 0-1 loss, $L(\bm{y}^{\prime};\bm{y})=\mathds{1}[\bm{y}^{\prime}\neq\bm{y}]$, the parameters in Assumption 2.1 are $\nu=2$ and $\gamma=\frac{1}{2}$. The logistic surrogate loss is a Fenchel–Young loss $S_{\Omega}$ generated by the entropy regularization function $\Omega=\mathsf{H}^{s}+I_{\Delta_{d}}$ (up to a constant factor), where $\mathsf{H}^{s}(\bm{y})\coloneqq-\sum_{i=1}^{d}y_{i}\log y_{i}$ and $\Delta_{d}$ is the $(d-1)$-dimensional probability simplex. Since $\Omega$ is a $1$-strongly convex function with respect to $\|\cdot\|_{1}$, we have $\lambda=1$.
Multilabel classification
Let $\mathcal{Y}=\{0,1\}^{d}$ and $\|\cdot\|=\|\cdot\|_{2}$. When using the Hamming loss as the target loss function $L(\bm{y}^{\prime};\bm{y})=\frac{1}{d}\sum_{i=1}^{d}\mathds{1}[y_{i}^{\prime}\neq y_{i}]$, Assumption 2.1 is satisfied with $\nu=1$ and $\gamma=\frac{1}{d}$. The SparseMAP surrogate loss $S_{\Omega}(\bm{\theta},\bm{y})=\frac{1}{2}\|\bm{y}-\bm{\theta}\|_{2}^{2}-\frac{1}{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{\theta}\|_{2}^{2}$ is a Fenchel–Young loss generated by $\Omega=\frac{1}{2}\|\cdot\|^{2}+I_{\operatorname{conv}(\mathcal{Y})}$. Since $\Omega$ is 1-strongly convex with respect to $\|\cdot\|_{2}$, we have $\lambda=1$.
Ranking
We consider predicting the ranking of $m$ items. Let $\|\cdot\|=\|\cdot\|_{1}$, $d=m^{2}$, and $\mathcal{Y}\subset\{0,1\}^{d}$ be the set of all vectors representing $m\times m$ permutation matrices. We use the target loss function that counts mismatches, $L(\bm{y}^{\prime};\bm{y})=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[y_{i,j_{i}}^{\prime}\neq y_{i,j_{i}}]$, where $j_{i}\in[m]$ is a unique index with $y_{ij_{i}}=1$ for each $i\in[m]$. In this case, the parameters in Assumption 2.1 satisfy $\nu=4$ and $\gamma=\frac{1}{2m}$. We use a surrogate loss given by $S_{\Omega}(\bm{\theta};\bm{y})=\langle\bm{\theta},\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{y}\rangle+\frac{1}{\zeta}\mathsf{H}^{s}(\widehat{\bm{y}}_{\Omega}(\bm{\theta}))$, where $\Omega=-\frac{1}{\zeta}\mathsf{H}^{s}+I_{\operatorname{conv}(\mathcal{Y})}$ and $\zeta$ is a parameter controlling the regularization strength. The first term in $S_{\Omega}$ measures the affinity between $\bm{\theta}$ and $\bm{y}$, while the second term evaluates the uncertainty of $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$. Since $\Omega$ is $\frac{1}{m\zeta}$-strongly convex, we have $\lambda=\frac{1}{m\zeta}$.
2.4 Structured Encoding Loss Function (SELF)
Here we introduce a common class of target loss functions, called the Structured Encoding Loss Function (SELF). A target loss function is SELF if it can be expressed as
$L(\bm{y}_{t};\widehat{\bm{y}}_{t})=\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}+\bm{b}\rangle+c(\bm{y}_{t}),$ (3)
where $\bm{b}\in\mathbb{R}^{d}$ is a constant vector, $\bm{V}\in\mathbb{R}^{d\times d}$ is a constant matrix, and $c:\mathcal{Y}\to\mathbb{R}$ is a function. The following loss examples, taken from Blondel (2019, Appendix A), belong to the SELF class:
  • Multiclass classification: the 0-1 loss is a SELF with $\bm{V}=\bm{1}\bm{1}^{\top}-\bm{I}$, $\bm{b}=\bm{0}$, and $c(\bm{y})=0$.
  • Multilabel classification: the Hamming loss, $L(\bm{y}^{\prime};\bm{y})=\frac{1}{d}\sum_{i=1}^{d}\mathds{1}[y^{\prime}_{t,i}\neq y_{i}]$, is a SELF with $\bm{V}=-\frac{2}{d}\bm{I}$, $\bm{b}=\frac{\bm{1}}{d}$, and $c(\bm{y})=\frac{1}{d}\langle\bm{y},\bm{1}\rangle$, where the last factor is constant if the number of correct labels is fixed.
  • Ranking: the Hamming loss $L(\bm{y}^{\prime};\bm{y})=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[y_{i,j_{i}}^{\prime}\neq y_{i,j_{i}}]$, where $j_{i}\in[m]$ is a unique index with $y_{i,j_{i}}=1$ for each $i\in[m]$, is a SELF with $\bm{V}=-\frac{1}{m}\bm{I}$, $\bm{b}=\bm{0}$, and $c(\bm{y})=1$.
Following Sakaue et al. (2024), this study assumes that the target loss function $L$ is a SELF.
2.5 Randomized Decoding
Algorithm 1 Randomized decoding $\phi_{\Omega}$
1:$\bm{\theta}\in\mathbb{R}^{d}$
2:$\widehat{\bm{y}}_{\Omega}(\bm{\theta})\leftarrow\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Psi(\bm{y})\>:\>\bm{y}\in\operatorname{conv}(\mathcal{Y})\}$
3:$\bm{y}^{\ast}\leftarrow\operatorname*{arg\,min}\{\|\bm{y}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|\>:\>\bm{y}\in\mathcal{Y}\}$
4:$\Delta^{\ast}\leftarrow\|\bm{y}^{\ast}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|,\>p\leftarrow\min\{1,2\Delta^{\ast}/\nu\}$
5:Sample $Z\sim\mathrm{Ber}(p)$
6:if $Z=0$ then $\bm{\widehat{y}}\leftarrow\bm{y}^{\ast}$
7:if $Z=1$ then $\bm{\widehat{y}}\leftarrow\bm{\widetilde{y}}$ where $\bm{\widetilde{y}}$ is randomly drawn from $\mathcal{Y}$ so that $\mathbb{E}\left[\bm{\widetilde{y}}|Z=1\right]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$
8:$\phi_{\Omega}(\bm{\theta})=\bm{\widehat{y}}$
The procedure of converting the estimated score $\bm{\theta}$ into a structured output $\bm{\widehat{y}}$ is called decoding. For this, we employ randomized decoding (Sakaue et al., 2024), which plays an essential role particularly in deriving an upper bound independent of the output set size $K=\lvert\mathcal{Y}\rvert$ in Section 3.4. The randomized decoding (Algorithm 1) selects either the closest $\bm{y}^{*}\in\mathcal{Y}$ to $\widehat{\bm{y}}_{\Omega}(\bm{\theta})\in\operatorname{conv}(\mathcal{Y})$ or a random $\widetilde{\bm{y}}\in\mathcal{Y}$ satisfying $\mathbb{E}[\widetilde{\bm{y}}\mid Z=1]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$, where $Z$ follows a Bernoulli distribution with a parameter $p$. Intuitively, the parameter $p$ is chosen so that if $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ is close to $\bm{y}^{*}$, the decoding is more likely to return $\bm{y}^{*}$; otherwise, it is more likely to return $\widetilde{\bm{y}}$, reflecting uncertainty. An important property satisfied by the randomized decoding is the following lemma, which we use in the subsequent analysis:
Lemma 2.4 (Sakaue et al. 2024, Lemma 4).
For any $(\bm{\theta},\bm{y})\in\mathbb{R}^{d}\times\mathcal{Y}$, the randomized decoding $\phi_{\Omega}$ satisfies
$\mathbb{E}[L(\phi_{\Omega}(\bm{\theta});\bm{y})]\leq\frac{4\gamma}{\lambda\nu}S_{\Omega}(\bm{\theta};\bm{y}),$
where the expectation is taken with respect to the internal randomness of the randomized decoding.
3 Bandit Feedback
In this section, we present two online structured prediction algorithms in the bandit feedback setup and analyze their regret. Our results here are mostly special cases of the regret bounds obtained when handling bandit and delayed feedback (Section 5). Nevertheless, by focusing on the case without delay, we provide a clearer exposition of the core ideas.
3.1 Randomized Decoding with Uniform Exploration
Algorithm 2 Randomized decoding with uniform exploration (RDUE) $\psi_{\Omega}$
1:$\bm{\theta}\in\mathbb{R}^{n}$, $q\in[0,1]$
2:Sample $X\sim\mathrm{Ber}(q)$
3:if $X=0$ then $\bm{\widehat{y}}\leftarrow\phi_{\Omega}(\bm{\theta})$
4:if $X=1$ then Sample $\bm{y}^{\ast}\sim\mathrm{Unif}(\mathcal{Y})$ and $\bm{\widehat{y}}\leftarrow\bm{y}^{\ast}$
5:$\psi_{\Omega}(\bm{\theta})=\bm{\widehat{y}}$
Here, we discuss the properties of the decoding function, Randomized Decoding with Uniform Exploration (RDUE), which will be used in both algorithms to convert scores into outputs. As discussed in Section 2.5, in online structured prediction with full information feedback, the randomized decoding (Algorithm 1) was introduced as a decoding function (Sakaue et al., 2024). However, naively applying the randomized decoding does not lead to a satisfactory regret bound under bandit feedback. We extend the framework of the randomized decoding to handle bandit feedback effectively.
RDUE (Algorithm 2) is a procedure that, with probability $q\in[0,1]$, selects $\widehat{\bm{y}}$ uniformly at random from $\mathcal{Y}$, and with probability $1-q$, selects the output of the randomized decoding. Using RDUE, we define $p_{t}(\bm{y})$ as the probability that $\widehat{\bm{y}}_{t}$ coincides with $\bm{y}$ at round $t$. Note that for any $\bm{y}\in\mathcal{Y}$, it holds that $p_{t}(\bm{y})\geq\frac{q}{K}.$ Furthermore, similar to the property of the randomized decoding in Lemma 2.4, RDUE satisfies the following property:
Lemma 3.1.
For any $(\bm{\theta},\bm{y})\in\mathbb{R}^{d}\times\mathcal{Y}$, RDUE $\psi_{\Omega}$ satisfies
$\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]\leq\frac{4\gamma}{\lambda\nu}(1-q)S_{\Omega}(\bm{\theta};\bm{y})+q\frac{K-1}{K},$
where the expectation is taken with respect to the internal randomness of RDUE.
Proof.
Since the output of the randomized decoding is chosen with probability $1-q$, and a uniformly random output is chosen with probability $q$, we have $\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]=(1-q)\mathbb{E}\left[L(\phi_{\Omega}(\bm{\theta});\bm{y})\right]+q\frac{K-1}{K},$ where we used $L(\cdot;\cdot)\leq 1$ and $\phi_{\Omega}$ is the randomized decoding. Hence, combining this with Lemma 2.4, we obtain the desired bound. ∎
Additionally, this section makes the following assumption:
Assumption 3.2.
There exists $a\in(0,1)$ such that
$\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t})+q.$
Here, $\mathbb{E}_{t}\left[{\cdot}\right]$ denotes the conditional expectation given the random variables $\widehat{\bm{y}}_{1},\dots,\widehat{\bm{y}}_{t-1}$.
This assumption can be satisfied by using RDUE and letting $a\leq 1-\frac{4\gamma}{\lambda\nu}(1-q)$ if $\lambda>\frac{4\gamma}{\nu}(1-q)$, according to Lemma 3.1. In what follows, let $a=1-\frac{4\gamma}{\lambda\nu}$. Note that $\lambda\geq\frac{4\gamma}{\nu}$ holds in the cases of multiclass classification, multilabel classification, and ranking, as discussed in Section 2.3.
3.2 Online Gradient Descent
The algorithm in this section uses the adaptive Online Gradient Descent (OGD, Streeter & McMahan 2010) as $\mathsf{ALG}$. OGD updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ using a gradient $\bm{G}_{t}$ and learning rate $\eta_{t}$ by $\bm{W}_{t+1}\leftarrow\Pi_{\mathcal{W}}\left(\bm{W}_{t}-\eta_{t}\bm{G}_{t}\right),$ where $\Pi_{\mathcal{W}}(\bm{Z})=\operatorname*{arg\,min}_{\bm{X}\in\mathcal{W}}\|\bm{X}-\bm{Z}\|_{\mathrm{F}}$. OGD with appropriately chosen learning rate $\eta_{t}$ achieves the following bound:
Lemma 3.3 (Orabona 2019, Theorem 4.14).
Suppose that we set the learning rate to $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\bm{G}_{i}\|_{\mathrm{F}}^{2}}}$ and do not update on rounds when $\bm{G}_{t}$ is the all-zero matrix. Then, for any $\bm{U}\in\mathcal{W}$, OGD achieves $\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}}.$
3.3 $O(\sqrt{KT})$ Regret Algorithm
Here, we present an algorithm that achieves a regret upper bound of $O(\sqrt{KT})$.
Algorithm based on inverse-weighted estimator
In the bandit setting, the true output $\bm{y}_{t}$ is not observed, and thus it is necessary to estimate the gradient required for updating $\bm{W}_{t}$. To deal with this, we use the following inverse-weighted gradient estimator:
$\widetilde{\bm{G}}_{t}\coloneqq\frac{\mathds{1}[\widehat{\bm{y}}_{t}=\bm{y}_{t}]}{p_{t}(\bm{y}_{t})}\bm{G}_{t},$ (4)
where we recall that $\bm{G}_{t}=\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$. Note that $\widetilde{\bm{G}}_{t}$ is unbiased, i.e., $\mathbb{E}\big{[}\widetilde{\bm{G}}_{t}\big{]}=\bm{G}_{t}$. We use RDUE with $q=B\sqrt{K/T}$ as the decoding function (assuming $T\geq B^{2}K$ for simplicity). For $\mathsf{ALG}$, we employ the adaptive OGD in Section 3.2 with the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
Remark 3.4.
This study defines the bandit feedback as the value of the target loss function $L_{t}(\widehat{\bm{y}}_{t})$. Note, however, that the above algorithm operates using only the weaker feedback of $\mathds{1}\left[\widehat{\bm{y}}_{t}\neq\bm{y}_{t}\right]$.
Regret bounds and analysis
The above algorithm achieves the following regret bound:
Theorem 3.5.
The regret of the above algorithm is bounded by $\mathbb{E}[\mathcal{R}_{T}]\leq\left(\frac{b}{2a}+1\right)B\sqrt{KT}.$
The $O(\sqrt{KT})$ bound has an optimal dependency on $T$ and matches the $\sqrt{T}$ lower bound in the special case of online multiclass classification with bandit feedback (Van der Hoeven et al., 2021, Corollary 1). Furthermore, our bound improves the existing $O(\sqrt{KT})$ bound by Van der Hoeven (2020) by a factor of $\sqrt{K}$. Note that, due to differences in the target loss function, our result is not directly comparable to the $\sqrt{KT}$ bound in Van der Hoeven et al. (2021). A more detailed discussion can be found in Appendix C.
Proof.
From the convexity of $S_{t}$ and the unbiasedness of $\widetilde{\bm{G}}_{t}$, we have $\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]=\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right].$ From Lemma 3.3, this is further upper bounded as $\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]\leq\sqrt{2}B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}}\right]}\leq B\sqrt{\frac{2bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]},$ where in the first inequality we used Jensen’s inequality and in the last inequality we used $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}),$ which follows from $p_{t}(\bm{y})\geq K/q$ and (2). Therefore, from Assumption 3.2, we have $\mathbb{E}[\mathcal{R}_{T}]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\left((1-a)S_{t}(\bm{W}_{t})-S_{t}(\bm{U})\right)}\right]+qT\leq B\sqrt{\frac{2bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT\leq\frac{bB^{2}K}{2aq}+qT,$ where the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Finally, substituting $q=B\sqrt{K/T}$ into the last inequality yields the desired bound. ∎
We can also prove the following high-probability bound:
Theorem 3.6.
Let $\delta\in(0,1)$. Then with probability at least $1-\delta$, the same algorithm as in Theorem 3.5, but with a different choice of $q$, achieves the regret bound of $\mathcal{R}_{T}=O\Big{(}\sqrt{KT\log(1/\delta)}+\log(1/\delta)\Big{)}$, where we omit the dependencies on parameters other than $K$, $T$, and $\delta$.
A more precise statement and proof of this theorem are provided in Section D.2. To prove this theorem, we follow the analysis of Theorem 3.5 and use Bernstein’s inequality. To address the challenges posed by the randomness introduced by bandit feedback, we adopt an approach similar to that used by Van der Hoeven et al. (2021), and arguably, we have successfully simplified their analysis.
3.4 $O(T^{2/3})$ Regret Algorithm
While the $O(\sqrt{KT})$ regret bound given above is desirable in terms of the dependence on $T$, the dependence on $K=\lvert\mathcal{Y}\rvert$ is undesirable for general structured prediction. In fact, we have $K=\binom{d}{m}$ in multilabel classification with $m$ correct labels and $K=m!$ in ranking with $m$ items. To address this issue, we present an algorithm that significantly improves the dependence on $K$ when the target loss function belongs to a special class of SELF satisfying the following assumptions:
Assumption 3.7.
(i) $\bm{V}$ is invertible, and $\bm{b}$ and $c(\cdot)$ are known and non-negative. (ii) Let $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where $\mu$ is the uniform distribution over $\mathcal{Y}$. At least one of the following two conditions holds: (ii-a) $\bm{Q}$ is invertible, or (ii-b) for any $\bm{y}\in\mathcal{Y}$, $\bm{V}\bm{y}$ lies in the linear subspace spanned by vectors in $\mathcal{Y}$. (iii) For some $\omega>0$, it holds that
$\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)\leq\omega.$
The first condition is true in the examples in Section 2.4, assuming that the number of correct labels is fixed in multilabel classification. The second one is satisfied if $\mathcal{Y}$ consists of $d$ linearly independent vectors or $\bm{V}$ is proportional to the identity matrix; either is true in those examples. It is also not difficult to derive bounds on $\omega$ in those examples.
Algorithm based on pseudo-inverse matrix estimator
As in the case of Section 3.3, we begin by providing a method to estimate the gradient. Define $\bm{P}_{t}\coloneqq\mathbb{E}_{\bm{y}\sim p_{t}}[\bm{y}\bm{y}^{\top}]$. Then, we define the estimator $\widetilde{\bm{y}}_{t}$ of $\bm{y}_{t}$ by
$\widetilde{\bm{y}}_{t}\coloneqq{\bm{V}}^{-1}\bm{P}_{t}^{+}\widehat{\bm{y}}_{t}\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle,$
where $\bm{P}_{t}^{+}$ is the Moore–Penrose pseudo-inverse matrix of $\bm{P}_{t}$. It is important to note that, given that $\bm{b}$ and $c(\cdot)$ are known, $\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle=L_{t}(\widehat{\bm{y}}_{t})-\langle\widehat{\bm{y}}_{t},\bm{b}\rangle-c(\bm{y}_{t})$ can be computed. Note that $\widetilde{\bm{y}}_{t}$ is unbiased, i.e., $\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]=\bm{y}_{t}$ from the first requirement of Assumption 3.7.
Using this $\widetilde{\bm{y}}_{t}$, we define the gradient estimator $\widetilde{\bm{G}}_{t}$ by
$\widetilde{\bm{G}}_{t}\coloneqq\left(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t}\right)\bm{x}_{t}^{\top},$ (5)
whose expectation is $\mathbb{E}\big{[}\widetilde{\bm{G}}_{t}\big{]}=\bm{G}_{t}.$ Our estimator is based on the estimators used in adversarial linear bandits and adversarial combinatorial full-bandits (Dani et al., 2007; Abernethy et al., 2008; Cesa-Bianchi & Lugosi, 2012).
We use RDUE with $q=\left(\frac{4\omega B^{2}C_{x}^{2}}{T}\right)^{1/3}$ as the decoding function (assuming $T\geq 4\omega B^{2}C_{x}^{2}$ for simplicity). For updating $\bm{W}_{t}$, we employ the adaptive OGD in Section 3.2 as $\mathsf{ALG}$ with the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
Regret bounds
The above algorithm achieves the following regret bound, which does not directly depend on $K$:
Theorem 3.8.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\big{(}\omega^{1/3}\left(BC_{x}T\right)^{2/3}\big{)}.$
The proof can be found in Section D.3. By using the estimator based on the pseudo-inverse matrix, we can upper bound the second moment of the gradient estimator $\widetilde{\bm{G}}_{t}$ without $K$, which allows us to establish the improved regret bound that does not explicitly depend on $K$.
The regret bound in Theorem 3.8 yields the different bounds on each problem setup as follows:
Corollary 3.9.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((BC_{x}dT)^{2/3}\right)$ in multiclass classification with the 0-1 loss, $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((d^{5}/m(d-m))^{1/3}(BC_{x}T)^{2/3}\right)$ in multilabel classification with $m$ correct labels and the Hamming loss, and $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(m^{5/3}(BC_{x}T)^{2/3}\right)$ in ranking with the number of items $m$ and the Hamming loss.
The proof of Corollary 3.9 is deferred to Section D.4. The bound for multilabel classification with $m$ correct labels significantly improved the $O(\sqrt{KT})$ bound in Section 3.3, which has a dependency of $\sqrt{\binom{d}{m}}$, and the bound for ranking significantly improved the $O(\sqrt{KT})$ bound in Section 3.3, which has a dependency of $\sqrt{m!}$.
4 Delayed Full-Information Feedback
This section discusses online structured prediction with delayed full-information feedback and provides an algorithm that achieves a surrogate regret bound of $O(D^{2/3}T^{1/3})$, a better bound than $O(\sqrt{DT})$ that can be achieved with a standard OCO algorithm under delayed feedback (Joulani et al., 2013). Below, we make the following assumption.
Assumption 4.1.
There exists a constant $a\in(0,1)$ which satisfies
$\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t}).$
From Lemma 2.4, if $\lambda>\frac{4\gamma}{\nu}$, this condition is satisfied with $a=1-\frac{4\gamma}{\lambda\nu}$ by using the randomized decoding. We suppose that such a decoding function is used in this section.
Algorithm
For updating $\bm{W}_{t}$, we employ the Optimistic Delayed Adaptive FTRL (ODAFTRL) algorithm proposed by Flaspohler et al. (2021). In ODAFTRL, given a gradient $\bm{G}_{t}$ at round $t$, $\bm{W}_{t}$ is updated as
$\bm{W}_{t+1}=\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}\left\{\sum_{i=1}^{t-D}\left\langle\bm{G}_{i},\bm{W}\right\rangle+\frac{\lambda_{t}\|\bm{W}\|_{\mathrm{F}}^{2}}{2}\right\},$
where $\lambda_{t}\geq 0$ is the regularization parameter. Due to space constraints, the details of the algorithm are provided in Section E.1. By updating $\lambda_{t}$ using an AdaHedge-type algorithm called AdaHedgeD, ODAFTRL achieves the following regret upper bound:
Lemma 4.2 (Informal version of Flaspohler et al. 2021, Theorem 12).
Consider the setting with delayed full-information feedback. Then, for any $\bm{U}\in\mathcal{W}$, ODAFTRL with the AdaHedgeD update of $\lambda_{t}$ achieves $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})\!-\!S_{t}(\bm{U}))}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}=O\Big{(}\sqrt{\sum_{t=1}^{T}{(\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\!+\!D\|\bm{G}_{t}\|_{\mathrm{F}})}}\Big{)}.$
Regret bounds and analysis
The algorithm described above achieves the following bound:
Theorem 4.3.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(D^{2/3}T^{1/3}).$
Here, we provide a proof sketch; the complete proof can be found in Section E.2.
Proof sketch.
Lemma 4.2 with $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$ in (2) and Cauchy–Schwarz yields $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}=O(\sqrt{S_{1:T}}+(D^{2}TS_{1:T})^{1/4})$, where $S_{1:T}=\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$. Hence, from Assumption 4.1, we have $\mathbb{E}[\mathcal{R}_{T}]\leq\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}=O(\sqrt{S_{1:T}}+(D^{2}TS_{1:T})^{1/4})-aS_{1:T}=O(D^{2/3}T^{1/3}),$ where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. ∎
We can also prove the following high-probability bound; see Section E.3 for the proof:
Theorem 4.4.
For any $\delta\in(0,1)$, with probability at least $1-\delta$, the above algorithm achieves $\mathcal{R}_{T}=O(\log({1}/{\delta})+D^{2/3}T^{1/3}).$
5 Delayed Bandit Feedback
Given the results so far, it is natural to explore online structured prediction with delayed bandit feedback. We construct algorithms for this setup by combining the theoretical developments from Section 3 and Section 4. This section assumes Assumption 3.2, as in Section 3.
5.1 $O(\sqrt{DKT})$ Regret Algorithm
We use RDUE with $q=B\sqrt{DK/T}$ as the decoding function (assuming $T\geq DB^{2}K$ for simplicity), the gradient estimator $\widetilde{\bm{G}}_{t}$ in (4), and ODAFTRL with the AdaHedgeD update as $\mathsf{ALG}$. This algorithm attains the following bound:
Theorem 5.1.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(\sqrt{DKT}).$
The proof can be found in Section F.2. Due to the introduction of the delay, the regret bound worsens by a factor of $\sqrt{D}$ compared to the non-delayed case, which is natural when considering analyses of adversarial cases in delayed feedback (Joulani et al., 2013; Zimmert & Seldin, 2020; Manwani & Agarwal, 2023).
5.2 $O(D^{1/3}T^{2/3})$ Regret Algorithm
Here, we make the same assumptions on the target loss function as in Section 3.4. We provide an algorithm that improves the dependence on $K$ from Section 5.1. We use RDUE with $q=\left(\frac{\omega B^{2}C_{x}^{2}D}{T}\right)^{1/3}$ as the decoding function (assuming $T\geq\omega B^{2}C_{x}^{2}D$ for simplicity), the gradient estimator $\widetilde{\bm{G}}_{t}$ in (5), and ODAFTRL with the AdaHedgeD update as $\mathsf{ALG}$. This algorithm achieves the following bound:
Theorem 5.2.
The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(D^{1/3}T^{2/3}).$
The proof can be found in Section F.3. Due to the introduction of the delay, the regret bound worsens by a factor of $D^{1/3}$ compared to the non-delayed bandit feedback case.
6 Experiments
Figure 1: A box plot of error rate of the MNIST experiment for multiclass classification with bandit feedback.
This section presents numerical experiment results for online multiclass classification under bandit feedback. We compare three algorithms: Gaptron (Van der Hoeven, 2020) with logistic loss and hinge loss as surrogate losses, Gappletron (Van der Hoeven et al., 2021) with logistic loss as the surrogate loss, and our proposed algorithm from Section 5.1. The parameters for each algorithm are set based on their theoretical values. We use the MNIST dataset (LeCun et al., 2010), a dataset of digit images. The diameter $B$ of $\mathcal{W}$ is fixed at $10$. We repeated the experiment for 20 times, and the boxplot of the obtained misclassification rates is summarized in Figure 1. From Figure 1, we observe that our method achieves the lowest misclassification rate. Despite not being specialized for multiclass classification, our approach outperforms existing algorithms designed for multiclass classification on real data with $K=10$. Further experiments can be found in Appendix G, and related discussions are provided in Appendix C.
7 Conclusion
In this study, we developed multiple algorithms for online structured prediction under bandit and delayed feedback and analyzed their surrogate regret. Among several contributions, a noteworthy one is the algorithm for bandit feedback whose surrogate regret bound does not explicitly depend on the label set size $K$, which is achieved by leveraging the pseudo-inverse matrix estimator. Several interesting research directions remain. One is to investigate surrogate regret lower bounds under bandit feedback and delayed feedback. In the existing $\Omega(\sqrt{T})$ lower bound for bandit feedback, the optimal dependence on $d$ is unexplored. Investigating the tightness of the $O(D^{2/3}T^{1/3})$ upper bound for delayed and full information feedback is also an important question. Another key direction is extending our results to settings where the delay time is not constant in delayed feedback, which is crucial for some practical applications.
Acknowledgments
TT is supported by JST ACT-X Grant Number JPMJAX210E and JSPS KAKENHI Grant Number JP24K23852, SS is supported by supported by JST ERATO Grant Number JPMJER1903, and KY is supported by JSPS KAKENHI Grant Number JP24H00703.
Appendix A Notation
Table 2 summarizes the symbols used in this paper.
Table 2: Notation
Symbol Meaning
$T\in\mathbb{N}$ time horizon
$d\in\mathbb{N}$ dimension of output space $\mathcal{Y}$
$B\coloneqq\mathrm{diam}(\mathcal{W})$ diameter of $\mathcal{W}$
$C_{x}\coloneqq\mathrm{diam}(\mathcal{X})$ diameter of input space $\mathcal{X}$
$C_{y}\coloneqq\mathrm{diam}(\mathcal{Y})$ diameter of output space $\mathcal{Y}$
$K\coloneqq|\mathcal{Y}|$ cardinality of $\mathcal{Y}$
$D\in\mathbb{N}$ fixed delayed time
$\mathsf{ALG}$ algorithm for updating linear estimators
$\widehat{\bm{y}}_{t}\in\mathcal{Y}$ output chosen by the learner at round $t$
$p_{t}(\bm{y})$ probability that $\bm{y}$ is chosen as $\widehat{\bm{y}}_{t}$ at time $t$
$L:\mathcal{Y}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ target loss function
$L_{t}(\widehat{\bm{y}}_{t})\coloneqq L(\widehat{\bm{y}}_{t};\bm{y}_{t})$ the value of target loss $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$
$S_{\Omega}:\mathbb{R}^{n}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ Fenchel–Young loss generated by $\Omega$
$S_{t}(\bm{W})\coloneqq S_{\Omega}(\bm{W}\bm{x}_{t};\bm{y}_{t})$ shorthand of surrogate loss $S_{\Omega}(\bm{W}\bm{x}_{t};\bm{y}_{t})$
$L\coloneqq\max_{\bm{W}\in\mathcal{W}}\|\nabla S_{t}(\bm{W})\|_{\mathrm{F}}$ upper bound of the gradient
$\bm{G}_{t}\coloneqq\nabla S_{t}(\bm{W}_{t})$ gradient of surrogate loss
$\mathcal{R}_{T}\coloneqq\sum_{t=1}^{T}{(L_{t}(\widehat{\bm{y}}_{t})-S_{t}(\bm{U}))}$ surrogate regret
$\mu$ uniform distribution over $\mathcal{Y}$
$\bm{P}_{t}=\mathbb{E}_{\bm{y}\sim p_{t}}[\bm{y}\bm{y}^{\top}]$ second moment matrix under $p_{t}$
$\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$ second moment matrix under $\mu$
$\lambda_{\min}(\bm{A})$ minimum eigenvalue of matrix $\bm{A}$
$\omega$ upper bound of $\operatorname*{tr}({\bm{V}}^{-1}\bm{Q}^{+}\bm{V})$
$\mathbb{E}_{t}\left[{\cdot}\right]$ conditional expectation given $\widehat{\bm{y}}_{1},\dots,\widehat{\bm{y}}_{t-1}$
Appendix B Additional Related Work
We discuss additional related work that could not be included above.
Structured prediction
Before the introduction of the Fenchel–Young loss framework, Niculae et al. (2018) proposed SparseMAP, which used the squared $\ell_{2}$-norm regularization. The Fenchel–Young loss, described in Section 2.2, is built upon the idea of SparseMAP. The Structure Encoding Loss Function (SELF) was introduced by Ciliberto et al. (2016, 2020) to analyze the relationship between surrogate and target losses, a concept known as Fisher consistency. For more extensive literature, we refer the reader to Sakaue et al. (2024, Appendix A).
Online classification with full and bandit feedback
In the full information setup, PERCEPTRON is one of the most representative algorithms for binary classification (Rosenblatt, 1958), and the multiclass setting has also been extensively studied (Crammer & Singer, 2003; Fink et al., 2006). Online logistic regression is another relevant research stream, with Foster et al. (2018) being a particularly representative study. The study of the bandit setup was initiated by Kakade et al. (2008), and it has since been extensively explored in subsequent research (Hazan & Kale, 2011; Beygelzimer et al., 2017; Foster et al., 2018). However, to the best of our knowledge, no prior work has addressed general structured prediction under bandit feedback. A most related study is the work by Gentile & Orabona (2014b), which investigated online multilabel classification and ranking. However, their setting assumes access to feedback of the form $\{\mathds{1}[\bm{y}_{t,i}\neq\widehat{\bm{y}}_{t,i}]\}_{i}$, which is more informative than bandit feedback and differs from our setup. Van der Hoeven (2020) explicitly introduced the surrogate regret in the context of online multiclass classification. This study has been extended to the setting where observations are determined by a directed graph (Van der Hoeven et al., 2021) and to structured prediction scenarios (Sakaue et al., 2024). For a more extensive overview of the literature on online classification, we refer the reader to Van der Hoeven (2020).
Delayed feedback
The study of delayed feedback began with Weinberger & Ordentlich (2002). Since then, it has been extensively explored in various online learning settings, primarily in the full information setup of online convex optimization (Mesterharm, 2005; Joulani et al., 2013, 2016; Flaspohler et al., 2021). Algorithms for delayed bandit feedback have been studied mainly in the context of multi-armed bandits and their variants (Cesa-Bianchi et al., 2016; Zimmert & Seldin, 2020; Ito et al., 2020; Masoudian et al., 2022; Van der Hoeven et al., 2023). In the context of online classification, research considering delay is scarce; the only work is Manwani & Agarwal (2023) to our knowledge.
Appendix C Discussion on the Difference in Surrogate Loss Functions
As in (1), the surrogate regret $\mathcal{R}_{T}$ is defined by $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}$, which means the choice of the surrogate loss $S$ affects the bound on the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$. Van der Hoeven et al. (2021, Theorem 1), which applies to a more general setup than bandit feedback, implies $\mathcal{R}_{T}=O(K\sqrt{T})$ for the bandit setup with $S$ being a logistic loss defined with the base-$K$ logarithm. On the other hand, our bound of $\mathcal{R}_{T}=O(\sqrt{KT})$ applies to the logistic loss $S$ defined with the base-$2$ logarithm. As a result, while our bound on $\mathcal{R}_{T}$ is better, the $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$ term can be worse; this is why we cannot directly compare our $O(\sqrt{KT})$ bound with the $O(K\sqrt{T})$ bound in Van der Hoeven et al. (2021, Theorem 1). We may use the decoding procedure in Van der Hoeven et al. (2021), instead of RDUE, to recover their bound that applies to the base-$K$ logistic loss. It should be noted that their method is specific to multiclass classification; naively extending their method to structured prediction formulated as $|\mathcal{Y}|$-class classification results in the undesirable dependence on $K=|\mathcal{Y}|$, as is also discussed in Sakaue et al. (2024). By contrast, our pseudo-inverse estimator, combined with RDUE, can rule out the explicit dependence on $K$, at the cost of the increase from $\sqrt{T}$ to $T^{2/3}$.
Appendix D Omitted Details of Section 3
This section provides the omitted details of Section 3.
D.1 Concentration inequality
To prove high probability regret bounds, we use the following concentration inequality.
Lemma D.1 (Bernstein’s inequality, e.g., Cesa-Bianchi & Lugosi 2006, Lemma A.8).
Let $Z_{1},\ldots,Z_{T}$ be a martingale difference sequence and $\delta\in(0,1)$. If there exist $a$ and $v$ which satisfy $|Z_{t}|\leq a$ for any $t\in[T]$ and $\sum_{t=1}^{T}{\mathbb{E}_{t}\left[{Z_{t}^{2}}\right]}\leq v$ , then with probability at least $1-\delta$, it holds that
$\sum_{t=1}^{T}{Z_{t}}\leq\sqrt{2v\log\frac{1}{\delta}}+\frac{\sqrt{2}}{3}a\log\frac{1}{\delta}.$
D.2 Proof of Theorem 3.6
Here, we provide the proof of Theorem 3.6. Hereafter, we let $S_{\max}=\max_{\bm{W}\in\mathcal{W}}S_{t}(\bm{W})$ and $\widehat{S}_{t}(\bm{W})=v_{t}S_{t}(\bm{W})=\frac{\mathds{1}[\bm{y}_{t}=\widehat{\bm{y}}_{t}]}{p_{t}(\widehat{\bm{y}}_{t})}S_{t}(\bm{W})$. The following theorem is the formal version of Theorem 3.6:
Theorem D.2 (Formal version of Theorem 3.6).
Consider the bandit and non-delayed setup. Let
$\mathcal{C}=\left(\frac{3}{2(a+\xi-1)}+1\right)KS_{\max}\log(2/\delta)+\frac{B^{2}Kb}{2(1-\xi)}.$
Then, for any $T\geq\mathcal{C}$ and $\delta\in(0,1/2)$, with probability at least $1-\delta$, the algorithm in Section 3.3 with $q=\sqrt{\mathcal{C}/T}$ achieves
$\mathcal{R}_{T}\leq 2\sqrt{\mathcal{C}T}+\sqrt{2\log(2/\delta)}(\mathcal{C}T)^{1/4}+\left(\frac{1-a}{2(a+\xi-1)}+2\right)\log(2/\delta).$
Before proving this theorem, we provide the following lemma:
Lemma D.3.
It holds that
$\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-\widehat{S}_{t}(\bm{U})\right)\leq\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{W}_{t})\right)+qT+\sqrt{2}B\sqrt{\frac{b}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}.$
Proof.
We have
$\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-\widehat{S}_{t}(\bm{U})\right)=\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-\widehat{S}_{t}(\bm{W}_{t})\right)+\sum_{t=1}^{T}\left(\widehat{S}_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{U})\right).$
From Assumption 3.2, the first term is bounded as
$\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-\widehat{S}_{t}(\bm{W}_{t})\right)\leq\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{W}_{t})\right)+qT,$
and the second term is bounded as
$\sum_{t=1}^{T}\left(\widehat{S}_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{U})\right)$ $\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}=\sqrt{2}B\sqrt{\sum_{t=1}^{T}v_{t}^{2}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}$
$\leq\sqrt{2}B\sqrt{b\sum_{t=1}^{T}v_{t}^{2}S_{t}(\bm{W}_{t})}\leq\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})},$
where we used Lemma 3.3 and $v_{t}\leq K/q$. Combining the above three, we obtain
$\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-\widehat{S}_{t}(\bm{U})\right)\leq\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{W}_{t})\right)+qT+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})},$
which completes the proof. ∎
Proof of Theorem D.2.
The surrogate regret can be decomposed as
$\mathcal{R}_{T}=\sum_{t=1}^{T}\left(L_{t}(\widehat{\bm{y}}_{t})-\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]\right)+\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-S_{t}(\bm{U})\right).$ (6)
We first upper bound the first term in (6). Let $Z_{t}=L_{t}(\widehat{\bm{y}}_{t})-\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]$ for simplicity. Then, we have $Z_{t}\leq 1$, $\mathbb{E}_{t}\left[Z_{t}\right]=0$, and $\mathbb{E}_{t}\left[Z_{t}^{2}\right]\leq\mathbb{E}_{t}\left[(L_{t}(\widehat{\bm{y}}_{t}))^{2}\right]\leq(1-a)S_{t}(\bm{W}_{t})+q.$ Hence, from Bernstein’s inequality in Lemma D.1, for any $\delta^{\prime}\in(0,1)$, at least $1-\delta^{\prime}$ we have
$\sum_{t=1}^{T}Z_{t}\leq\sqrt{2\log(1/\delta^{\prime})\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})+q\right)}+\frac{\sqrt{2}}{3}\log(1/\delta^{\prime}).$ (7)
We next consider the second term in (6). Define $r_{t}=S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})$ for some $\xi\in(0,1)$, which will be determined later, and let $v_{t}=\mathds{1}[\bm{y}_{t}=\widehat{\bm{y}}_{t}]/p_{t}(\widehat{\bm{y}}_{t})\leq K/q$ for simplicity. Then, we have $\mathbb{E}_{t}[v_{t}r_{t}-r_{t}]=0$, $\lvert v_{t}r_{t}-r_{t}\rvert\leq KS_{\max}/q$, and
$\mathbb{E}_{t}[(v_{t}r_{t}-r_{t})^{2}]\leq\mathbb{E}_{t}[(v_{t}r_{t})^{2}]\leq\frac{KS_{\max}}{q}\lvert r_{t}\rvert\leq\frac{KS_{\max}}{q}\left(S_{t}(\bm{U})+S_{t}(\bm{W}_{t})\right).$
Hence from Bernstein’s inequality in Lemma D.1, for any $\delta^{\prime\prime}\in(0,1)$, with probability at least $1-\delta^{\prime\prime}$ we have
$\sum_{t=1}^{T}(v_{t}r_{t}-r_{t})\leq\sqrt{3\log(1/\delta^{\prime\prime})\sum_{t=1}^{T}\frac{KS_{\max}}{q}(S_{t}(\bm{U})+S_{t}(\bm{W}_{t}))}+\frac{\sqrt{2}KS_{\max}}{3q}\log(1/\delta^{\prime\prime}).$ (8)
When $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$. We first consider the case of $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$. From Lemma D.3, we have
$\sum_{t=1}^{T}\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-qT\leq\sum_{t=1}^{T}v_{t}S_{t}(\bm{U})+\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})-v_{t}S_{t}(\bm{W}_{t})\right)+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
$=\sum_{t=1}^{T}v_{t}\underbrace{\left(S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})\right)}_{=r_{t}}-(1-\xi)\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
$\leq\sum_{t=1}^{T}v_{t}r_{t}+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)},$
where the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. From the concentration result provided in (8), this is further bounded as
$\sum_{t=1}^{T}\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-qT$ $\leq\sum_{t=1}^{T}(S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t}))+\sqrt{3\log(1/\delta^{\prime\prime})\sum_{t=1}^{T}\frac{KS_{\max}}{q}(S_{t}(\bm{U})+S_{t}(\bm{W}_{t}))}$
$\qquad+\frac{\sqrt{2}KS_{\max}}{3q}\log(1/\delta^{\prime\prime})+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)},$
where we recall that $r_{t}=S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})$. Rearranging the last inequality and using $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$ give
$\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-S_{t}(\bm{U})\right)$ $\leq qT+\sqrt{6\log(1/\delta^{\prime\prime})\sum_{t=1}^{T}\frac{KS_{\max}}{q}S_{t}(\bm{W}_{t})}+\frac{\sqrt{2}KS_{\max}}{3q}\log(1/\delta^{\prime\prime})$
$\qquad+(1-a-\xi)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)}.$
In what follows, we let $\delta^{\prime}=\delta^{\prime\prime}=\delta/2$ and $\xi=\frac{(4+c)\gamma}{\lambda\nu}$ for a sufficiently small constant $c>0$, which satisfies $a+\xi>1$. Then, plugging (7) and the last inequality in (6), with probability at least $1-\delta$, we obtain
$\mathcal{R}_{T}$ $\leq\sqrt{2\log(2/\delta)\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})+q\right)}+\frac{\sqrt{2}}{3}\log(2/\delta)+qT+\sqrt{6\log(2/\delta)\sum_{t=1}^{T}\frac{KS_{\max}}{q}S_{t}(\bm{W}_{t})}$
$\qquad+\frac{\sqrt{2}KS_{\max}}{3q}\log(2/\delta)+(1-a-\xi)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)}$
$\leq\frac{1}{2(a+\xi-1)}\left((1-a)+\frac{3KS_{\max}}{q}\right)\log(2/\delta)+\sqrt{2qT\log(2/\delta)}+\frac{\sqrt{2}}{3}\log(2/\delta)+qT$
$\qquad+\frac{\sqrt{2}KS_{\max}}{3q}\log(2/\delta)+\frac{B^{2}b}{2q(1-\xi)}$
$\leq\frac{1}{q}\left(\frac{3KS_{\max}\log(2/\delta)}{2(a+\xi-1)}+KS_{\max}\log(2/\delta)+\frac{B^{2}Kb}{2(1-\xi)}\right)+qT+\sqrt{2qT\log(2/\delta)}$
$\qquad+\frac{1}{2(a+\xi-1)}(1-a)\log(2/\delta)+\frac{\sqrt{2}}{3}\log(2/\delta)$
$=\frac{\mathcal{C}}{q}+qT+\sqrt{2qT\log(2/\delta)}+\frac{1}{2(a+\xi-1)}(1-a)\log(2/\delta)+\frac{\sqrt{2}}{3}\log(2/\delta).$
Using the definition of $q=\sqrt{\mathcal{C}/T}$ with the last inequality, we obtain
$\mathcal{R}_{T}\leq 2\sqrt{\mathcal{C}T}+(\mathcal{C}T)^{1/4}\sqrt{\log(2/\delta)}+\left(\frac{1-a}{2(a+\xi-1)}+\frac{\sqrt{2}}{3}\right)\log(2/\delta).$
When $\sum_{t=1}^{T}S_{t}(\bm{U})>\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$. We next consider the case of $\sum_{t=1}^{T}S_{t}(\bm{U})>\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$. We have
$\mathcal{R}_{T}$ $=\sum_{t=1}^{T}\left(L_{t}(\widehat{\bm{y}}_{t})-\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]\right)+\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-S_{t}(\bm{U})\right)$
$\leq\sqrt{2\log(1/\delta^{\prime})\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})+q\right)}+\frac{\sqrt{2}}{3}\log(1/\delta^{\prime})+\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-S_{t}(\bm{W}_{t})\right)$
$\leq\sqrt{2\log(1/\delta^{\prime})\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})+q\right)}+\frac{\sqrt{2}}{3}\log(1/\delta^{\prime})+\sum_{t=1}^{T}\left(-aS_{t}(\bm{W}_{t})+q\right)$
$\leq\frac{(1-a)\log(1/\delta^{\prime})}{2a}+\sqrt{2qT\log(1/\delta^{\prime})}+\frac{\sqrt{2}}{3}\log(1/\delta^{\prime})+qT,$
where the first inequality follows from (7) and $\sum_{t=1}^{T}S_{t}(\bm{U})>\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$, and the second inequality follows from Assumption 3.2, the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Substituting $q=\sqrt{\mathcal{C}/2}$ and $\delta^{\prime}=\delta/2$ and in the last inequality, we obtain
$\mathcal{R}_{T}\leq\frac{(1-a)\log(2/\delta)}{2a}+\sqrt{2\log(2/\delta)}(\mathcal{C}T)^{1/4}+\frac{\sqrt{2}}{3}\log(2/\delta)+\sqrt{\mathcal{C}T}.$
This completes the proof. ∎
D.3 Proof of Theorem 3.8
Here, we provide the proof of Theorem 3.8. We recall that $\bm{P}_{t}=\mathbb{E}_{t}\left[{\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}}\right]$. We then estimate $\bm{y}_{t}$ by $\widetilde{\bm{y}}_{t}={\bm{V}}^{-1}\bm{P}^{+}_{t}\widehat{\bm{y}}_{t}\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle$ and $\bm{G}_{t}$ by $\widetilde{\bm{G}}_{t}\coloneqq(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}$ under Assumption 3.7. This $\widetilde{\bm{G}}_{t}$ satisfies $\mathbb{E}_{t}\left[{\widetilde{\bm{G}}_{t}}\right]=\bm{G}_{t}.$ To prove Theorem 3.8, we will upper bound $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$. To do so, we begin by proving the following lemma:
Lemma D.4.
Let $\bm{A}$ and $\bm{B}$ positive semi-definite matrices with $\operatorname*{Im}(\bm{A})=\operatorname*{Im}(\bm{B})$ with $\bm{A}\succeq\bm{B}$. Then, it holds that $\bm{A}^{+}\preceq\bm{B}^{+}$.
Proof.
Since $\operatorname*{Im}(\bm{A})=\operatorname*{Im}(\bm{B})$, there exists an orthogonal matrix $\bm{R}$, a diagonal matrix $\bm{\Lambda}$, and an invertible matrix $\bm{B}^{\prime}$ that has same dimensions as $\bm{\Lambda}$ such that
$\bm{A}=\bm{R}\begin{pmatrix}O&O\\ O&\bm{\Lambda}\end{pmatrix}\bm{R}^{\top}\quad\mbox{and}\quad\bm{B}=\bm{R}\begin{pmatrix}O&O\\ O&\bm{B}^{\prime}\end{pmatrix}\bm{R}^{\top}.$
Then,
$\bm{A}^{+}=\bm{R}\begin{pmatrix}O&O\\ O&\bm{\Lambda}^{-1}\end{pmatrix}\bm{R}^{\top}\quad\mbox{and}\quad\bm{B}^{+}=\bm{R}\begin{pmatrix}O&O\\ O&{\bm{B}^{\prime}}^{-1}\end{pmatrix}\bm{R}^{\top}.$ (9)
From $\bm{A}\succeq\bm{B}$, we have $\bm{\Lambda}\succeq\bm{B}^{\prime}$, which implies $\bm{\Lambda}^{-1}\preceq{\bm{B}^{\prime}}^{-1}$. From this and (9), we have $\bm{A}^{+}\preceq\bm{B}^{+}$, as desired. ∎
Using this lemma we prove a property of $\bm{P}_{t}$ and an upper bound of $\mathbb{E}_{t}\left[{\operatorname*{tr}\left(\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}\right)}\right]$. In what follows, we use $\lambda_{\min}(\bm{A})$ to denote the minimum eigenvalue of a matrix $\bm{A}$.
Lemma D.5.
Suppose that $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where we recall $\mu$ is the uniform distribution over $\mathcal{Y}$. Then, we have
$\mathbb{E}_{t}\left[{\operatorname*{tr}(\widetilde{\bm{y}}_{t}\widetilde{\bm{y}}_{t}^{\top})}\right]\leq\frac{\omega}{q}.$
Proof.
By the linearity of expectation and the trace property, we have
$\mathbb{E}_{t}\left[{\operatorname*{tr}(\widetilde{\bm{y}}_{t}\widetilde{\bm{y}}_{t}^{\top})}\right]$ $\leq\operatorname*{tr}\left({\bm{V}}^{-1}\bm{P}^{+}_{t}\mathbb{E}_{t}\left[{\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}}\right]\bm{P}^{+}_{t}\left({\bm{V}}^{-1}\right)^{\top}\right)=\operatorname*{tr}\left({\bm{V}}^{-1}\bm{P}^{+}_{t}\bm{P}_{t}\bm{P}^{+}_{t}\left({\bm{V}}^{-1}\right)^{\top}\right)$
$=\operatorname*{tr}\left({\bm{V}}^{-1}\bm{P}^{+}_{t}\left({\bm{V}}^{-1}\right)^{\top}\right),$
where the first inequality follows from $\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle=L_{t}(\widehat{\bm{y}}_{t})-\langle\widehat{\bm{y}}_{t},\bm{b}\rangle-c(\bm{y}_{t})\leq L_{t}(\widehat{\bm{y}}_{t})\leq 1$ since $\bm{b}\geq 0$ and $c(\cdot)$ is non-negative and the last equality follows from $\bm{P}^{+}_{t}\bm{P}_{t}\bm{P}^{+}_{t}=\bm{P}^{+}_{t}$. Hence,
$\operatorname*{tr}\left({\bm{V}}^{-1}\bm{P}^{+}_{t}\left({\bm{V}}^{-1}\right)^{\top}\right)$ $=\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}{\bm{V}}^{-1}\bm{P}^{+}_{t}\left({\bm{V}}^{-1}\right)^{\top}\bm{\mathrm{e}}_{i}\leq\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}{\bm{V}}^{-1}\left(q\bm{Q}\right)^{+}\left({\bm{V}}^{-1}\right)^{\top}\bm{\mathrm{e}}_{i}$
$\leq\operatorname*{tr}\left(\left({\bm{V}}^{-1}\right)^{\top}{\bm{V}}^{-1}(q\bm{Q})^{+}\right)=\frac{1}{q}\operatorname*{tr}\left({\bm{V}}^{-1}\bm{Q}^{+}({\bm{V}}^{-1})^{\top}\right)\leq\frac{\omega}{q},$
where in the first inequality we used Lemma D.4 and in the last inequality we used the assumption that $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)\leq\omega$. This completes the proof. ∎
Now, we are ready to upper bound $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$.
Lemma D.6.
Under the same assumption as Lemma D.5, it holds that
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]\leq 2bS_{t}(\bm{W}_{t})+\frac{2C_{x}^{2}\omega}{q}.$
Proof.
We have
$\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}$ $=\|\left(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t}\right)\bm{x}_{t}^{\top}\|_{\mathrm{F}}^{2}\leq 2\|(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}\|_{\mathrm{F}}^{2}+2\|(\bm{y}_{t}-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}\|_{\mathrm{F}}^{2}$
$\leq 2\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2C_{x}^{2}\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}^{2},$
where we recall $C_{x}=\mathrm{diam}(\mathcal{X})$. From this inequality,
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$ $\leq 2\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2C_{x}^{2}\mathbb{E}_{t}\left[{\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\leq 2bS_{t}(\bm{W}_{t})+2C_{x}^{2}\left(\|\bm{y}_{t}\|_{2}^{2}-2\bm{y}_{t}^{\top}\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]+\mathbb{E}_{t}\left[{\|\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\right)$
$=2bS_{t}(\bm{W}_{t})+2C_{x}^{2}\left(\|\bm{y}_{t}\|_{2}^{2}-2\|\bm{y}_{t}\|_{2}^{2}+\mathbb{E}_{t}\left[{\|\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\right)$
$\leq 2bS_{t}(\bm{W}_{t})+2C_{x}^{2}\mathbb{E}_{t}\left[{\operatorname*{tr}(\widetilde{\bm{y}}_{t}\widetilde{\bm{y}}_{t}^{\top})}\right]\leq 2bS_{t}(\bm{W}_{t})+\frac{2C_{x}^{2}\omega}{q},$
where in the second inequality we used $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$, in the equality we used $\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]=\bm{y}_{t}$, and in the last inequality we used Lemma D.5. ∎
Finally, we are ready to prove Theorem 3.8.
Proof of Theorem 3.8.
From Assumption 3.2, we have
$\mathbb{E}[\mathcal{R}_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\left\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\right\rangle}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT.$
From Lemma D.6 and the unbiasedness of $\widetilde{\bm{G}}_{t}$, the first term in the last inequality is further bounded as
$\mathbb{E}\left[\sum_{t=1}^{T}{\left\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\right\rangle}\right]$ $=\mathbb{E}\left[\sum_{t=1}^{T}{\left\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\right\rangle}\right]\leq\sqrt{2}B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}$
$\leq 2B\sqrt{b\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BC_{x}\sqrt{\omega/q},$
where the first inequality follows from Lemma 3.3 and the last inequality follows from Lemma D.6 and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, by combining with the last inequality, we have
$\mathbb{E}[\mathcal{R}_{T}]$ $\leq 2B\sqrt{b\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BC_{x}\sqrt{\omega/q}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$ (10)
$\leq\frac{bB^{2}}{a}+2BC_{x}\sqrt{\omega/q}+qT,$
where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Finally, substituting $q=\left(\frac{4\omega B^{2}C_{x}^{2}}{T}\right)^{1/3}$ in the last inequality gives
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2^{5/3}\omega^{1/3}\left(BC_{x}T\right)^{2/3},$
which is the desired bound. ∎
D.4 Proof of Corollary 3.9
Here, we derive the regret upper bounds provided by the algorithm established in Theorem 3.8 for online multiclass classification, online multilabel classification, and ranking. Recall that we can achieve
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(\omega^{1/3}\left(BC_{x}T\right)^{2/3}\right),$ (11)
where we recall that $\omega$ is defined as $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$. Note that when $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, then the matrix $\bm{Q}$ is invertible and $\lambda_{\min}(\bm{Q})>0$, and thus
$\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)=\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\leq\frac{1}{\lambda_{\min}(\bm{Q})}\sum_{i=1}^{d}\|(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\|_{2}^{2}\leq\frac{1}{\lambda_{\min}(\bm{Q})}\|\bm{V}^{-1}\|_{\mathrm{F}}^{2}.$ (12)
In each problem setting, this regret upper bound can be reduced to the following bounds:
Multiclass classification with 0-1 loss
We first consider multiclass classification with the 0-1 loss. Since $\bm{V}=\bm{1}\bm{1}^{\top}-\bm{I}$, we have $\|{\bm{V}}^{-1}\|_{\mathrm{F}}^{2}\leq d$ for $d\geq 2$. Recalling that $\mu$ is the uniform distribution over $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$, we have $\mathbb{E}_{\bm{y}\sim\mu}[(\bm{y}^{\top}\bm{x})^{2}]=\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}$ for any $\bm{x}\in\mathbb{R}^{d}$. Hence, $\lambda_{\min}(\bm{Q})=\min_{\|\bm{x}\|_{2}=1}\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]=\frac{1}{d}$, where the first equality is from Cesa-Bianchi & Lugosi (2012, Lemma 2). Since $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$ in this case, from (12) we can let $\omega=d/\lambda_{\min}(\bm{Q})=d^{2}$. Substituting these into our upper bound in (11), we obtain
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((BC_{x}dT)^{2/3}\right).$
Online multilabel classification with $m$ correct labels and the Hamming loss
We next consider online multilabel classification with the number of correct labels $m$ and the Hamming loss. Since $\bm{V}=-\frac{2}{d}\bm{I}$, we have $\|{\bm{V}}^{-1}\|_{\mathrm{F}}^{2}=\frac{d^{3}}{4}$. Let $\mathcal{Y}\subset\{0,1\}^{d}$ be the set of all vectors where exactly $m$ components are $1$, and the remaining components are all 00. By drawing $\bm{y}\in\mathcal{Y}$ according to the uniform distribution over $\mathcal{Y}$, the probability that a given component of $\bm{y}$ is $1$ is $\binom{m-1}{d-1}/\binom{m}{d}=\frac{m}{d}$. Hence, for any $\bm{x}\in\mathbb{R}^{d}$ with $\|\bm{x}\|_{2}=1$, we have
$\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]=\frac{m}{d}\sum_{i=1}^{d}x_{i}^{2}+\frac{m^{2}}{d^{2}}\sum_{i\neq j}x_{i}x_{j}=\left(\frac{m}{d}\sum_{i=1}^{d}x_{i}\right)^{2}+\frac{m(d-m)}{d^{2}}\|\bm{x}\|_{2}^{2}\geq\frac{m(d-m)}{d^{2}}.$
Hence, we have $\lambda_{\min}(\bm{Q})=\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]\geq\frac{m(d-m)}{d^{2}}$, where the equality is from Cesa-Bianchi & Lugosi (2012, Lemma 2). Since $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, from (12) we can choose $\omega=d^{3}/(4\lambda_{\min}(\bm{Q}))=\frac{d^{5}}{4m(d-m)}$. Therefore, our regret upper bound in (11) is reduced to
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(\left(\frac{B^{2}C_{x}^{2}d^{5}}{m(d-m)}\right)^{1/3}T^{2/3}\right).$
Ranking with the Hamming loss and the number of items $m$
We finally consider online ranking with the Hamming loss and the number of items $m$. From Cesa-Bianchi & Lugosi (2012, Proposition 4), the smallest positive eigenvalue is at least $1/m$. Hence, since $\bm{V}=-\frac{1}{m}\bm{I}$, we have
$\operatorname*{tr}({\bm{V}}^{-1}\bm{Q}^{+}({\bm{V}}^{-1})^{\top})=m^{2}\operatorname*{tr}(\bm{Q}^{+})\leq m^{2}\sum_{i=1}^{\mathrm{rank}(\bm{Q}^{+})}m\leq m^{5},$
where we used $\mathrm{rank}(\bm{Q}^{+})\leq d=m^{2}$, and this allows us to choose $\omega=m^{5}$. Substituting these values into our regret upper bound in (11) , we obtain
$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(m^{5/3}(BC_{x}T)^{2/3}\right).$
Appendix E Omitted Details of Section 4
This section provides the proofs of the theorems in Section 4.
E.1 Details of Optimistic Delayed Adaptive FTRL (ODAFTRL)
We provide a more detailed explanation of the Optimistic Delayed Adaptive FTRL (ODAFTRL) algorithm used for updating $\bm{W}_{t}$ in Section 4. We recall that ODAFTRL computes $\bm{W}_{t}$ by the following update rule:
$\bm{W}_{t+1}=\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}\left\{\sum_{i=1}^{t-D}\langle\bm{G}_{i},\bm{W}\rangle+\frac{\lambda_{t}\|\bm{W}\|_{\mathrm{F}}^{2}}{2}\right\},$ (13)
where $\lambda_{t}\geq 0$ is a regularization parameter. The ODAFTRL algorithm, when using this parameter update called AdaHedgeD, satisfies the following lemma:
Lemma E.1 (Flaspohler et al. 2021, Theorem 12).
Fix $\alpha>0$. Let $S_{t}:\mathcal{W}\to\mathbb{R}$ be a convex function for each $t=1,\dots,T$. Suppose that we update $\lambda_{t}$ in (13) by the following AdaHedgeD update:
$\lambda_{t+1}$ $=\frac{1}{\alpha}\sum_{s=1}^{t-D}\delta_{s},$
$\delta_{t}$ $=\min\left\{F_{t+1}(\bm{W}_{t})-F_{t+1}(\bar{\bm{W}}_{t}),\langle\bm{G}_{t},\bm{W}_{t}-\bar{\bm{W}}_{t}\rangle,F_{t+1}(\widehat{\bm{W}}_{t})-F_{t+1}(\bar{\bm{W}}_{t})+\langle\bm{G}_{t},\bm{W}_{t}-\widehat{\bm{W}}_{t}\rangle\right\}_{+},$
$\bar{\bm{W}}_{t}$ $=\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}F_{t+1}(\bm{W}),$
$\widehat{\bm{W}}_{t}$ $=\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}\left\{F_{t+1}(\bm{W})-\min\left\{\frac{\|\bm{G}_{t}\|_{\mathrm{F}}}{\|\bm{G}_{t-D:t}\|_{\mathrm{F}}},1\right\}\langle\bm{G}_{t-D:t},\bm{W}\rangle\right\},\text{ and}$
$F_{t+1}(\bm{W})$ $=\frac{\lambda_{t}\|\bm{W}\|_{\mathrm{F}}^{2}}{2}+\langle\bm{G}_{1:t},\bm{W}\rangle.$
Then, for any $\bm{U}\in\mathcal{W}$, ODAFTRL achieves
$\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\leq\sum_{t=1}^{T}{S_{t}(\bm{U})}+\left(\frac{B^{2}}{2\alpha}+1\right)\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+2\alpha b_{t}}\right),$
where
$a_{t}$ $=B\min\{\|\bm{G}_{t-D:t}\|_{\mathrm{F}},\|\bm{G}_{t}\|_{\mathrm{F}}\},$
$b_{t}$ $=\operatorname{huber}(\|\bm{G}_{t-D:t}\|_{\mathrm{F}},\|\bm{G}_{t}\|_{\mathrm{F}}),\>\text{and}\>\operatorname{huber}(x,y)=\frac{1}{2}x^{2}-\frac{1}{2}(|x|-|y|)^{2}_{+}\leq\min\left\{\frac{1}{2}x^{2},|x||y|\right\}.$
In the following, we let $\alpha=\max_{\bm{W}\in\mathcal{W}}\frac{\|\bm{W}\|_{\mathrm{F}}^{2}}{2}=\frac{B^{2}}{2}$ for simplicity. Then,
$\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\leq\sum_{t=1}^{T}{S_{t}(\bm{U})}+2\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+B^{2}b_{t}}\right),$
E.2 Proof of Theorem 4.3
We present Theorem 4.3 in a more detailed form and provide its proof.
Theorem E.2 (Formal version of Theorem 4.3).
Let $\alpha=\frac{B^{2}}{2}$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a delay of $D$ achieves
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BDL+\frac{2bB^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3}.$
Proof.
From Lemma E.1 and the definition of $a_{t}$ and $b_{t}$, we have
$\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}$ $\leq 2\left(2B\max_{s\in[T]}\sum_{i=s-D}^{s-1}\|\bm{G}_{i}\|_{\mathrm{F}}+\sqrt{\sum_{t=1}^{T}{\left(B^{2}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+B^{2}\|\bm{G}_{t-D:t}\|_{\mathrm{F}}\|\bm{G}_{t}\|_{\mathrm{F}}\right)}}\right)$
$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right),$
where we used the Cauchy–Schwarz inequality, $\|\bm{G}_{t}\|_{\mathrm{F}}\leq L$, $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$, and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$ in the last inequality. From this inequality and Assumption 4.1, we can evaluate surrogate regret as
$\mathbb{E}\left[\mathcal{R}_{T}\right]$ $\leq\sum_{t=1}^{T}{\left((1-a)S_{t}(\bm{W}_{t})-S_{t}(\bm{U})\right)}$
$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right)-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
$\leq 4BDL+\frac{2bB^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3}.$
We used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq({3}/{4})\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$ that hold for any $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$ in the last inequality. ∎
E.3 Proof of Theorem 4.4
We present Theorem 4.4 in a more detailed form and provide its proof.
Theorem E.3.
Let $\alpha=\frac{B^{2}}{2}$ and $\delta\in(0,1)$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a delay of $D$ achieves
$\mathcal{R}_{T}\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3},$
with probability at least $1-\delta$.
Proof.
We decompose $\mathcal{R}_{T}$ into
$\mathcal{R}_{T}=\sum_{t=1}^{T}{\left(L_{t}(\widehat{\bm{y}}_{t})-\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\right)}+\sum_{t=1}^{T}{(\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]-S_{t}(\bm{U}))}.$ (14)
Let $Z_{t}=L_{t}(\widehat{\bm{y}}_{t})-\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]$. Then, we have $|Z_{t}|\leq 1$ and $\mathbb{E}_{t}\left[{Z_{t}^{2}}\right]\leq(1-a)S_{t}(\bm{W}_{t})$ from Assumption 4.1. Hence, from Lemma D.1, with probability at least $1-\delta$, the first term in (14) is upper bounded as
$\sum_{t=1}^{T}{Z_{t}}\leq\sqrt{2(1-a)\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\log\frac{1}{\delta}}+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}.$ (15)
From from Assumption 4.1 and Lemma E.1, the second term in (14) is also upper bounded as
$\sum_{t=1}^{T}{(\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]-S_{t}(\bm{U}))}$
$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right)-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})},$ (16)
where we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, substituting (15) and (16) into (14) gives
$\mathcal{R}_{T}$ $\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\left(\sqrt{2(1-a)\log\frac{1}{\delta}}+2\sqrt{bB^{2}}\right)\sqrt{\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
$\quad+2\left(bB^{4}(D+1)^{2}L^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
$\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3},$
where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq({3}/{4})\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$ in the last inequality. This is the desired bound. ∎
Appendix F Omitted Details of Section 5
This section provides the omitted proofs of the theorems in Section 5.
F.1 Common analysis
We provide the analysis that is commonly used in the proofs of Theorems 5.1 and 5.2. We use ODAFTRL with the AdaHedgeD update in Section E.1 for $\mathsf{ALG}$. From Lemma E.1, it holds that
$\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq 2\left(2\max_{t\in[T]}a_{t-D:t-1}+\sqrt{\sum_{t=1}^{T}{a_{t}^{2}}+B^{2}b_{t}}\right),$ (17)
where
$a_{t}=B\min\{\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}},\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\}\quad\mbox{and}\quad b_{t}\leq\min\left\{\frac{1}{2}\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}^{2},\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right\}.$
By the definition of $a_{t}$, we have
$\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]\leq B\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]\leq B\mathbb{E}\left[\sqrt{D\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}^{2}}\right]\leq B\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]},$ (18)
from the Cauchy–Schwarz inequality and Jensen’s inequality. Hence, from (17) and (18), we have
$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2\left(2\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{a_{t}^{2}}\right]}+B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
$\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right),$ (19)
where we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. The last term in the last inequality is further bounded as
$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
$=\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right],$ (20)
where the second inequality follows from the triangle inequality and the equality follows from the law of total expectation.
F.2 Proof of Theorem 5.1
We provide the complete version of Theorem 5.1:
Theorem F.1 (Formal version of Theorem 5.1).
The algorithm in Section 5.1 achieves
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq\frac{4bB}{a}\left(D^{1/4}+D^{-1/4}\right)^{2}\sqrt{KT}+2BL\sqrt{DT}+B\sqrt{DKT}=O(D\sqrt{KT}).$
Proof.
First, we will upper bound $\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$. From (F.1), we have
$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+L\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DL^{2}T,$ (21)
where the second and third inequality follow from the inequality $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]=\|\bm{G}_{t}\|_{\mathrm{F}}\leq L$. Hence, from (F.1) and (F.2), it holds that
$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DL^{2}T}}\right)$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2BL\sqrt{DT}$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BL\sqrt{DT},$ (22)
where in the second inequality we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$ and in the last inequality we used
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}).$
Therefore, combining all the above arguments yields
$\mathbb{E}\left[\mathcal{R}_{T}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BL\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq\frac{4bB^{2}}{a}\left(\sqrt{D}+1\right)^{2}\frac{bK}{q}+2BL\sqrt{DT}+qT,$
where the first inequality follows from Assumption 3.2, the second inequality follows from the convexity of $S_{t}$ and the unbiasedness of $\widetilde{\bm{G}}_{t}$, and the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Finally, by substituting $q=B\sqrt{DK/T}$, we obtain
$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq\frac{4bB}{a}\left(D^{1/4}+D^{-1/4}\right)^{2}\sqrt{KT}+2BL\sqrt{DT}+B\sqrt{DKT},$
which is the desired bound. ∎
F.3 Proof of Theorem 5.2
We provide the complete version of Theorem 5.2:
Theorem F.2 (Formal version of Theorem 5.2).
The algorithm in Section 5.2 achieves
$\mathbb{E}[\mathcal{R}_{T}]=\frac{8bB^{2}}{a}\Big{(}\sqrt{D}+1\Big{)}^{2}+2B(L+C_{x}C_{y})\sqrt{DT}+O\left(B^{2/3}D^{1/3}T^{2/3}\right).$
Proof.
First, we will derive an upper bound of $\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$. We first observe that
$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]$ $=\mathbb{E}_{t}\left[{\|(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}\|_{\mathrm{F}}}\right]\leq\mathbb{E}_{t}\left[{\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}}\right]\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\mathbb{E}_{t}\left[{\|\bm{y}_{t}\|_{2}+\|\widetilde{\bm{y}}_{t}\|_{2}}\right]$
$\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+C_{x}\mathbb{E}_{t}\left[{\sqrt{\operatorname*{tr}\left(\widetilde{\bm{y}}_{t}\widetilde{\bm{y}}_{t}^{\top}\right)}}\right]\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+C_{x}{\sqrt{\mathbb{E}_{t}\left[{\operatorname*{tr}\left(\widetilde{\bm{y}}_{t}\widetilde{\bm{y}}_{t}^{\top}\right)}\right]}}$
$\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\leq L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q},$
where the first inequality follows from $C_{x}=\mathrm{diam}(\mathcal{X})$, the third inequality follows from $C_{y}=\mathrm{diam}(\mathcal{Y})$, the fourth inequality follows from Jensen’s inequality, and the fifth inequality follows from Lemma D.5. Thus, combining (F.1) with the last inequality, we have
$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\left(L+C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DT\left(L+C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)^{2}.$ (23)
Hence, from (F.1) and (23), we have
$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$
$\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT}$
$\leq 4B\left(\sqrt{D}+1\right)\sqrt{2\sum_{t=1}^{T}{\left(bS_{t}(\bm{W}_{t})+\frac{C_{x}^{2}\omega}{q}\right)}}+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT},$ (24)
where the first inequality follows from (F.1), the second inequality follows from (23) and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$, and the last inequality follows from Lemma D.6. Therefore, combining all the above arguments yields
$\mathbb{E}\left[\mathcal{R}_{T}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq 4B\left(\sqrt{D}+1\right)\left(\sqrt{2b\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\sqrt{2C_{x}^{2}\omega T/q}\right)$
$\qquad+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
$\leq\frac{8bB^{2}}{a}\Big{(}\sqrt{D}+1\Big{)}^{2}+2B(L+C_{x}C_{y})\sqrt{DT}+2BC_{x}\big{(}(2\sqrt{2}+1)\sqrt{D}+2\sqrt{2}\big{)}\sqrt{\omega T/q}+qT,$
where the first inequality follows from Assumption 3.2, the third inequality follows from (F.3) and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$, and the last inequality follows from the definition of $\varepsilon$ and $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$. Finally, substituting $q=\left(\frac{\omega B^{2}C_{x}^{2}D}{T}\right)^{1/3}$ gives the desired bound. ∎
Appendix G Additional Experiments
(a) $n^{\prime}=2,p=0.0$
(b) $n^{\prime}=2,p=0.1$
(c) $n^{\prime}=4,p=0.0$
(d) $n^{\prime}=4,p=0.1$
Figure 2: Results of the synthetic experiments in multiclass classification with bandit feedback. In all figures, the horizontal axis represents the number of classes $K$ and the vertical axis represents the cumulative prediction loss.
This section provides the results of numerical experiments for online multiclass classification with bandit feedback on synthetic data, which could not be included in the main text due to space constraints.
G.1 Setup
Here, we describe the experimental setup. As Section 6, we compare three algorithms: Gaptron (Van der Hoeven, 2020) with logistic loss, Gappletron (Van der Hoeven et al., 2021) with hinge loss, and our algorithm in Section 5.1. All experiments were run on a system with 16GB of RAM, Apple M3 CPU, and in Python 3.11.7 on a macOS Sonoma 14.6.1.
Data generation
We describe the procedure for generating synthetic data. The synthetic data was generated by using the same method as Van der Hoeven et al. (2021). The input vector consists of a binary vector with entries of 00 and $1$, and is composed of two parts. The first part corresponds to a unique feature vector associated with the label, and the second part is randomly selected and unrelated to the label. Specifically, the data are generated as follows: Let $n^{\prime}$, $K\in\mathbb{N}$, $r\in[0,1]$, and $T>0$. We generate $K$ unique feature vectors of length $10n^{\prime}$ using the following as follows. We first randomly select an integer $s$ uniformly from the range $n^{\prime}$ to $5n^{\prime}$. After that, we randomly choose $s$ elements from a zero vector of length $10n^{\prime}$ and set them to 1. The input vector is obtained by concatenating the feature vector of a randomly chosen class with a vector of length $30n^{\prime}$, in which exactly $5n^{\prime}$ elements are randomly set to $1$. Additionally, as noise, with probability $r$, the corresponding class is replaced with a randomly selected class instead of the original one. The length of the generated input vector $n$ is $40n^{\prime}$. These input vectors are generated for $T$. Based on the above procedure, we create data for $n^{\prime}\in\{2,4\}$ and $r\in\{0.0,0.1\}$.
Details of algorithms
As the algorithm $\mathsf{ALG}$ to update the linear estimator, we employ the adaptive OGD in Section 3.2. Following Van der Hoeven et al. (2021), we use the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\left(10^{-8}+\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}\right)}}$ and no orthogonal projection in conducted in OGD. Since $B=\mathrm{diam}(\mathcal{W})$ is unknown, we fixed $B=10$ for the experiments, regardless of whether this value represents the actual diameter. All other parameters were set according to theoretical values. Under these parameter settings, we repeated experiments for $20$ times.
G.2 Results
The results on synthetic data are provided in Figure 2. Our algorithm achieves comparable or better results than the comparison methods for datasets when $K\leq 24$. By contrast, when $K=48,96$, the cumulative losses of our algorithm are larger compared to the other two algorithms, Gaptron with the hinge loss and Gappletron with the logistic loss. Note that these experimental results are consistent with existing theoretical results since for larger values of $K$ the upper bound on the cumulative loss of Gappletron is better than our algorithm due to differences in the surrogate loss function (see the discussion in Appendix C for details). Nevertheless, in this experiment, the advantages of structural prediction algorithms are not fully exploited, making the setup more favorable for the algorithms in previous studies. It is also worth noting that by using the same decoding function as theirs, our approach can achieve the same order of the cumulative losses in online multiclass classification with bandit feedback.

### v2

@@
 Online structured prediction is a task of sequentially predicting outputs with complex structures based on inputs and past observations, encompassing online classification.
-Recent studies showed that in the full information setup, we can achieve finite bounds on the surrogate regret, i.e., the extra target loss relative to the best possible surrogate loss.
-In practice, however, full information feedback is often unrealistic as it requires immediate access to the whole structure of complex outputs.
+Recent studies showed that in the full-information setting, we can achieve finite bounds on the surrogate regret, i.e., the extra target loss relative to the best possible surrogate loss.
+In practice, however, full-information feedback is often unrealistic as it requires immediate access to the whole structure of complex outputs.
 Motivated by this, we propose algorithms that work with less demanding feedback, bandit and delayed feedback.
-For the bandit setting, using a standard inverse-weighted gradient estimator, we achieve a surrogate regret bound of $O(\sqrt{KT})$ for the time horizon $T$ and the size of the output set $K$.
-However, $K$ can be extremely large when outputs are highly complex, making this result less desirable.
-To address this, we propose an algorithm that achieves a surrogate regret bound of $O(T^{2/3})$, which is independent of $K$.
-This is enabled with a carefully designed pseudo-inverse matrix estimator.
-Furthermore, for the delayed full information feedback setup, we obtain a surrogate regret bound of $O(D^{2/3}T^{1/3})$ for the delay time $D$.
-We also provide algorithms for the delayed bandit feedback setup.
-Finally, we numerically evaluate the performance of the proposed algorithms in online classification with bandit feedback.
+For bandit feedback, by using a standard inverse-weighted gradient estimator, we achieve a surrogate regret bound of $O(\sqrt{KT})$ for the time horizon $T$ and the size of the output set $K$.
+However, $K$ can be extremely large when outputs are highly complex, resulting in an undesirable bound.
+To address this issue, we propose another algorithm that achieves a surrogate regret bound of $O(T^{2/3})$, which is independent of $K$.
+This is achieved with a carefully designed pseudo-inverse matrix estimator.
+Furthermore, we numerically compare the performance of these algorithms, as well as existing ones.
+Regarding delayed feedback, we provide algorithms and regret analyses that cover various scenarios, including full-information and bandit feedback, as well as fixed and variable delays.
 1 Introduction
-In many machine learning problems, given an input vector from a vector space $\mathcal{X}$, we aim to predict a vector in a finite output space $\mathcal{Y}$.
-Multiclass classification is one of the simplest examples, while in other cases output spaces may have more complex structures.
-Structured prediction refers to such a class of problems with structured output spaces, including multiclass classification, multilabel classification, ranking, and ordinal regression, and it has applications in various fields, ranging from natural language processing to bioinformatics (Tsochantaridis et al., 2005; BakIr et al., 2007).
+In many machine learning problems, given an input vector from a set $\mathcal{X}$ of input vectors, we aim to predict a vector in a finite output space $\mathcal{Y}$.
+Multiclass classification is one of the simplest examples, while in other cases, output spaces may have more complex structures.
+Structured prediction refers to such a class of problems with structured output spaces, including multiclass classification, multilabel classification, ranking, and ordinal regression, and it has applications in various fields, ranging from natural language processing to bioinformatics [37, 2].
 In structured prediction, training models that directly predict outputs in complex discrete output spaces is typically challenging.
-Therefore, we often adopt the surrogate loss framework (Bartlett et al., 2006)—define an intermediate space of score vectors and train models that estimate score vectors from inputs based on surrogate loss functions.
-Examples of surrogate losses include squared, logistic, and hinge losses, and a general framework encompassing them is the Fenchel–Young loss (Blondel et al., 2020), which we rely on in this study.
-Structured prediction can be naturally extended to the online setting (Sakaue et al., 2024).
-In this online setting, at each round $t=1,\dots,T$, the environment selects an input-output pair $(\bm{x}_{t},\bm{y}_{t})\in\mathcal{X}\times\mathcal{Y}$.
-A learner then predicts $\widehat{\bm{y}}_{t}\in\mathcal{Y}$ based on the input $\bm{x}_{t}$ and incurs a loss $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$, where $L:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is the target loss function.
-As with Sakaue et al.
-(2024), we focus on the simple yet fundamental case where the learner’s model for estimating score vectors is linear.
-Table 1: Surrogate regret upper and lower bounds in online multiclass classification and online structured prediction.
-Here, $T$ is the time horizon, $K$ is the size of the set, $\mathcal{Y}$, of output vectors, and $D$ is the fixed delayed time.
-“FI” is the abbreviation of full information.
+Therefore, we often adopt the surrogate loss framework [3]—define an intermediate space of score vectors and train models that estimate score vectors from inputs based on surrogate loss functions.
+Examples of surrogate losses include squared, logistic, and hinge losses, and a general framework encompassing them is the Fenchel–Young loss [6], which we rely on in this study.
+Structured prediction can be naturally extended to the online setting, called Online Structured Prediction (OSP) [35].
+In OSP, at each round $t=1,\dots,T$, an environment selects an input–output pair $(\bm{x}_{t},\bm{y}_{t})\in\mathcal{X}\times\mathcal{Y}$.
+A learner then predicts $\widehat{\bm{y}}_{t}\in\mathcal{Y}$ based on the input $\bm{x}_{t}$ and incurs a loss $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$, where $L\colon\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is the target loss function.
+Following prior work [38, 39, 35], we focus on the simple yet fundamental case where the learner’s model for estimating score vectors is linear.
+Table 1: Upper and lower bounds on the surrogate regret in online multiclass classification and OSP.
+Here, $T$ is the time horizon, $K=\lvert\mathcal{Y}\rvert$ is the size of the output space, and $D$ is the fixed delay time.
 Delayed feedback is considered only when “Delayed” appears in the feedback column.
@@
 Note that the $O(T^{2/3})$ bounds for SELF* in lines 6 and 9 do not explicitly depend on $K$ but on $d$; in the case of multiclass classification with the 0-1 loss, the dependence on $K$ appears as $d=K$.
-Problem setup Feedback Target loss Surrogate regret bound
-Van der Hoeven et al. (2021, Cor. 1) Binary classification Bandit 0-1 loss $\Omega(\sqrt{T})$ ($d=2$)
-Van der Hoeven (2020, Thm. 4) Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
-Van der Hoeven et al. (2021, Thm. 1) Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
-Sakaue et al. (2024, Thms. 7 and 8) Structured prediction FI SELF $O(1)$
-This work (Theorems 3.5 and 3.6) Structured prediction Bandit SELF $O(\sqrt{KT})$
+Problem setting Feedback Target loss Surrogate regret bound
+van der Hoeven et al. [39, Cor. 1] Binary classification Graph bandit 0-1 loss $\Omega(\sqrt{T})$ ($d=2$)
+van der Hoeven [38, Thm. 4] Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
+van der Hoeven et al. [39, Thm. 1] Multiclass classification Bandit 0-1 loss $O(K\sqrt{T})$
+Sakaue et al. [35, Thms. 7 and 8] Structured prediction Full-info SELF $O(1)$
+This work (Theorems 3.5 and D.2) Structured prediction Bandit SELF $O(\sqrt{KT})$
 This work (Theorem 3.7) Structured prediction Bandit SELF* $O(T^{2/3})$
-This work (Theorems 4.3 and 4.4) Structured prediction FI & Delayed SELF $O(D^{2/3}T^{1/3})$
-This work (Theorem 5.1) Structured prediction Bandit & Delayed SELF $O(\sqrt{DKT})$
+This work (Theorems 4.3 and E.3) Structured prediction Full-info & Delayed SELF $O(\min\{D^{2},D^{2/3}T^{1/3}\})$
+This work (Theorem 5.1) Structured prediction Bandit & Delayed SELF $O(\sqrt{(K+D)T})$
 This work (Theorem 5.2) Structured prediction Bandit & Delayed SELF* $O(D^{1/3}T^{2/3})$
 The goal of the learner is to minimize the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$.
-On the other hand, the best the learner can do in the surrogate loss framework is to minimize the cumulative surrogate loss, namely $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$, where $\bm{U}:\mathcal{X}\to\mathbb{R}^{d}$ is the best offline linear estimator and $S:\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a surrogate loss, which measures the discrepancy between the score vector $\bm{U}\bm{x}_{t}\in\mathbb{R}^{d}$ and $\bm{y}_{t}\in\mathcal{Y}$.
-Given this, a natural performance measure of the learner’s predictions is the following surrogate regret $\mathcal{R}_{T}$:
-$\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}.$ (1)
-The surrogate regret was introduced in the seminal paper by Van der Hoeven (2020) in the context of online multiclass classification.
+On the other hand, the best the learner can do in the surrogate loss framework is to minimize the cumulative surrogate loss, namely $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$, where $\bm{U}\colon\mathcal{X}\to\mathbb{R}^{d}$ is the best offline linear estimator and $S\colon\mathbb{R}^{d}\times\mathcal{Y}\to\mathbb{R}_{\geq 0}$ is a surrogate loss, which measures the discrepancy between the score vector $\bm{U}\bm{x}_{t}\in\mathbb{R}^{d}$ and $\bm{y}_{t}\in\mathcal{Y}$.
+Given this, a natural performance measure of the learner’s predictions is the surrogate regret, $\mathcal{R}_{T}$, such that $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}$.
+The surrogate regret was introduced in the seminal paper by van der Hoeven [38] in the context of online multiclass classification.
 Recently, Sakaue et al.
-(2024) showed that a finite surrogate regret bound can be achieved for online structured prediction under full information feedback, i.e., the learner can observe $\bm{y}_{t}$ at the end of each round $t$.
-However, the assumption that full information feedback is available is often demanding, especially when outputs have complex structures.
-For example, in sequential ad assortment on an advertising platform, we may be able to observe only the click-through rate but not which ads were clicked, which boils down to the bandit feedback setting (Kakade et al., 2008; Gentile & Orabona, 2014a).
-Also, we may only have access to feedback from a while ago when designing an ad assortment for a new user—namely, delayed feedback (Weinberger & Ordentlich, 2002; Manwani & Agarwal, 2023).
+[35] showed that a finite surrogate regret bound can be achieved for OSP under full-information feedback, i.e., the learner can observe $\bm{y}_{t}$ at the end of each round $t$.
+However, the assumption that full-information feedback is available is often demanding, especially when outputs have complex structures.
+For example, in sequential ad assortment on an advertising platform, we may be able to observe only the click-through rate but not which ads were clicked, which boils down to the bandit feedback setting [24, 19].
+Also, we may only have access to feedback from a while ago when designing an ad assortment for a new user—namely, delayed feedback [41, 27].
 Similar situations have led to a plethora of studies in various online learning settings.
-In combinatorial bandits, algorithms under bandit feedback (referred to as full-bandit feedback in their context), instead of full information feedback, have been widely studied (Cesa-Bianchi & Lugosi, 2012; Combes et al., 2015; Rejwan & Mansour, 2020; Du et al., 2021).
-Delayed feedback is also explored in various settings, including full information and bandit feedback (Joulani et al., 2013; Cesa-Bianchi et al., 2016).
+In combinatorial bandits, algorithms under bandit feedback (referred to as full-bandit feedback in their context), instead of full-information feedback, have been widely studied [8, 12, 33, 15].
+Delayed feedback is also explored in various other settings, including full-information and bandit feedback [22, 9].
 Due to space limitations, we defer a further discussion of the background to Appendix B.
 Our contributions
-To extend the applicability of online structured prediction, this study develops online structured prediction algorithms that can handle weaker feedback—bandit feedback and delayed feedback—instead of full information feedback.
-As with Sakaue et al.
-(2024), we consider the case where target loss functions belong to a class called the Structured Encoding Loss Function (SELF) (Ciliberto et al., 2016; Blondel, 2019), a general class including the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking (see Section 2.4 for the definition).
-Table 1 summarizes the surrogate regret bounds provided in this study and comparisons with existing results.
-One of the challenges of bandit feedback is that the true output $\bm{y}_{t}$ is not observed, making it impossible to compute the true gradient of the surrogate loss.
-To deal with this, we use an inverse-weighted gradient estimator, a typical approach that assigns weights to gradients by the inverse of choosing each output, establishing an $O(\sqrt{KT})$ surrogate regret upper bounds (Theorems 3.5 and 3.6), where $K=\lvert\mathcal{Y}\rvert$ is the cardinality of $\mathcal{Y}$.
-The $O(\sqrt{KT})$ bound has an optimal dependence on $T$; it matches the $\sqrt{T}$ lower bound in the special case of online multiclass classification with bandit feedback (Van der Hoeven et al., 2021, Corollary 1).
-Furthermore, our bound is better than the existing $O(K\sqrt{T})$ bound of Van der Hoeven (2020) by a factor of $\sqrt{K}$, while it is not directly comparable to the latest $O(K\sqrt{T})$ bound in Van der Hoeven et al.
-(2021) due to differences in surrogate loss functions.
-See Appendix C for a more detailed discussion.
-While the $O(\sqrt{KT})$ bound is satisfactory when $K=\lvert\mathcal{Y}\rvert$ is small, $K$ can be extremely large in some structured prediction problems: in multilabel classification with $m$ correct labels, we have $K=\binom{d}{m}$, and in ranking problems with $m$ items, we have $K=m!$.
+To extend the applicability of OSP, this study develops OSP algorithms that can handle weaker feedback—bandit feedback and delayed feedback—instead of full-information feedback.
+Following the work of Sakaue et al.
+[35], we consider the case where target loss functions belong to a class called the Structured Encoding Loss Function (SELF) [10, 5], a general class that includes the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking (see Section 2.4 for the definition).
+Table 1 summarizes the surrogate regret bounds provided in this study and comparisons with the existing results.
+One of the major challenges in the bandit feedback setting is that the true output $\bm{y}_{t}$ is not observed, making it impossible to compute the true gradient of the surrogate loss.
+To address this, we use an inverse-weighted gradient estimator, a common approach that assigns weights to gradients by the inverse probability of choosing each output, establishing an $O(\sqrt{KT})$ surrogate regret upper bound (Theorems 3.5 and D.2), where $K=\lvert\mathcal{Y}\rvert$ denotes the cardinality of $\mathcal{Y}$.
+This $O(\sqrt{KT})$ bound has a desirable dependence on $T$, matching an $\Omega(\sqrt{T})$ lower bound known in a problem closely related to online multiclass classification with bandit feedback [39, Corollary 1].
+Furthermore, our bound is better than the existing $O(K\sqrt{T})$ bounds [38, 39] by a factor of $\sqrt{K}$, although the bound of van der Hoeven et al.
+[39] applies to a broader class of surrogate losses and thus it is not directly comparable to ours (see Appendix C for a more detailed discussion).
+We also conduct numerical experiments on online multiclass classification and find that our methods, which apply to general OSP, are comparable to the existing ones specialized for multiclass classification (see Section H.1).
+While the $O(\sqrt{KT})$ bound is satisfactory when $K=\lvert\mathcal{Y}\rvert$ is small, $K$ can be extremely large in some structured prediction problems.
+In multilabel classification with $m$ correct labels, we have $K=\binom{d}{m}$, and in ranking problems with $m$ items, we have $K=m!$.
 To address this issue, we consider a special case of SELF (denoted by SELF* in Table 1), which still includes the aforementioned examples: the 0-1 loss in multiclass classification and the Hamming loss in multilabel classification and ranking.
 A technical challenge to resolve the issue lies in designing an appropriate gradient estimator used in online learning methods.
-To this end, we draw inspiration from pseudo-inverse estimators used in the adversarial linear/combinatorial bandit literature (Dani et al., 2007; Abernethy et al., 2008; Cesa-Bianchi & Lugosi, 2012).
-While we cannot naively apply the existing estimators, we design a new gradient estimator that applies to various specific structured prediction problems belonging to the special SELF framework.
-Armed with this gradient estimator, we achieve a surrogate regret upper bound of $O(T^{2/3})$, which does not explicitly depend on $K$ (Theorem 3.8).
-For the delayed feedback setting with a known fixed delay time of $D$, it is actually not difficult to obtain a surrogate regret bound of $O(\sqrt{DT})$ with standard Online Convex Optimization (OCO) algorithms for delayed feedback.
-Our finding is that we can achieve a surrogate regret bound of $O(D^{2/3}T^{1/3})$ in online structured prediction under delayed full information feedback (Theorems 4.3 and 4.4) by leveraging ODAFTRL (Flaspohler et al., 2021), a Follow-the-Regularized-Leader-type algorithm that achieves an AdaGrad-type regret upper bound in OCO under delayed feedback.
-This bound is better than $O(\sqrt{DT})$ as $D\leq T$.
-Given the contributions so far, it is natural to explore online structured prediction in environments where both delay and bandit feedback are present.
-We obtain algorithms for this setup by combining the theoretical developments for bandit feedback without delay and delayed full information feedback, offering surrogate regret bounds of $O(\sqrt{DKT})$ (Theorem 5.1) and $O(D^{1/3}T^{2/3})$ (Theorem 5.2).
-We validate our algorithms through numerical experiments using both synthetic and real-world data.
-Specifically, we consider online multiclass classification with bandit feedback.
-We observe that, depending on the number of classes and the dataset, our algorithm, designed for general structured prediction, can achieve accuracy comparable to existing algorithms specialized for multiclass classification.
+To this end, we draw inspiration from pseudo-inverse estimators used in the adversarial linear/combinatorial bandit literature [14, 1, 8].
+Indeed, we cannot naively use the existing estimators, and hence we design a new gradient estimator that applies to various structured prediction problems with target losses belonging to the special SELF class.
+Armed with this gradient estimator, we achieve a surrogate regret upper bound of $O(T^{2/3})$ (Theorem 3.7).
+This successfully eliminates the explicit dependence on $K$, although the dependence on $T$ increases compared to the $O(\sqrt{KT})$ bound.
+We also numerically observe the benefit of this approach when $K$ is large, aligning with the implication of the theoretical bounds (see Section H.2).
+For the delayed feedback setting, the surrogate regret bounds depend on whether the delay time is fixed or variable.
+We here describe our results for the known fixed delay time, denoted by $D$, under the full-information setting.
+It is relatively straightforward to obtain a surrogate regret bound of $O(\sqrt{DT})$ with standard Online Convex Optimization (OCO) algorithms for the delayed feedback.
+We improve upon this bound and achieve a surrogate regret bound of $O(\min\{D^{2},D^{2/3}T^{1/3}\})$, which is notably independent of $T$ (Theorems 4.3 and E.3).
+The proof requires carefully bridging different lines of research: an AdaGrad-type OCO algorithm for delayed feedback [17] and surrogate regret analysis in OSP.
+Given the contributions so far, it is natural to explore OSP in environments where both delay and bandit feedback are present.
+We develop algorithms for this setting by combining the theoretical developments for bandit feedback without delay and delayed full-information feedback, achieving surrogate regret bounds of $O(\sqrt{(D+K)T})$ (Theorem 5.1) and $O(D^{1/3}T^{2/3})$ (Theorem 5.2).
+It is worth noting that similar surrogate regret bounds can also be obtained in the variable-delay setting—where the delay may differ for each round of feedback—for both full-information and bandit feedback.
+Due to space constraints, most of the details are provided in Appendix G.
 2 Preliminaries
-We describe the detailed setup of online structured prediction and key tools used in this work: the Fenchel–Young loss, SELF, and randomized decoding.
+This section describes the detailed setting of OSP and key tools used in this work: the Fenchel–Young loss, SELF, and randomized decoding.
 Notation
 […]
-2.1 Online Structured Prediction
-Here, we describe the problem setting of online structured prediction.
-Let $\mathcal{X}$ be the input vector space and $\mathcal{Y}$ be the output vector space.
+2.1 Online structured prediction (OSP)
+We describe the problem setting of OSP.
+Let $\mathcal{X}$ be the space of input vectors and $\mathcal{Y}$ be the finite set of outputs.
 Define $K\coloneqq|\mathcal{Y}|$.
-Following Blondel et al.
-(2020) and Sakaue et al.
-(2024), we assume that $\mathcal{Y}$ is embedded into $\mathbb{R}^{d}$ in a standard manner.
-For example, $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$ in $d$-class multiclass classification.
-A linear estimator $\bm{W}\in\mathcal{W}$ for a convex domain $\mathcal{W}$ is used to transform the input vector $\bm{x}$ into the score vector $\bm{W}\bm{x}$.
-In online structured prediction, at each round $t=1,\dots,T$:
-  1. 1.
-    The environment selects an input $\bm{x}_{t}\in\mathcal{X}$ and the true output $\bm{y}_{t}\in\mathcal{Y}$;
-  2. 2.
-    The learner receives $\bm{x}_{t}$ and computes the score vector $\bm{\theta}_{t}=\bm{W}_{t}\bm{x}_{t}$ using the linear estimator $\bm{W}_{t}$;
-  3. 3.
-    The learner selects a predicted output $\widehat{\bm{y}}_{t}$ based on $\bm{\theta}_{t}$ and incurs a loss of $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$;
-  4. 4.
-    The learner receives feedback based on the problem setup and updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ using an online learning algorithm, $\mathsf{ALG}$.
-The goal of the learner is to minimize the cumulative prediction loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$, which is equivalent to minimizing the surrogate regret $\mathcal{R}_{T}$ in (1).
-We assume that the input and output are generated in an oblivious manner.
-Note that when $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\dots,\bm{\mathrm{e}}_{d}\}$ and $L(\widehat{\bm{y}}_{t};\bm{y}_{t})=\mathds{1}[\widehat{\bm{y}}_{t}\neq\bm{y}_{t}]$, the above setting reduces to online multiclass classification, which was studied by Van der Hoeven (2020) and Van der Hoeven et al.
-(2021).
-We will use $B\coloneqq\mathrm{diam}(\mathcal{W})$, $C_{x}\coloneqq\mathrm{diam}(\mathcal{X})$, and $C_{y}\coloneqq\mathrm{diam}(\mathcal{Y})$ to denote the diameters of the sets $\mathcal{W}$, $\mathcal{X}$, and $\mathcal{Y}$, respectively.
+Following the literature [6, 35], we assume that $\mathcal{Y}$ is embedded into $\mathbb{R}^{d}$ in a standard manner, e.g., $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$ in multiclass classification with $d$ classes.
+Let $\mathcal{W}$ be a closed convex domain.
+A linear estimator $\bm{W}\in\mathcal{W}$ transforms the input vector $\bm{x}$ into the score vector $\bm{W}\bm{x}$.
+In OSP, at each round $t=1,\dots,T$, the environment selects an input $\bm{x}_{t}\in\mathcal{X}$ and the true output $\bm{y}_{t}\in\mathcal{Y}$.
+The learner receives $\bm{x}_{t}$ and computes the score vector $\bm{\theta}_{t}=\bm{W}_{t}\bm{x}_{t}$ using the linear estimator $\bm{W}_{t}$.
+Then, the learner selects a prediction $\widehat{\bm{y}}_{t}$ based on $\bm{\theta}_{t}$, which is called decoding, and incurs a loss of $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$.
+Finally, the learner receives feedback, which depends on the problem setting, and updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ using some online learning algorithm, denoted by $\mathsf{ALG}$, which is applied to the surrogate loss function $\bm{W}\mapsto S(\bm{W}\bm{x}_{t};\bm{y}_{t})$.
+The goal of the learner is to minimize the cumulative target loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$, which is equivalent to minimizing the surrogate regret $\mathcal{R}_{T}=\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}-\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}.$ We assume that the input and output are generated in an oblivious manner.
+Note that when $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\dots,\bm{\mathrm{e}}_{d}\}$ and $L(\widehat{\bm{y}}_{t};\bm{y}_{t})=\mathds{1}[\widehat{\bm{y}}_{t}\neq\bm{y}_{t}]$, the above setting reduces to online multiclass classification, which was studied in prior work [38, 39].
+Let $B\coloneqq\mathrm{diam}(\mathcal{W})$ denote the diameter of $\mathcal{W}$ measured by $\|\cdot\|_{\mathrm{F}}$ and $C_{x}\coloneqq\max_{\bm{x}\in\mathcal{X}}\|\bm{x}\|_{2}$ the maximum Euclidean norm of input vectors in $\mathcal{X}$.
 The feedback observed by the learner depends on the problem setting.
-The most fundamental setting is the full information setup, where the true output $\bm{y}_{t}$ is observed as feedback at the end of each round $t$.
-This setup was extensively investigated in Sakaue et al.
-(2024).
-By contrast, our study investigates the following weaker feedback:
-  • Bandit feedback: Only the value of the loss function is observed.
+The most basic setting is the full-information setting, where the true output $\bm{y}_{t}$ is observed as feedback at the end of each round $t$, extensively investigated by Sakaue et al.
+[35].
+By contrast, we study the following weaker feedback: In the bandit feedback setting, only the value of the loss function is observed.
 Specifically, at the end of each round $t$, the learner observes the target loss value $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$ as feedback.
-  • Delayed feedback: The feedback is observed with a certain delay.
-We consider a fixed $D$-round delay setting, i.e., no feedback is received for round $t\leq D$, and for $t>D$, the learner observes either full information feedback $\bm{y}_{t-D}$ or bandit feedback $L(\widehat{\bm{y}}_{t-D};\bm{y}_{t-D})$.
+In the delayed feedback setting, the feedback is observed with a certain delay.
+In this paper, we consider especially a fixed $D$-round delay setting, i.e., no feedback is received for round $t\leq D$, and for $t>D$, the learner observes either full-information feedback $\bm{y}_{t-D}$ or bandit feedback $L(\widehat{\bm{y}}_{t-D};\bm{y}_{t-D})$.
+We also discuss the variable-delay setting in Appendix G.
 In this paper, we make the following assumptions:
@@
 (1) There exists $\nu>0$ such that for any distinct $\bm{y},\bm{y}^{\prime}\in\mathcal{Y}$, it holds that $\|\bm{y}-\bm{y}^{\prime}\|\geq\nu$.
-(2) For each $\bm{y}\in\mathcal{Y}$, the target loss function $L(\cdot;\bm{y})$ is defined on $\operatorname{conv}(\mathcal{Y})$, is non-negative, and is affine with respect to its first argument.
+(2) For each $\bm{y}\in\mathcal{Y}$, the target loss function $L(\cdot;\bm{y})$ is defined on $\operatorname{conv}(\mathcal{Y})$, non-negative, and affine w.r.t.
+its first argument.
 (3) There exists $\gamma$ such that for any $\bm{y}^{\prime}\in\operatorname{conv}(\mathcal{Y})$ and $\bm{y}\in\mathcal{Y}$, it holds that $L(\bm{y}^{\prime};\bm{y})\leq\gamma\|\bm{y}^{\prime}-\bm{y}\|$ and $L(\bm{y}^{\prime};\bm{y})\leq 1$.
-(4) It holds that $L(\bm{y}^{\prime};\bm{y})=0$ only if $\bm{y}^{\prime}=\bm{y}$.
-As discussed in Sakaue et al.
-(2024, Section 2.3), these assumptions are natural and hold for a broad range of problem settings and target loss functions, including SELF (see Section 2.4 for the formal definition).
+(4) It holds that $L(\bm{y}^{\prime};\bm{y})=0$ if and only if $\bm{y}^{\prime}=\bm{y}$.
+As discussed in [35, Section 2.3], these assumptions are natural and hold for various structured prediction problems and target loss functions, including SELF (see Section 2.4 for the formal definition).
 2.2 Fenchel–Young loss
-We use the Fenchel–Young loss (Blondel et al., 2020) as the surrogate loss, which subsumes many representative surrogate losses, such as logistic loss, Conditional Random Field (CRF) loss (Lafferty et al., 2001), and SparseMAP (Niculae et al., 2018).
-See Blondel et al.
-(2020, Table 1) for more examples.
-Definition 2.2 (Blondel et al. 2020, Fenchel–Young loss).
-Let $\Omega:\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a regularization function with $\mathcal{Y}\subset\operatorname{dom}(\Omega)$.
-The Fenchel–Young loss generated by $\Omega$, denoted by $S_{\Omega}:\operatorname{dom}(\Omega^{\ast})\times\operatorname{dom}(\Omega)\rightarrow\mathbb{R}_{\geq 0}$, is defined as
-$S_{\Omega}(\bm{\theta};\bm{y})\coloneqq\Omega^{\ast}(\bm{\theta})+\Omega(\bm{y})-\langle\bm{\theta},\bm{y}\rangle.$
-The Fenchel–Young loss has the following properties, which will be useful in the subsequent discussion:
-Proposition 2.3 (Blondel et al. 2020, Propositions 2 and 3 and Sakaue et al. 2024, Proposition 3).
+We use the Fenchel–Young loss [6] as the surrogate loss, which subsumes many representative surrogate losses, such as the logistic loss, Conditional Random Field (CRF) loss [25], and SparseMAP [30].
+See [6, Table 1] for more examples.
+Definition 2.2 ([6, Fenchel–Young loss]).
+Let $\Omega\colon\mathbb{R}^{d}\rightarrow\mathbb{R}\cup\{+\infty\}$ be a regularization function with $\mathcal{Y}\subset\operatorname{dom}(\Omega)$.
+The Fenchel–Young loss generated by $\Omega$, denoted by $S_{\Omega}\colon\operatorname{dom}(\Omega^{\ast})\times\operatorname{dom}(\Omega)\rightarrow\mathbb{R}_{\geq 0}$, is defined as $S_{\Omega}(\bm{\theta};\bm{y})\coloneqq\Omega^{\ast}(\bm{\theta})+\Omega(\bm{y})-\langle\bm{\theta},\bm{y}\rangle.$
+The Fenchel–Young loss has the following useful properties:
+Proposition 2.3 ([6, Propositions 2 and 3] and [35, Proposition 3]).
 […]
-$\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ $\coloneqq\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Omega(\bm{y})\>:\>\bm{y}\in\mathbb{R}^{d}\}$
-$=\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Psi(\bm{y})\>:\>\bm{y}\in\operatorname{conv}(\mathcal{Y})\}.$
-Then, for any $\bm{y}\in\mathcal{Y}$, $S_{\Omega}(\bm{\theta},\bm{y})$ is differentiable with respect to $\bm{\theta}$, and it satisfies $\nabla S_{\Omega}(\bm{\theta};\bm{y})=\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{y}.$ Furthermore, it holds that $S_{\Omega}(\bm{\theta};\bm{y})\geq\frac{\lambda}{2}\|\bm{y}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|^{2}.$
 In what follows, let $S_{t}(\bm{W})\coloneqq S_{\Omega}(\bm{W}\bm{x}_{t};\bm{y}_{t})$ for simplicity.
@@
 $\|\nabla S_{t}(\bm{W})\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}).$ (1)
-Indeed, from Proposition 2.3 and Assumption 2.1, we have $\|\nabla S_{t}(\bm{W}_{t})\|_{\mathrm{F}}^{2}=\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|_{2}^{2}\|\bm{x}_{t}\|^{2}\leq C_{x}^{2}\kappa^{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|^{2}\leq\frac{2C_{x}^{2}\kappa^{2}}{\lambda}S_{t}(\bm{W}_{t})$, where we used $\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$ and $\|\cdot\|_{2}\leq\kappa\|\cdot\|$.
+Indeed, from Proposition 2.3, we have $\|\nabla S_{t}(\bm{W}_{t})\|_{\mathrm{F}}^{2}=\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|_{2}^{2}\|\bm{x}_{t}\|_{2}^{2}\leq C_{x}^{2}\kappa^{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|^{2}\leq\frac{2C_{x}^{2}\kappa^{2}}{\lambda}S_{t}(\bm{W}_{t})$, where we used $\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$ and $\|\cdot\|_{2}\leq\kappa\|\cdot\|$.
 Thus, (1) holds with $b=\frac{2C_{x}^{2}\kappa^{2}}{\lambda}$.
-Below, let $L_{t}(\bm{y})\coloneqq L(\bm{y};\bm{y}_{t})$, $\bm{G}_{t}\coloneqq\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$, and $L\coloneqq\max_{\bm{W}\in\mathcal{W}}\|\nabla S_{t}(\bm{W})\|_{\mathrm{F}}$.
+Below, let $L_{t}(\bm{y})\coloneqq L(\bm{y};\bm{y}_{t})$ and $\bm{G}_{t}\coloneqq\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$.
 2.3 Examples of structured prediction
-We present several special cases of structured prediction along with specific parameter values introduced so far; see Sakaue et al. (2024, Section 2.3) for further details.
+We present several instances of structured prediction along with specific parameter values introduced so far; see [35, Section 2.3] for further details.
 Multiclass classification
 Let $\mathcal{Y}=\{\mathbf{e}_{1},\dots,\mathbf{e}_{d}\}$ and $\|\cdot\|=\|\cdot\|_{1}$.
-When using the 0-1 loss, $L(\bm{y}^{\prime};\bm{y})=\mathds{1}[\bm{y}^{\prime}\neq\bm{y}]$, the parameters in Assumption 2.1 are $\nu=2$ and $\gamma=\frac{1}{2}$.
-The logistic surrogate loss is a Fenchel–Young loss $S_{\Omega}$ generated by the entropy regularization function $\Omega=\mathsf{H}^{s}+I_{\Delta_{d}}$ (up to a constant factor), where $\mathsf{H}^{s}(\bm{y})\coloneqq-\sum_{i=1}^{d}y_{i}\log y_{i}$ and $\Delta_{d}$ is the $(d-1)$-dimensional probability simplex.
-Since $\Omega$ is a $1$-strongly convex function with respect to $\|\cdot\|_{1}$, we have $\lambda=1$.
+As the 0-1 loss satisfies $L(\bm{y};\mathbf{e}_{i})=\mathds{1}[\bm{y}\neq\bm{y}]=\sum_{j\neq i}y_{j}=\frac{1}{2}(1-y_{i}+\sum_{j\neq i}y_{j})=\frac{1}{2}\|\mathbf{e}_{i}-\bm{y}\|_{1}$, we have $\gamma=\frac{1}{2}$ in Assumption 2.1.
+Also, $\nu=2$ holds since $\|\mathbf{e}_{i}-\mathbf{e}_{j}\|_{1}=2$ if $i\neq j$.
+The logistic surrogate loss is a Fenchel–Young loss $S_{\Omega}$ generated by the negative Shannon entropy $\Omega=\mathsf{H}^{s}+I_{\Delta_{d}}$ (up to a constant factor originating from the base of $\log$), where $\mathsf{H}^{s}(\bm{y})\coloneqq-\sum_{i=1}^{d}y_{i}\log y_{i}$ and $\Delta_{d}$ is the $(d-1)$-dimensional probability simplex.
+Since $\Omega$ is a $1$-strongly convex function w.r.t.
+$\|\cdot\|_{1}$ on $\Delta_{d}$, we have $\lambda=1$.
 Multilabel classification
@@
 The SparseMAP surrogate loss $S_{\Omega}(\bm{\theta},\bm{y})=\frac{1}{2}\|\bm{y}-\bm{\theta}\|_{2}^{2}-\frac{1}{2}\|\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{\theta}\|_{2}^{2}$ is a Fenchel–Young loss generated by $\Omega=\frac{1}{2}\|\cdot\|^{2}+I_{\operatorname{conv}(\mathcal{Y})}$.
-Since $\Omega$ is 1-strongly convex with respect to $\|\cdot\|_{2}$, we have $\lambda=1$.
+Since $\Omega$ is 1-strongly convex w.r.t.
+$\|\cdot\|_{2}$, we have $\lambda=1$.
 Ranking
@@
 We use the target loss function that counts mismatches, $L(\bm{y}^{\prime};\bm{y})=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[y_{i,j_{i}}^{\prime}\!\neq\!y_{i,j_{i}}]$, where $j_{i}\in[m]$ is a unique index with $y_{ij_{i}}=1$ for each $i\in[m]$.
-In this case, the parameters in Assumption 2.1 satisfy $\nu=4$ and $\gamma=\frac{1}{2m}$.
-We use a surrogate loss given by $S_{\Omega}(\bm{\theta};\bm{y})=\langle\bm{\theta},\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{y}\rangle+\frac{1}{\zeta}\mathsf{H}^{s}(\widehat{\bm{y}}_{\Omega}(\bm{\theta}))$, where $\Omega=-\frac{1}{\zeta}\mathsf{H}^{s}+I_{\operatorname{conv}(\mathcal{Y})}$ and $\zeta$ is a parameter controlling the regularization strength.
+In this case, Assumption 2.1 is satisfied with $\nu=4$ and $\gamma=\frac{1}{2m}$.
+We use a surrogate loss given by $S_{\Omega}(\bm{\theta};\bm{y})\!=\!\langle\bm{\theta},\widehat{\bm{y}}_{\Omega}(\bm{\theta})-\bm{y}\rangle\!+\!\frac{1}{\zeta}\mathsf{H}^{s}(\widehat{\bm{y}}_{\Omega}(\bm{\theta}))$, where $\Omega\!=\!-\frac{1}{\zeta}\mathsf{H}^{s}\!+\!I_{\operatorname{conv}(\mathcal{Y})}$ and $\zeta$ controls the regularization strength.
 The first term in $S_{\Omega}$ measures the affinity between $\bm{\theta}$ and $\bm{y}$, while the second term evaluates the uncertainty of $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$.
@@
 2.4 Structured encoding loss function (SELF)
-Here we introduce a common class of target loss functions, called the Structured Encoding Loss Function (SELF). A target loss function is SELF if it can be expressed as
+We introduce a general class of target loss functions, called the (generalized) Structured Encoding Loss Function (SELF) [10, 11, 5]. A target loss function is SELF if it can be expressed as
 $L(\bm{y}_{t};\widehat{\bm{y}}_{t})=\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}+\bm{b}\rangle+c(\bm{y}_{t}),$ (2)
-where $\bm{b}\in\mathbb{R}^{d}$ is a constant vector, $\bm{V}\in\mathbb{R}^{d\times d}$ is a constant matrix, and $c:\mathcal{Y}\to\mathbb{R}$ is a function.
-The following loss examples, taken from Blondel (2019, Appendix A), belong to the SELF class:
+where $\bm{b}\in\mathbb{R}^{d}$ is a constant vector, $\bm{V}\in\mathbb{R}^{d\times d}$ is a constant matrix, and $c\colon\mathcal{Y}\to\mathbb{R}$ is a function.
+The following examples of target losses, taken from [5, Appendix A], belong to the SELF class:
   • Multiclass classification: the 0-1 loss is a SELF with $\bm{V}=\bm{1}\bm{1}^{\top}-\bm{I}$, $\bm{b}=\bm{0}$, and $c(\bm{y})=0$.
-  • Multilabel classification: the Hamming loss, $L(\bm{y}^{\prime};\bm{y})=\frac{1}{d}\sum_{i=1}^{d}\mathds{1}[y^{\prime}_{t,i}\neq y_{i}]$, is a SELF with $\bm{V}=-\frac{2}{d}\bm{I}$, $\bm{b}=\frac{\bm{1}}{d}$, and $c(\bm{y})=\frac{1}{d}\langle\bm{y},\bm{1}\rangle$, where the last factor is constant if the number of correct labels is fixed.
+  • Multilabel classification: the Hamming loss, $L(\bm{y}^{\prime};\bm{y})\!=\!\frac{1}{d}\sum_{i=1}^{d}\mathds{1}[y^{\prime}_{t,i}\!\neq\!y_{i}]$, is a SELF with $\bm{V}\!=\!-\frac{2}{d}\bm{I}$, $\bm{b}\!=\!\frac{1}{d}\bm{1}$, and $c(\bm{y})\!=\!\frac{1}{d}\langle\bm{y},\bm{1}\rangle$, where $c(\bm{y})$ is constant if the number of correct labels is fixed.
   • Ranking: the Hamming loss $L(\bm{y}^{\prime};\bm{y})=\frac{1}{m}\sum_{i=1}^{m}\mathds{1}[y_{i,j_{i}}^{\prime}\neq y_{i,j_{i}}]$, where $j_{i}\in[m]$ is a unique index with $y_{i,j_{i}}=1$ for each $i\in[m]$, is a SELF with $\bm{V}=-\frac{1}{m}\bm{I}$, $\bm{b}=\bm{0}$, and $c(\bm{y})=1$.
-Following Sakaue et al. (2024), this study assumes that the target loss function $L$ is a SELF.
+Following the work by Sakaue et al. [35], we assume that the target loss function $L$ is a SELF.
 2.5 Randomized decoding
@@
 1:$\bm{\theta}\in\mathbb{R}^{d}$
-2:$\widehat{\bm{y}}_{\Omega}(\bm{\theta})\leftarrow\operatorname*{arg\,max}\{\langle\bm{\theta},\bm{y}\rangle-\Psi(\bm{y})\>:\>\bm{y}\in\operatorname{conv}(\mathcal{Y})\}$
-3:$\bm{y}^{\ast}\leftarrow\operatorname*{arg\,min}\{\|\bm{y}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|\>:\>\bm{y}\in\mathcal{Y}\}$
-4:$\Delta^{\ast}\leftarrow\|\bm{y}^{\ast}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|,\>p\leftarrow\min\{1,2\Delta^{\ast}/\nu\}$
-5:Sample $Z\sim\mathrm{Ber}(p)$
-6:if $Z=0$ then $\bm{\widehat{y}}\leftarrow\bm{y}^{\ast}$
-7:if $Z=1$ then $\bm{\widehat{y}}\leftarrow\bm{\widetilde{y}}$ where $\bm{\widetilde{y}}$ is randomly drawn from $\mathcal{Y}$ so that $\mathbb{E}\left[\bm{\widetilde{y}}|Z=1\right]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$
+2:Compute $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ defined in Proposition 2.3.
+3:$\bm{y}^{\ast}\leftarrow\operatorname*{arg\,min}\{\|\bm{y}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|\>:\>\bm{y}\in\mathcal{Y}\}$.
+4:$\Delta^{\ast}\!\leftarrow\!\|\bm{y}^{\ast}-\widehat{\bm{y}}_{\Omega}(\bm{\theta})\|,\>p\!\leftarrow\!\min\{1,2\Delta^{\ast}/\nu\}$.
+5:Sample $Z\sim\mathrm{Ber}(p)$.
+6:if $Z=0$ then $\bm{\widehat{y}}\leftarrow\bm{y}^{\ast}$.
+7:if $Z=1$ then $\bm{\widehat{y}}\leftarrow\bm{\widetilde{y}}$ where $\bm{\widetilde{y}}$ is randomly drawn from $\mathcal{Y}$ so that $\mathbb{E}\left[\bm{\widetilde{y}}|Z=1\right]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$.
 8:$\phi_{\Omega}(\bm{\theta})=\bm{\widehat{y}}$
-The procedure of converting the estimated score $\bm{\theta}$ into a structured output $\bm{\widehat{y}}$ is called decoding.
-For this, we employ randomized decoding (Sakaue et al., 2024), which plays an essential role particularly in deriving an upper bound independent of the output set size $K=\lvert\mathcal{Y}\rvert$ in Section 3.4.
-The randomized decoding (Algorithm 1) selects either the closest $\bm{y}^{*}\in\mathcal{Y}$ to $\widehat{\bm{y}}_{\Omega}(\bm{\theta})\in\operatorname{conv}(\mathcal{Y})$ or a random $\widetilde{\bm{y}}\in\mathcal{Y}$ satisfying $\mathbb{E}[\widetilde{\bm{y}}\mid Z=1]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$, where $Z$ follows a Bernoulli distribution with a parameter $p$.
-Intuitively, the parameter $p$ is chosen so that if $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ is close to $\bm{y}^{*}$, the decoding is more likely to return $\bm{y}^{*}$; otherwise, it is more likely to return $\widetilde{\bm{y}}$, reflecting uncertainty.
-An important property satisfied by the randomized decoding is the following lemma, which we use in the subsequent analysis:
-Lemma 2.4 (Sakaue et al. 2024, Lemma 4).
-For any $(\bm{\theta},\bm{y})\in\mathbb{R}^{d}\times\mathcal{Y}$, the randomized decoding $\phi_{\Omega}$ satisfies
-$\mathbb{E}[L(\phi_{\Omega}(\bm{\theta});\bm{y})]\leq\frac{4\gamma}{\lambda\nu}S_{\Omega}(\bm{\theta};\bm{y}),$
-where the expectation is taken with respect to the internal randomness of the randomized decoding.
+We employ the randomized decoding [35], which plays an essential role, particularly in deriving an upper bound independent of the output set size $K=\lvert\mathcal{Y}\rvert$ in Section 3.4.
+The randomized decoding (Algorithm 1) returns either the closest $\bm{y}^{*}\in\mathcal{Y}$ to $\widehat{\bm{y}}_{\Omega}(\bm{\theta})\in\operatorname{conv}(\mathcal{Y})$ (see Proposition 2.3) or a random $\widetilde{\bm{y}}\in\mathcal{Y}$ satisfying $\mathbb{E}[\widetilde{\bm{y}}\mid Z=1]=\widehat{\bm{y}}_{\Omega}(\bm{\theta})$, where $Z$ follows the Bernoulli distribution with a parameter $p$.
+Intuitively, the parameter $p$ is chosen so that if $\widehat{\bm{y}}_{\Omega}(\bm{\theta})$ is close to $\bm{y}^{*}$, the decoding more likely returns $\bm{y}^{*}$; otherwise, it more likely returns $\widetilde{\bm{y}}$, reflecting uncertainty.
+An crucial property of the randomized decoding is the following lemma, which we use in the subsequent analysis:
+Lemma 2.4 ([35, Lemma 4]).
+For any $(\bm{\theta},\bm{y})\!\!\in\!\mathbb{R}^{d}\times\mathcal{Y}$, the randomized decoding $\phi_{\Omega}$ satisfies $\mathbb{E}[L(\phi_{\Omega}(\bm{\theta});\bm{y})]\!\leq\frac{4\gamma}{\lambda\nu}S_{\Omega}(\bm{\theta};\bm{y})$, where the expectation is taken w.r.t.
+the randomness of $\phi_{\Omega}$.
 3 Bandit feedback
-In this section, we present two online structured prediction algorithms in the bandit feedback setup and analyze their regret.
-Our results here are mostly special cases of the regret bounds obtained when handling bandit and delayed feedback (Section 5).
-Nevertheless, by focusing on the case without delay, we provide a clearer exposition of the core ideas.
+In this section, we present two OSP algorithms for the bandit feedback setting and analyze their surrogate regret.
+Our results can be extended to handle bandit and delayed feedback; see Section 5.
+Here, we focus on the simpler case without delay to provide a clearer exposition of our core ideas.
 3.1 Randomized decoding with uniform exploration
@@
 1:$\bm{\theta}\in\mathbb{R}^{n}$, $q\in[0,1]$
-2:Sample $X\sim\mathrm{Ber}(q)$
-3:if $X=0$ then $\bm{\widehat{y}}\leftarrow\phi_{\Omega}(\bm{\theta})$
-4:if $X=1$ then Sample $\bm{y}^{\ast}\sim\mathrm{Unif}(\mathcal{Y})$ and $\bm{\widehat{y}}\leftarrow\bm{y}^{\ast}$
+2:Sample $X\sim\mathrm{Ber}(q)$.
+3:if $X=0$ then ${\widehat{\bm{y}}}\leftarrow\phi_{\Omega}(\bm{\theta})$.
+4:if $X=1$ then Sample $\bm{y}^{\ast}$ from $\mathcal{Y}$ uniformly at random and $\widehat{\bm{y}}\!\leftarrow\!\bm{y}^{\ast}$.
 5:$\psi_{\Omega}(\bm{\theta})=\bm{\widehat{y}}$
-Here, we discuss the properties of the decoding function, Randomized Decoding with Uniform Exploration (RDUE), which will be used in both algorithms to convert scores into outputs.
-As discussed in Section 2.5, in online structured prediction with full information feedback, the randomized decoding (Algorithm 1) was introduced as a decoding function (Sakaue et al., 2024).
-However, naively applying the randomized decoding does not lead to a satisfactory regret bound under bandit feedback.
-We extend the framework of the randomized decoding to handle bandit feedback effectively.
+We discuss the properties of our decoding function, called Randomized Decoding with Uniform Exploration (RDUE), which will be used in subsequent sections.
+As discussed in Section 2.5, the randomized decoding (Algorithm 1) was introduced as a decoding function [35] for OSP with full-information feedback.
+However, naively applying it does not lead to a desired bound under bandit feedback due to the lack of exploration.
+We extend the randomized-decoding framework to handle bandit feedback effectively.
 RDUE (Algorithm 2) is a procedure that, with probability $q\in[0,1]$, selects $\widehat{\bm{y}}$ uniformly at random from $\mathcal{Y}$, and with probability $1-q$, selects the output of the randomized decoding.
-Using RDUE, we define $p_{t}(\bm{y})$ as the probability that $\widehat{\bm{y}}_{t}$ coincides with $\bm{y}$ at round $t$.
-Note that for any $\bm{y}\in\mathcal{Y}$, it holds that $p_{t}(\bm{y})\geq\frac{q}{K}.$ Furthermore, similar to the property of the randomized decoding in Lemma 2.4, RDUE satisfies the following property:
+Let $p_{t}(\bm{y})$ be the probability that the output of RDUE, $\widehat{\bm{y}}_{t}$, coincides with $\bm{y}$ at round $t$.
+Note that for any $\bm{y}\in\mathcal{Y}$, it holds that $p_{t}(\bm{y})\geq\frac{q}{K}$ thanks to the uniform exploration.
+Furthermore, similar to the property of the randomized decoding in Lemma 2.4, RDUE satisfies the following property:
 Lemma 3.1.
-For any $(\bm{\theta},\bm{y})\in\mathbb{R}^{d}\times\mathcal{Y}$, RDUE $\psi_{\Omega}$ satisfies
-$\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]\leq\frac{4\gamma}{\lambda\nu}(1-q)S_{\Omega}(\bm{\theta};\bm{y})+q\frac{K-1}{K},$
-where the expectation is taken with respect to the internal randomness of RDUE.
+For any $(\bm{\theta},\bm{y})\in\mathbb{R}^{d}\times\mathcal{Y}$, RDUE $\psi_{\Omega}$ satisfies $\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]\leq\frac{4\gamma}{\lambda\nu}(1-q)S_{\Omega}(\bm{\theta};\bm{y})+q\frac{K-1}{K},$ where the expectation is taken w.r.t.
+the internal randomness of RDUE.
 Proof.
-Since the output of the randomized decoding is chosen with probability $1-q$, and a uniformly random output is chosen with probability $q$, we have $\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]=(1-q)\mathbb{E}\left[L(\phi_{\Omega}(\bm{\theta});\bm{y})\right]+q\frac{K-1}{K},$ where we used $L(\cdot;\cdot)\leq 1$ and $\phi_{\Omega}$ is the randomized decoding.
-Hence, combining this with Lemma 2.4, we obtain the desired bound.
-∎
-Additionally, this section makes the following assumption:
+With probability $1-q$, the randomized decoding is used; otherwise, a uniformly random output is chosen.
+Thus, $\mathbb{E}\left[L(\psi_{\Omega}(\bm{\theta});\bm{y})\right]=(1-q)\mathbb{E}\left[L(\phi_{\Omega}(\bm{\theta});\bm{y})\right]+q\frac{K-1}{K}$ holds, where $\phi_{\Omega}$ is the randomized decoding and we used $L(\cdot;\cdot)\leq 1$.
+Combining this with Lemma 2.4 completes the proof.∎
+Based on this lemma, we make the following assumption for convenience:
 Assumption 3.2.
-There exists $a\in(0,1)$ such that
-$\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t})+q.$
-Here, $\mathbb{E}_{t}\left[{\cdot}\right]$ denotes the conditional expectation given the random variables $\widehat{\bm{y}}_{1},\dots,\widehat{\bm{y}}_{t-1}$.
-This assumption can be satisfied by using RDUE and letting $a\leq 1-\frac{4\gamma}{\lambda\nu}(1-q)$ if $\lambda>\frac{4\gamma}{\nu}(1-q)$, according to Lemma 3.1.
+There exists $a\in(0,1)$ such that $\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t})+q.$ Here, $\mathbb{E}_{t}\left[{\cdot}\right]$ denotes the conditional expectation given the past outputs, $\widehat{\bm{y}}_{1},\dots,\widehat{\bm{y}}_{t-1}$.
+This assumption can be satisfied by using RDUE for $a\leq 1-\frac{4\gamma}{\lambda\nu}(1-q)$ if $\lambda>\frac{4\gamma}{\nu}(1-q)$, due to Lemma 3.1.
 In what follows, let $a=1-\frac{4\gamma}{\lambda\nu}$.
-Note that $\lambda\geq\frac{4\gamma}{\nu}$ holds in the cases of multiclass classification, multilabel classification, and ranking, as discussed in Section 2.3.
+Note that $\lambda>\frac{4\gamma}{\nu}\geq\frac{4\gamma}{\nu}(1-q)$ holds in the cases of multiclass classification, multilabel classification, and ranking (see Section 2.3 and [35] for details).
 3.2 Online gradient descent
-The algorithm in this section uses the adaptive Online Gradient Descent (OGD, Streeter & McMahan 2010) as $\mathsf{ALG}$.
-OGD updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ using a gradient $\bm{G}_{t}$ and learning rate $\eta_{t}$ by $\bm{W}_{t+1}\leftarrow\Pi_{\mathcal{W}}\left(\bm{W}_{t}-\eta_{t}\bm{G}_{t}\right),$ where $\Pi_{\mathcal{W}}(\bm{Z})=\operatorname*{arg\,min}_{\bm{X}\in\mathcal{W}}\|\bm{X}-\bm{Z}\|_{\mathrm{F}}$.
-OGD with appropriately chosen learning rate $\eta_{t}$ achieves the following bound:
-Lemma 3.3 (Orabona 2019, Theorem 4.14).
-Suppose that we set the learning rate to $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\bm{G}_{i}\|_{\mathrm{F}}^{2}}}$ and do not update on rounds when $\bm{G}_{t}$ is the all-zero matrix.
-Then, for any $\bm{U}\in\mathcal{W}$, OGD achieves $\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}}.$
-3.3 $O(\sqrt{KT})$ Regret Algorithm
-Here, we present an algorithm that achieves a regret upper bound of $O(\sqrt{KT})$.
-Algorithm based on inverse-weighted estimator
-In the bandit setting, the true output $\bm{y}_{t}$ is not observed, and thus it is necessary to estimate the gradient required for updating $\bm{W}_{t}$.
-To deal with this, we use the following inverse-weighted gradient estimator:
-$\widetilde{\bm{G}}_{t}\coloneqq\frac{\mathds{1}[\widehat{\bm{y}}_{t}=\bm{y}_{t}]}{p_{t}(\bm{y}_{t})}\bm{G}_{t},$ (4)
-where we recall that $\bm{G}_{t}=\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$.
-Note that $\widetilde{\bm{G}}_{t}$ is unbiased, i.e., $\mathbb{E}\big{[}\widetilde{\bm{G}}_{t}\big{]}=\bm{G}_{t}$.
+We use the adaptive Online Gradient Descent (OGD) [36] as $\mathsf{ALG}$, which we apply to surrogate loss $S_{t}$.
+OGD updates $\bm{W}_{t}$ to $\bm{W}_{t+1}$ by using the gradient $\bm{G}_{t}=\nabla S_{t}(\bm{W}_{t})$ and learning rate $\eta_{t}$ as $\bm{W}_{t+1}\leftarrow\Pi_{\mathcal{W}}\left(\bm{W}_{t}-\eta_{t}\bm{G}_{t}\right),$ where $\Pi_{\mathcal{W}}(\bm{Z})=\operatorname*{arg\,min}_{\bm{X}\in\mathcal{W}}\|\bm{X}-\bm{Z}\|_{\mathrm{F}}$.
+OGD achieves the following bound:
+Lemma 3.3 (e.g., [31, Theorem 4.14]).
+Let $\eta_{t}={B}/{\sqrt{2\sum_{i=1}^{t}\|\bm{G}_{i}\|_{\mathrm{F}}^{2}}}$.
+Then, adaptive OGD achieves $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}}$ for any $\bm{U}\in\mathcal{W}$.
+3.3 Algorithm based on inverse-weighted gradient estimator with $O(\sqrt{KT})$ regret
+We present an algorithm that achieves a surrogate regret upper bound of $O(\sqrt{KT})$.
+Algorithm based on inverse-weighted gradient estimator
+In the bandit setting, the true output $\bm{y}_{t}$ is not observed, and thus we need to estimate the gradient of $S_{t}(\bm{W}_{t})$ required for updating $\bm{W}_{t}$.
+To do this, we use the inverse-weighted gradient estimator $\widehat{\bm{G}}_{t}\coloneqq\frac{\mathds{1}[\widehat{\bm{y}}_{t}=\bm{y}_{t}]}{p_{t}(\bm{y}_{t})}\bm{G}_{t},$ where $\bm{G}_{t}=\nabla S_{t}(\bm{W}_{t})=(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t})\bm{x}_{t}^{\top}$.
+Note that $\widehat{\bm{G}}_{t}$ is unbiased, i.e., $\mathbb{E}\big{[}\widehat{\bm{G}}_{t}\big{]}=\bm{G}_{t}$.
 We use RDUE with $q=B\sqrt{K/T}$ as the decoding function (assuming $T\geq B^{2}K$ for simplicity).
-For $\mathsf{ALG}$, we employ the adaptive OGD in Section 3.2 with the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
+For $\mathsf{ALG}$, we employ the adaptive OGD in Section 3.2 with the learning rate of $\eta_{t}={B}/{\sqrt{2\sum_{i=1}^{t}\|\widehat{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
 Remark 3.4.
@@
 Regret bounds and analysis
-The above algorithm achieves the following regret bound:
+The above algorithm achieves the following surrogate regret bound:
 Theorem 3.5.
-The regret of the above algorithm is bounded by $\mathbb{E}[\mathcal{R}_{T}]\leq\left(\frac{b}{2a}+1\right)B\sqrt{KT}.$
-The $O(\sqrt{KT})$ bound has an optimal dependency on $T$ and matches the $\sqrt{T}$ lower bound in the special case of online multiclass classification with bandit feedback (Van der Hoeven et al., 2021, Corollary 1).
-Furthermore, our bound improves the existing $O(\sqrt{KT})$ bound by Van der Hoeven (2020) by a factor of $\sqrt{K}$.
-Note that, due to differences in the target loss function, our result is not directly comparable to the $\sqrt{KT}$ bound in Van der Hoeven et al.
-(2021).
-A more detailed discussion can be found in Appendix C.
+The above algorithm achieves the surrogate regret of $\mathbb{E}[\mathcal{R}_{T}]\leq\left(\frac{b}{2a}+1\right)B\sqrt{KT}.$
+This upper bound achieves the rate of $\sqrt{T}$, which matches the existing surrogate regret upper bound for bandit multiclass classification in [39].
+Regarding the dependence on $K$, our bound improves the existing $O(K\sqrt{T})$ bound in [38, 39] by a factor of $\sqrt{K}$.
+Note, however, that the $O(K\sqrt{T})$ bound in [39] applies to a broader class of surrogate loss functions.
+For example, in $K$-class classification, their bound applies to the base-$k$ logistic loss for $k\leq K$, while ours is restricted to the baes-2 logistic loss.
+A more detailed discussion is given in Appendix C.
+As for tightness, an $\Omega(\sqrt{T})$ lower bound is provided in [39] for the graph feedback setting, a variant of the bandit feedback model.
+This suggests that the $\sqrt{T}$ dependence would be close to being tight, although this lower bound does not directly apply to the bandit setting.
+Therefore, whether the rate of $\sqrt{T}$ is improvable or not is left open.
 Proof.
-From the convexity of $S_{t}$ and the unbiasedness of $\widetilde{\bm{G}}_{t}$, we have $\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]=\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right].$ From Lemma 3.3, this is further upper bounded as $\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]\leq\sqrt{2}B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}}\right]}\leq B\sqrt{\frac{2bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]},$ where in the first inequality we used Jensen’s inequality and in the last inequality we used $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}),$ which follows from $p_{t}(\bm{y})\geq K/q$ and (2).
-Therefore, from Assumption 3.2, we have $\mathbb{E}[\mathcal{R}_{T}]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\left((1-a)S_{t}(\bm{W}_{t})-S_{t}(\bm{U})\right)}\right]+qT\leq B\sqrt{\frac{2bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT\leq\frac{bB^{2}K}{2aq}+qT,$ where the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
-Finally, substituting $q=B\sqrt{K/T}$ into the last inequality yields the desired bound.
-∎
-We can also prove the following high-probability bound:
-Theorem 3.6.
-Let $\delta\in(0,1)$.
-Then with probability at least $1-\delta$, the same algorithm as in Theorem 3.5, but with a different choice of $q$, achieves the regret bound of $\mathcal{R}_{T}=O\Big{(}\sqrt{KT\log(1/\delta)}+\log(1/\delta)\Big{)}$, where we omit the dependencies on parameters other than $K$, $T$, and $\delta$.
-A more precise statement and proof of this theorem are provided in Section D.2.
-To prove this theorem, we follow the analysis of Theorem 3.5 and use Bernstein’s inequality.
-To address the challenges posed by the randomness introduced by bandit feedback, we adopt an approach similar to that used by Van der Hoeven et al.
-(2021), and arguably, we have successfully simplified their analysis.
-3.4 $O(T^{2/3})$ Regret Algorithm
-While the $O(\sqrt{KT})$ regret bound given above is desirable in terms of the dependence on $T$, the dependence on $K=\lvert\mathcal{Y}\rvert$ is undesirable for general structured prediction.
+From the convexity of $S_{t}$ and the unbiasedness of $\widehat{\bm{G}}_{t}$, we have $\mathbb{E}[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}]\leq\mathbb{E}[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}].$ By applying Lemma 3.3 and Jensen’s inequality, this term can be further upper bounded by $\mathbb{E}[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}]\leq\sqrt{2}B\sqrt{\mathbb{E}[\sum_{t=1}^{T}{{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}}]}\leq B\sqrt{\frac{2bK}{q}\mathbb{E}[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}]},$ where we used $\mathbb{E}[\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}),$ which follows from $p_{t}(\bm{y})\geq K/q$ and (1).
+From Assumption 3.2, it holds that $\mathbb{E}[\mathcal{R}_{T}]\leq\mathbb{E}[\sum_{t=1}^{T}{\left((1-a)S_{t}(\bm{W}_{t})-S_{t}(\bm{U})\right)}]+qT\leq B\sqrt{\frac{2bK}{q}\mathbb{E}[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}]}-a\mathbb{E}[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}]+qT\leq\frac{bB^{2}K}{2aq}+qT$, where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
+Plugging $q=B\sqrt{K/T}$ into the last inequality completes the proof.∎
+We can also prove a surrogate regret bound of $O(\sqrt{KT\log(1/\delta)}+\log(1/\delta))$, which holds with probabilty $1-\delta$.
+The precise statement and proof are provided in Section D.2.
+To prove this high-probability bound, we follow the analysis of Theorem 3.5 and use Bernstein’s inequality.
+To address the challenges posed by the randomness introduced by bandit feedback, we adopt an approach similar to that used in [39], and arguably, we have simplified their analysis.
+3.4 Algorithm based on pseudo-inverse matrix estimator with $O(T^{2/3})$ regret
+While the surrogate regret bound of $O(\sqrt{KT})$ achieves the presumably tight dependence on $T$, the dependence on $K=\lvert\mathcal{Y}\rvert$ is undesirable for general structured prediction.
 In fact, we have $K=\binom{d}{m}$ in multilabel classification with $m$ correct labels and $K=m!$ in ranking with $m$ items.
-To address this issue, we present an algorithm that significantly improves the dependence on $K$ when the target loss function belongs to a special class of SELF satisfying the following assumptions:
+To address this issue, we present an algorithm that significantly improves the dependence on $K$ when the target loss function belongs to a special class of SELF (2) with the following additional assumptions:
 Assumption 3.6.
-(i) $\bm{V}$ is invertible, and $\bm{b}$ and $c(\cdot)$ are known and non-negative.
+(i) $\bm{V}$ is known and invertible, and $\bm{b}$ and $c(\cdot)$ are known and non-negative.
 (ii) Let $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where $\mu$ is the uniform distribution over $\mathcal{Y}$.
 At least one of the following two conditions holds: (ii-a) $\bm{Q}$ is invertible, or (ii-b) for any $\bm{y}\in\mathcal{Y}$, $\bm{V}\bm{y}$ lies in the linear subspace spanned by vectors in $\mathcal{Y}$.
-(iii) For some $\omega>0$, it holds that
-$\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)\leq\omega.$
-The first condition is true in the examples in Section 2.4, assuming that the number of correct labels is fixed in multilabel classification.
-The second one is satisfied if $\mathcal{Y}$ consists of $d$ linearly independent vectors or $\bm{V}$ is proportional to the identity matrix; either is true in those examples.
-It is also not difficult to derive bounds on $\omega$ in those examples.
+(iii) For some $\omega>0$, it holds that $\operatorname*{tr}(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top})\leq\omega.$ (iv) For any $\widehat{\bm{y}}_{t}\in\mathcal{Y}$ and $\bm{y}_{t}\in\mathcal{Y}$, it holds that $|\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle|\leq 1$.
+The first and fourth conditions are true in the examples in Section 2.4, assuming that the number of correct labels, $m$, is fixed in multilabel classification.
+(While the fourth condition does not hold when $m>d/2$ in multilabel classification, we can flip 00 and $1$ in the labels to redefine the problem that satisfies the condition.) The second one holds if $\mathcal{Y}$ consists of $d$ linearly independent vectors or $\bm{V}$ is proportional to the identity matrix; either is true in the examples.
+Also, deriving reasonable bounds on $\omega$ in those examples is not difficult; see also Section D.4 for details.
 Algorithm based on pseudo-inverse matrix estimator
-As in the case of Section 3.3, we begin by providing a method to estimate the gradient.
-Define $\bm{P}_{t}\coloneqq\mathbb{E}_{\bm{y}\sim p_{t}}[\bm{y}\bm{y}^{\top}]$.
-Then, we define the estimator $\widetilde{\bm{y}}_{t}$ of $\bm{y}_{t}$ by
-$\widetilde{\bm{y}}_{t}\coloneqq{\bm{V}}^{-1}\bm{P}_{t}^{+}\widehat{\bm{y}}_{t}\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle,$
-where $\bm{P}_{t}^{+}$ is the Moore–Penrose pseudo-inverse matrix of $\bm{P}_{t}$.
-It is important to note that, given that $\bm{b}$ and $c(\cdot)$ are known, $\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle=L_{t}(\widehat{\bm{y}}_{t})-\langle\widehat{\bm{y}}_{t},\bm{b}\rangle-c(\bm{y}_{t})$ can be computed.
-Note that $\widetilde{\bm{y}}_{t}$ is unbiased, i.e., $\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]=\bm{y}_{t}$ from the first requirement of Assumption 3.7.
-Using this $\widetilde{\bm{y}}_{t}$, we define the gradient estimator $\widetilde{\bm{G}}_{t}$ by
-$\widetilde{\bm{G}}_{t}\coloneqq\left(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t}\right)\bm{x}_{t}^{\top},$ (5)
-whose expectation is $\mathbb{E}\big{[}\widetilde{\bm{G}}_{t}\big{]}=\bm{G}_{t}.$ Our estimator is based on the estimators used in adversarial linear bandits and adversarial combinatorial full-bandits (Dani et al., 2007; Abernethy et al., 2008; Cesa-Bianchi & Lugosi, 2012).
-We use RDUE with $q=\left(\frac{4\omega B^{2}C_{x}^{2}}{T}\right)^{1/3}$ as the decoding function (assuming $T\geq 4\omega B^{2}C_{x}^{2}$ for simplicity).
-For updating $\bm{W}_{t}$, we employ the adaptive OGD in Section 3.2 as $\mathsf{ALG}$ with the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
+As with Section 3.3, we consider estimating the gradient.
+Let $\bm{P}_{t}\coloneqq\mathbb{E}_{\bm{y}\sim p_{t}}[\bm{y}\bm{y}^{\top}]$ and define the estimator $\widetilde{\bm{y}}_{t}$ of $\bm{y}_{t}$ by $\widetilde{\bm{y}}_{t}\coloneqq{\bm{V}}^{-1}\bm{P}_{t}^{+}\widehat{\bm{y}}_{t}\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle,$ where $\bm{P}_{t}^{+}$ is the Moore–Penrose pseudo-inverse matrix of $\bm{P}_{t}$.
+Note that, given that $\bm{b}$ and $c(\cdot)$ are known, we can compute $\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle=L_{t}(\widehat{\bm{y}}_{t})-\langle\widehat{\bm{y}}_{t},\bm{b}\rangle-c(\bm{y}_{t})$.
+Importantly, $\widetilde{\bm{y}}_{t}$ is unbiased, i.e., $\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]=\bm{y}_{t}$ from the second requirement of Assumption 3.6.
+By using this $\widetilde{\bm{y}}_{t}$, we define the pseudo-inverse matrix estimator $\widetilde{\bm{G}}_{t}$ by $\widetilde{\bm{G}}_{t}\coloneqq\left(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t}\right)\bm{x}_{t}^{\top},$ which is also unbiased, i.e., $\mathbb{E}\big{[}\widetilde{\bm{G}}_{t}\big{]}=\bm{G}_{t}.$ Our estimator is inspired by those used in adversarial linear bandits and adversarial combinatorial full-bandits [14, 1, 8].
+We use RDUE with $q=({4\omega B^{2}C_{x}^{2}}/{T})^{1/3}$ as the decoding function (assuming $T\geq 4\omega B^{2}C_{x}^{2}$ for simplicity).
+To update $\bm{W}_{t}$, we use OGD in Section 3.2 as $\mathsf{ALG}$ with $\eta_{t}={B}/{\sqrt{2\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{i}\|_{\mathrm{F}}^{2}}}.$
 Regret bounds
-The above algorithm achieves the following regret bound, which does not directly depend on $K$:
+This algorithm achieves the following surrogate regret bound independent of $K$:
 Theorem 3.7.
-The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\big{(}\omega^{1/3}\left(BC_{x}T\right)^{2/3}\big{)}.$
+The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O\big{(}\omega^{1/3}T^{2/3}\big{)}.$
 The proof can be found in Section D.3.
-By using the estimator based on the pseudo-inverse matrix, we can upper bound the second moment of the gradient estimator $\widetilde{\bm{G}}_{t}$ without $K$, which allows us to establish the improved regret bound that does not explicitly depend on $K$.
-The regret bound in Theorem 3.8 yields the different bounds on each problem setup as follows:
+Note that we leverage the structure of OSP when using the pseudo-inverse matrix estimator, which largely differs from the existing approaches to surrogate regret analysis for online classification and OSP [38, 39, 35].
+With the pseudo-inverse matrix estimator, we can upper bound the second moment of the gradient estimator $\widetilde{\bm{G}}_{t}$ without $K$, which allows for the surrogate regret bound that does not explicitly involves $K$.
+This is in contrast to the inverse-weighted gradient estimator in Section 3.3.
+The inverse-weighted gradient estimator involves division by $p_{t}$, whose lower bound comes from uniform exploration on $\mathcal{Y}$; consequently, its upper bound depends on $K=|\mathcal{Y}|$.
+In other words, the above pseudo-inverse matrix estimator offers an alternative way to obtain an unbiased gradient estimator while eschewing uniform exploration on $\mathcal{Y}$.
+However, this comes at the price of a somewhat looser bound on the second moment, which increases the dependence on $T$.
+By using Theorem 3.7 as a template, we can derive specific bounds for each problem as follows:
 Corollary 3.8.
-The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((BC_{x}dT)^{2/3}\right)$ in multiclass classification with the 0-1 loss, $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((d^{5}/m(d-m))^{1/3}(BC_{x}T)^{2/3}\right)$ in multilabel classification with $m$ correct labels and the Hamming loss, and $\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(m^{5/3}(BC_{x}T)^{2/3}\right)$ in ranking with the number of items $m$ and the Hamming loss.
+The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]\leq O(d^{2/3}T^{2/3})$ in multiclass classification with the 0-1 loss, $\mathbb{E}[\mathcal{R}_{T}]\leq O\big{(}(d^{5}/m(d-m))^{1/3}T^{2/3}\big{)}$ in multilabel classification with $m$ correct labels and the Hamming loss, and $\mathbb{E}[\mathcal{R}_{T}]\leq O(m^{5/3}T^{2/3})$ in ranking with the number of items $m$ and the Hamming loss.
 The proof of Corollary 3.8 is deferred to Section D.4.
-The bound for multilabel classification with $m$ correct labels significantly improved the $O(\sqrt{KT})$ bound in Section 3.3, which has a dependency of $\sqrt{\binom{d}{m}}$, and the bound for ranking significantly improved the $O(\sqrt{KT})$ bound in Section 3.3, which has a dependency of $\sqrt{m!}$.
+The bound for multilabel classification with $m$ correct labels can be significantly better than the $O(\sqrt{KT})$ bound in Section 3.3 since $K=\sqrt{\binom{d}{m}}$; similarly, the bound for ranking can be much better than the $O(\sqrt{KT})$ bound since $K=\sqrt{m!}$.
 4 Delayed full-information feedback
-This section discusses online structured prediction with delayed full-information feedback and provides an algorithm that achieves a surrogate regret bound of $O(D^{2/3}T^{1/3})$, a better bound than $O(\sqrt{DT})$ that can be achieved with a standard OCO algorithm under delayed feedback (Joulani et al., 2013).
-Below, we make the following assumption.
+This section discusses OSP with fixed-delay full-information feedback and provides an algorithm that achieves a surrogate regret bound of $O(\min\{D^{2},D^{2/3}T^{1/3}\})$, a better bound than $O(\sqrt{DT})$ that can be achieved with a standard OCO algorithm under delayed feedback [22].
+Below, we make the following assumption based on the randomized decoding of [35].
 Assumption 4.1.
-There exists a constant $a\in(0,1)$ which satisfies
-$\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t}).$
+There exists a constant $a\in(0,1)$ that satisfies $\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t}).$
 From Lemma 2.4, if $\lambda>\frac{4\gamma}{\nu}$, this condition is satisfied with $a=1-\frac{4\gamma}{\lambda\nu}$ by using the randomized decoding.
@@
 Algorithm
-For updating $\bm{W}_{t}$, we employ the Optimistic Delayed Adaptive FTRL (ODAFTRL) algorithm proposed by Flaspohler et al.
-(2021).
-In ODAFTRL, given a gradient $\bm{G}_{t}$ at round $t$, $\bm{W}_{t}$ is updated as
-$\bm{W}_{t+1}=\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}\left\{\sum_{i=1}^{t-D}\left\langle\bm{G}_{i},\bm{W}\right\rangle+\frac{\lambda_{t}\|\bm{W}\|_{\mathrm{F}}^{2}}{2}\right\},$
-where $\lambda_{t}\geq 0$ is the regularization parameter.
-Due to space constraints, the details of the algorithm are provided in Section E.1.
-By updating $\lambda_{t}$ using an AdaHedge-type algorithm called AdaHedgeD, ODAFTRL achieves the following regret upper bound:
-Lemma 4.2 (Informal version of Flaspohler et al. 2021, Theorem 12).
-Consider the setting with delayed full-information feedback.
-Then, for any $\bm{U}\in\mathcal{W}$, ODAFTRL with the AdaHedgeD update of $\lambda_{t}$ achieves $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})\!-\!S_{t}(\bm{U}))}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}=O\Big{(}\sqrt{\sum_{t=1}^{T}{(\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\!+\!D\|\bm{G}_{t}\|_{\mathrm{F}})}}\Big{)}.$
+We employ the Optimistic Delayed Adaptive FTRL algorithm (ODAFTRL) [17] as $\mathsf{ALG}$, which we detail in Section E.1 for completeness.
+ODAFTRL computes the linear estimator by $\bm{W}_{t+1}\in\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}\left\{\sum_{i=1}^{t-D}\langle\bm{G}_{i},\bm{W}\rangle+\frac{\lambda_{t}\|\bm{W}\|_{\mathrm{F}}^{2}}{2}\right\},$ where $\lambda_{t}\geq 0$ is the regularization parameter.
+By updating $\lambda_{t}$ using an AdaHedge-type algorithm (AdaHedgeD), ODAFTRL achieves the following regret upper bound:
+Lemma 4.2 (Informal version of [17, Theorem 12]).
+Consider the delayed full-information feedback setting.
+Then, for any $\bm{U}\in\mathcal{W}$, ODAFTRL with the AdaHedgeD update of $\lambda_{t}$ achieves a regret bound of $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}=O\Big{(}\sqrt{\sum_{t=1}^{T}{(\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+D\|\bm{G}_{t}\|_{\mathrm{F}})}}\Big{)}.$
 Regret bounds and analysis
-The algorithm described above achieves the following bound:
+The above algorithm achieves the following surrogate regret bound:
 Theorem 4.3.
-The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(D^{2/3}T^{1/3}).$
-Here, we provide a proof sketch; the complete proof can be found in Section E.2.
+The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(\min\{D^{2},D^{2/3}T^{1/3}\}).$
+Notably, this bound is independent of $T$, and we expect that it is fairly close to tight, though no lower bound is known in the prior literature [27].
+Note that the proof ideas for deriving surrogate regret bounds in the non-delayed setting [39, 35] differ from those in the standard OCO and multi-armed bandits.
+This means that we cannot naively extend the analyses of the algorithms for delayed feedback in those settings to our case.
+Below is the proof sketch, and the complete proof is given in Section E.2.
 Proof sketch.
-Lemma 4.2 with $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$ in (2) and Cauchy–Schwarz yields $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}=O(\sqrt{S_{1:T}}+(D^{2}TS_{1:T})^{1/4})$, where $S_{1:T}=\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$.
-Hence, from Assumption 4.1, we have $\mathbb{E}[\mathcal{R}_{T}]\leq\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}=O(\sqrt{S_{1:T}}+(D^{2}TS_{1:T})^{1/4})-aS_{1:T}=O(D^{2/3}T^{1/3}),$ where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
+From Lemma 4.2 with $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$ in (1) and Cauchy–Schwarz ineqaulity, we have $\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}=O(D\sqrt{S_{1:T}})$, where $S_{1:T}=\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$.
+Thus, Assumption 4.1 implies $\mathbb{E}[\mathcal{R}_{T}]\leq\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}=O(D\sqrt{S_{1:T}})-aS_{1:T}=O(D^{2})$, where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
 ∎
-We can also prove the following high-probability bound; see Section E.3 for the proof:
+We can also prove a high-probability surrogate regret bound of $O(\min\{D^{2},D^{2/3}T^{1/3}\}+\log({1}/{\delta}))$, which holds with probability at least $1-\delta$. See Section E.3 for the proof.
+5 Delayed bandit feedback
+Given the results so far, it is natural to explore OSP with delayed bandit feedback.
+We construct algorithms for this setting by combining the theoretical developments from Sections 3 and 4.
+This section assumes Assumption 3.2, as in Section 3.
+5.1 Algorithm for bandit delayed feedback with $O(\sqrt{(K+D)T})$ regret
+We adopt RDUE with $q=B\sqrt{K/T}$ for decoding (assuming $T\geq B^{2}K$), the inverse-weighted gradient estimator $\widehat{\bm{G}}_{t}$, and ODAFTRL with AdaHedgeD as $\mathsf{ALG}$.
+Then, the following bound holds:
 Theorem 5.1.
-For any $\delta\in(0,1)$, with probability at least $1-\delta$, the above algorithm achieves $\mathcal{R}_{T}=O(\log({1}/{\delta})+D^{2/3}T^{1/3}).$
-5 Delayed Bandit Feedback
-Given the results so far, it is natural to explore online structured prediction with delayed bandit feedback.
-We construct algorithms for this setup by combining the theoretical developments from Section 3 and Section 4.
-This section assumes Assumption 3.2, as in Section 3.
-5.1 $O(\sqrt{DKT})$ Regret Algorithm
-We use RDUE with $q=B\sqrt{DK/T}$ as the decoding function (assuming $T\geq DB^{2}K$ for simplicity), the gradient estimator $\widetilde{\bm{G}}_{t}$ in (4), and ODAFTRL with the AdaHedgeD update as $\mathsf{ALG}$.
-This algorithm attains the following bound:
-Theorem 5.1.
-The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(\sqrt{DKT}).$
+The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(\sqrt{(K+D)T}).$
 The proof can be found in Section F.2.
-Due to the introduction of the delay, the regret bound worsens by a factor of $\sqrt{D}$ compared to the non-delayed case, which is natural when considering analyses of adversarial cases in delayed feedback (Joulani et al., 2013; Zimmert & Seldin, 2020; Manwani & Agarwal, 2023).
-5.2 $O(D^{1/3}T^{2/3})$ Regret Algorithm
-Here, we make the same assumptions on the target loss function as in Section 3.4.
+This upper bound incurs an additional $O(\sqrt{DT})$ compared to the non-delayed case.
+Whether this surrogate regret upper bound is tight or not remains open, as it is not easy to establish a corresponding lower bound.
+As mentioned above, an $\Omega(\sqrt{T})$ surrogate regret lower bound exists in the graph feedback setting [39].
+However, no such lower bound is known for the bandit setting, and constructing lower bounds that take the delay into account would be more difficult.
+5.2 Algorithm for bandit delayed feedback with $O(D^{1/3}T^{2/3})$ regret
 We provide an algorithm that improves the dependence on $K$ from Section 5.1.
-We use RDUE with $q=\left(\frac{\omega B^{2}C_{x}^{2}D}{T}\right)^{1/3}$ as the decoding function (assuming $T\geq\omega B^{2}C_{x}^{2}D$ for simplicity), the gradient estimator $\widetilde{\bm{G}}_{t}$ in (5), and ODAFTRL with the AdaHedgeD update as $\mathsf{ALG}$.
-This algorithm achieves the following bound:
+We make the same assumptions on the target loss function as Section 3.4.
+We use RDUE with $q=\left({\omega B^{2}C_{x}^{2}D}/{T}\right)^{1/3}$ for decoding (assuming $T\geq\omega B^{2}C_{x}^{2}D$), the pseudo-inverse matrix estimator $\widetilde{\bm{G}}_{t}$, and ODAFTRL with the AdaHedgeD update as $\mathsf{ALG}$.
+Then, the following bound holds:
 Theorem 5.2.
 The above algorithm achieves $\mathbb{E}[\mathcal{R}_{T}]=O(D^{1/3}T^{2/3}).$
-The proof can be found in Section F.3. Due to the introduction of the delay, the regret bound worsens by a factor of $D^{1/3}$ compared to the non-delayed bandit feedback case.
-6 Experiments
-Figure 1: A box plot of error rate of the MNIST experiment for multiclass classification with bandit feedback.
-This section presents numerical experiment results for online multiclass classification under bandit feedback.
-We compare three algorithms: Gaptron (Van der Hoeven, 2020) with logistic loss and hinge loss as surrogate losses, Gappletron (Van der Hoeven et al., 2021) with logistic loss as the surrogate loss, and our proposed algorithm from Section 5.1.
-The parameters for each algorithm are set based on their theoretical values.
-We use the MNIST dataset (LeCun et al., 2010), a dataset of digit images.
-The diameter $B$ of $\mathcal{W}$ is fixed at $10$.
-We repeated the experiment for 20 times, and the boxplot of the obtained misclassification rates is summarized in Figure 1.
-From Figure 1, we observe that our method achieves the lowest misclassification rate.
-Despite not being specialized for multiclass classification, our approach outperforms existing algorithms designed for multiclass classification on real data with $K=10$.
-Further experiments can be found in Appendix G, and related discussions are provided in Appendix C.
+The proof can be found in Section F.3.
+Due to the presence of the delay, the surrogate regret bound worsens by a factor of $D^{1/3}$ compared to the non-delayed bandit setting.
+Additionally, we present algorithms for the variable-delay setting, which we defer to Appendix G due to space limitation.
 6 Conclusion
-In this study, we developed multiple algorithms for online structured prediction under bandit and delayed feedback and analyzed their surrogate regret.
-Among several contributions, a noteworthy one is the algorithm for bandit feedback whose surrogate regret bound does not explicitly depend on the label set size $K$, which is achieved by leveraging the pseudo-inverse matrix estimator.
-Several interesting research directions remain.
-One is to investigate surrogate regret lower bounds under bandit feedback and delayed feedback.
-In the existing $\Omega(\sqrt{T})$ lower bound for bandit feedback, the optimal dependence on $d$ is unexplored.
-Investigating the tightness of the $O(D^{2/3}T^{1/3})$ upper bound for delayed and full information feedback is also an important question.
-Another key direction is extending our results to settings where the delay time is not constant in delayed feedback, which is crucial for some practical applications.
+We have developed multiple algorithms for OSP under bandit and delayed feedback and analyzed their surrogate regret.
+Among these contributions, of particular note is the algorithm for bandit feedback whose surrogate regret bound does not explicitly depend on the output set size $K$, achieved by leveraging the pseudo-inverse matrix estimator.
+Let us discuss limitations regarding tightness.
+As for the bandit setting, the lower bound of $\Omega(\sqrt{T})$ in the graph-feedback setting [39] suggests that a similar lower bound likely exists, but this has not yet been proven, leaving the tightness of our upper bounds open.
+In the delayed full-information setting, our $O(\min\{D^{2},D^{2/3}T^{1/3}\})$ upper bound is independent of $T$, and its gap to the lower bound, if any, is considered small; however, it remains open.
+In the delayed bandit setting, although the dependence on $D$ is less critical than that on $T$, the optimal dependence on $D$ is to be explored in future work.
 Acknowledgments
-TT is supported by JST ACT-X Grant Number JPMJAX210E and JSPS KAKENHI Grant Number JP24K23852, SS is supported by supported by JST ERATO Grant Number JPMJER1903, and KY is supported by JSPS KAKENHI Grant Number JP24H00703.
+TT is supported by JST ACT-X Grant Number JPMJAX210E and JSPS KAKENHI Grant Number JP24K23852, SS was supported by JST ERATO Grant Number JPMJER1903, and KY is supported by JSPS KAKENHI Grant Number JP24H00703.
 Appendix A Notation
@@
 $B\coloneqq\mathrm{diam}(\mathcal{W})$ diameter of $\mathcal{W}$
-$C_{x}\coloneqq\mathrm{diam}(\mathcal{X})$ diameter of input space $\mathcal{X}$
-$C_{y}\coloneqq\mathrm{diam}(\mathcal{Y})$ diameter of output space $\mathcal{Y}$
+$C_{x}=\max_{\bm{x}\in\mathcal{X}}\|\bm{x}\|_{2}$ maximum norm of input vectors in $\mathcal{X}$
+$C_{y}$ the maximum of the largest Euclidean norm of vectors in $\operatorname{conv}(\mathcal{Y})$ or the diameter of $\operatorname{conv}(\mathcal{Y})$
 $K\coloneqq|\mathcal{Y}|$ cardinality of $\mathcal{Y}$
-$D\in\mathbb{N}$ fixed delayed time
 $\mathsf{ALG}$ algorithm for updating linear estimators
@@
 $p_{t}(\bm{y})$ probability that $\bm{y}$ is chosen as $\widehat{\bm{y}}_{t}$ at time $t$
-$L:\mathcal{Y}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ target loss function
+$L\colon\mathcal{Y}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ target loss function
 $L_{t}(\widehat{\bm{y}}_{t})\coloneqq L(\widehat{\bm{y}}_{t};\bm{y}_{t})$ the value of target loss $L(\widehat{\bm{y}}_{t};\bm{y}_{t})$
-$S_{\Omega}:\mathbb{R}^{n}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ Fenchel–Young loss generated by $\Omega$
+$S_{\Omega}\colon\mathbb{R}^{n}\times\mathcal{Y}\xrightarrow{}\mathbb{R}_{\geq 0}$ Fenchel–Young loss generated by $\Omega$
 $S_{t}(\bm{W})\coloneqq S_{\Omega}(\bm{W}\bm{x}_{t};\bm{y}_{t})$ shorthand of surrogate loss $S_{\Omega}(\bm{W}\bm{x}_{t};\bm{y}_{t})$
-$L\coloneqq\max_{\bm{W}\in\mathcal{W}}\|\nabla S_{t}(\bm{W})\|_{\mathrm{F}}$ upper bound of the gradient
+$\mathcal{R}_{T}\coloneqq\sum_{t=1}^{T}{(L_{t}(\widehat{\bm{y}}_{t})-S_{t}(\bm{U}))}$ surrogate regret
 $\bm{G}_{t}\coloneqq\nabla S_{t}(\bm{W}_{t})$ gradient of surrogate loss
-$\mathcal{R}_{T}\coloneqq\sum_{t=1}^{T}{(L_{t}(\widehat{\bm{y}}_{t})-S_{t}(\bm{U}))}$ surrogate regret
+${\widehat{\bm{G}}_{t}}$ inverse weighted estimator
+${\widetilde{\bm{G}}_{t}}$ pseudo-inverse matrix estimator
 $\mu$ uniform distribution over $\mathcal{Y}$
@@
 $\mathbb{E}_{t}\left[{\cdot}\right]$ conditional expectation given $\widehat{\bm{y}}_{1},\dots,\widehat{\bm{y}}_{t-1}$
+$D\in\mathbb{N}$ fixed delay time
+$\tau_{t}$ variable delay time
+$\tau_{\ast}=\max_{t}\tau_{t}$ maximum value of delay time
+$\rho(t)$ the time step of the $t$th feedback from SOLID to BASE
+$\widetilde{\tau}_{t}=t-1-\sum_{s=1}^{\rho(t)-1}\mathds{1}[s+\tau_{s}<\rho(t)]$ the number of feedback from SOLID to BASE pending during the $t$th feedback
 Appendix B Additional related work
-We discuss additional related work that could not be included above.
+We discuss additional related work that could not be included in the main text.
 Structured prediction
-Before the introduction of the Fenchel–Young loss framework, Niculae et al.
-(2018) proposed SparseMAP, which used the squared $\ell_{2}$-norm regularization.
+Before the development of the Fenchel–Young loss framework, Niculae et al.
+[30] proposed SparseMAP, which used the squared $\ell_{2}$-norm regularization.
 The Fenchel–Young loss, described in Section 2.2, is built upon the idea of SparseMAP.
 The Structure Encoding Loss Function (SELF) was introduced by Ciliberto et al.
-(2016, 2020) to analyze the relationship between surrogate and target losses, a concept known as Fisher consistency.
-For more extensive literature, we refer the reader to Sakaue et al.
-(2024, Appendix A).
+[10, 11] to analyze the relationship between surrogate and target losses, a concept known as Fisher consistency.
+For a more extensive literature review, we refer the reader to Sakaue et al.
+[35, Appendix A].
 Online classification with full and bandit feedback
-In the full information setup, PERCEPTRON is one of the most representative algorithms for binary classification (Rosenblatt, 1958), and the multiclass setting has also been extensively studied (Crammer & Singer, 2003; Fink et al., 2006).
+In the full-information setting, PERCEPTRON is one of the most representative algorithms for binary classification [34], and the multiclass setting has also been extensively studied [13, 16].
 Online logistic regression is another relevant research stream, with Foster et al.
-(2018) being a particularly representative study.
-The study of the bandit setup was initiated by Kakade et al.
-(2008), and it has since been extensively explored in subsequent research (Hazan & Kale, 2011; Beygelzimer et al., 2017; Foster et al., 2018).
+[18] being a particularly representative study.
+The study of the bandit setting was initiated by Kakade et al.
+[24], and it has since been extensively explored in subsequent research [20, 4, 18].
 However, to the best of our knowledge, no prior work has addressed general structured prediction under bandit feedback.
-A most related study is the work by Gentile & Orabona (2014b), which investigated online multilabel classification and ranking.
-However, their setting assumes access to feedback of the form $\{\mathds{1}[\bm{y}_{t,i}\neq\widehat{\bm{y}}_{t,i}]\}_{i}$, which is more informative than bandit feedback and differs from our setup.
-Van der Hoeven (2020) explicitly introduced the surrogate regret in the context of online multiclass classification.
-This study has been extended to the setting where observations are determined by a directed graph (Van der Hoeven et al., 2021) and to structured prediction scenarios (Sakaue et al., 2024).
-For a more extensive overview of the literature on online classification, we refer the reader to Van der Hoeven (2020).
+One of the most relevant studies is the work by Gentile and Orabona [19], who investigated online multilabel classification and ranking.
+However, their setting assumes access to feedback of the form $\{\mathds{1}[\bm{y}_{t,i}\neq\widehat{\bm{y}}_{t,i}]\}_{i}$, which is more informative than bandit feedback and differs from our setting.
+van der Hoeven [38] introduced the surrogate regret in the context of online multiclass classification.
+This study has been extended to the setting where observations are determined by a directed graph [39] and to structured prediction [35].
+For a more extensive overview of the literature on online classification, we refer the reader to van der Hoeven [38].
 Delayed feedback
-The study of delayed feedback began with Weinberger & Ordentlich (2002).
-Since then, it has been extensively explored in various online learning settings, primarily in the full information setup of online convex optimization (Mesterharm, 2005; Joulani et al., 2013, 2016; Flaspohler et al., 2021).
-Algorithms for delayed bandit feedback have been studied mainly in the context of multi-armed bandits and their variants (Cesa-Bianchi et al., 2016; Zimmert & Seldin, 2020; Ito et al., 2020; Masoudian et al., 2022; Van der Hoeven et al., 2023).
-In the context of online classification, research considering delay is scarce; the only work is Manwani & Agarwal (2023) to our knowledge.
+The study of delayed feedback was initiated by Weinberger and Ordentlich [41].
+Since then, it has been extensively explored in various online learning settings, primarily in the full-information setting of online convex optimization [29, 22, 23, 17].
+Algorithms for delayed bandit feedback have been studied mainly in the context of multi-armed bandits and their variants [9, 42, 21, 28, 40].
+In online classification, research considering delay is scarce; the only work is that of Manwani and Agarwal [27] to our knowledge.
+There are several differences between their work and ours.
+Among them, a key distinction is that their study focuses on multiclass classification, whereas we address the more general OSP.
 Appendix C Discussion on the difference in surrogate loss functions
-As in (1), the surrogate regret $\mathcal{R}_{T}$ is defined by $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}$, which means the choice of the surrogate loss $S$ affects the bound on the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$.
+As in Section 1, the surrogate regret, $\mathcal{R}_{T}$, is defined by $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}=\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}+\mathcal{R}_{T}$, which means the choice of the surrogate loss function, $S$, affects the bound on the cumulative loss $\sum_{t=1}^{T}{L(\widehat{\bm{y}}_{t};\bm{y}_{t})}$.
 van der Hoeven et al.
-(2021, Theorem 1), which applies to a more general setup than bandit feedback, implies $\mathcal{R}_{T}=O(K\sqrt{T})$ for the bandit setup with $S$ being a logistic loss defined with the base-$K$ logarithm.
+[39, Theorem 1], which applies to a more general setting than bandit feedback, implies $\mathcal{R}_{T}=O(K\sqrt{T})$ for the bandit setting with $S$ being a logistic loss defined with the base-$K$ logarithm.
 On the other hand, our bound of $\mathcal{R}_{T}=O(\sqrt{KT})$ applies to the logistic loss $S$ defined with the base-$2$ logarithm.
 As a result, while our bound on $\mathcal{R}_{T}$ is better, the $\sum_{t=1}^{T}{S(\bm{U}\bm{x}_{t};\bm{y}_{t})}$ term can be worse; this is why we cannot directly compare our $O(\sqrt{KT})$ bound with the $O(K\sqrt{T})$ bound in van der Hoeven et al.
-(2021, Theorem 1).
+[39, Theorem 1].
 We may use the decoding procedure in van der Hoeven et al.
-(2021), instead of RDUE, to recover their bound that applies to the base-$K$ logistic loss.
+[39], instead of RDUE, to recover their bound that applies to the base-$K$ logistic loss.
 It should be noted that their method is specific to multiclass classification; naively extending their method to structured prediction formulated as $|\mathcal{Y}|$-class classification results in the undesirable dependence on $K=|\mathcal{Y}|$, as is also discussed in Sakaue et al.
-(2024).
+[35].
 By contrast, our pseudo-inverse estimator, combined with RDUE, can rule out the explicit dependence on $K$, at the cost of the increase from $\sqrt{T}$ to $T^{2/3}$.
-Appendix D Omitted Details of Section 3
+Appendix D Details omitted from Section 3
 This section provides the omitted details of Section 3.
 D.1 Concentration inequality
-To prove high probability regret bounds, we use the following concentration inequality.
-Lemma D.1 (Bernstein’s inequality, e.g., Cesa-Bianchi & Lugosi 2006, Lemma A.8).
+To prove high probability regret bounds, we will use the following concentration inequality.
+Lemma D.1 (Bernstein’s inequality, e.g., [7, Lemma A.8]).
 Let $Z_{1},\ldots,Z_{T}$ be a martingale difference sequence and $\delta\in(0,1)$.
@@
 $\sum_{t=1}^{T}{Z_{t}}\leq\sqrt{2v\log\frac{1}{\delta}}+\frac{\sqrt{2}}{3}a\log\frac{1}{\delta}.$
-D.2 Proof of Theorem 3.6
-Here, we provide the proof of Theorem 3.6.
+D.2 Proof of high probability bound
+Here, we provide the proof of a high probability bound.
 Hereafter, we let $S_{\max}=\max_{\bm{W}\in\mathcal{W}}S_{t}(\bm{W})$ and $\widehat{S}_{t}(\bm{W})=v_{t}S_{t}(\bm{W})=\frac{\mathds{1}[\bm{y}_{t}=\widehat{\bm{y}}_{t}]}{p_{t}(\widehat{\bm{y}}_{t})}S_{t}(\bm{W})$.
-The following theorem is the formal version of Theorem 3.6:
-Theorem D.2 (Formal version of Theorem 3.6).
-Consider the bandit and non-delayed setup. Let
+The following theorem is the formal version of the high probability bound under the bandit feedback:
+Theorem D.2.
+Consider the bandit and non-delayed setting. Let
 $\mathcal{C}=\left(\frac{3}{2(a+\xi-1)}+1\right)KS_{\max}\log(2/\delta)+\frac{B^{2}Kb}{2(1-\xi)}.$
@@
 and the second term is bounded as
-$\sum_{t=1}^{T}\left(\widehat{S}_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{U})\right)$ $\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}=\sqrt{2}B\sqrt{\sum_{t=1}^{T}v_{t}^{2}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}$
+$\sum_{t=1}^{T}\left(\widehat{S}_{t}(\bm{W}_{t})-\widehat{S}_{t}(\bm{U})\right)$ $\leq\sqrt{2}B\sqrt{\sum_{t=1}^{T}\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}=\sqrt{2}B\sqrt{\sum_{t=1}^{T}v_{t}^{2}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}$
 $\leq\sqrt{2}B\sqrt{b\sum_{t=1}^{T}v_{t}^{2}S_{t}(\bm{W}_{t})}\leq\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})},$
@@
 $\mathbb{E}_{t}[(v_{t}r_{t}-r_{t})^{2}]\leq\mathbb{E}_{t}[(v_{t}r_{t})^{2}]\leq\frac{KS_{\max}}{q}\lvert r_{t}\rvert\leq\frac{KS_{\max}}{q}\left(S_{t}(\bm{U})+S_{t}(\bm{W}_{t})\right).$
-Hence from Bernstein’s inequality in Lemma D.1, for any $\delta^{\prime\prime}\in(0,1)$, with probability at least $1-\delta^{\prime\prime}$ we have
+Hence, from Bernstein’s inequality in Lemma D.1, for any $\delta^{\prime\prime}\in(0,1)$, with probability at least $1-\delta^{\prime\prime}$ we have
 $\sum_{t=1}^{T}(v_{t}r_{t}-r_{t})\leq\sqrt{3\log(1/\delta^{\prime\prime})\sum_{t=1}^{T}\frac{KS_{\max}}{q}(S_{t}(\bm{U})+S_{t}(\bm{W}_{t}))}+\frac{\sqrt{2}KS_{\max}}{3q}\log(1/\delta^{\prime\prime}).$ (5)
+Below, we proceed by case analysis.
 When $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$. We first consider the case of $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$. From Lemma D.3, we have
-$\sum_{t=1}^{T}\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-qT\leq\sum_{t=1}^{T}v_{t}S_{t}(\bm{U})+\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})-v_{t}S_{t}(\bm{W}_{t})\right)+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
-$=\sum_{t=1}^{T}v_{t}\underbrace{\left(S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})\right)}_{=r_{t}}-(1-\xi)\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
+$\sum_{t=1}^{T}\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-qT$
+$\leq\sum_{t=1}^{T}v_{t}S_{t}(\bm{U})+\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})-v_{t}S_{t}(\bm{W}_{t})\right)+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
+$=\sum_{t=1}^{T}v_{t}\underbrace{\left(S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})\right)}_{=r_{t}}-(1-\xi)\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})$
+$\qquad+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\sqrt{2}B\sqrt{\frac{bK}{q}\sum_{t=1}^{T}v_{t}S_{t}(\bm{W}_{t})}$
 $\leq\sum_{t=1}^{T}v_{t}r_{t}+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)},$
@@
 $\qquad+\frac{\sqrt{2}KS_{\max}}{3q}\log(1/\delta^{\prime\prime})+(1-a)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)},$
-where we recall that $r_{t}=S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})$. Rearranging the last inequality and using $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$ give
+where we recall that $r_{t}=S_{t}(\bm{U})-\xi S_{t}(\bm{W}_{t})$. Rearranging the last inequality and using $\sum_{t=1}^{T}S_{t}(\bm{U})\leq\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$ yield
 $\sum_{t=1}^{T}\left(\mathbb{E}_{t}\left[L_{t}(\widehat{\bm{y}}_{t})\right]-S_{t}(\bm{U})\right)$ $\leq qT+\sqrt{6\log(1/\delta^{\prime\prime})\sum_{t=1}^{T}\frac{KS_{\max}}{q}S_{t}(\bm{W}_{t})}+\frac{\sqrt{2}KS_{\max}}{3q}\log(1/\delta^{\prime\prime})$
@@
 Then, plugging (4) and the last inequality in (3), with probability at least $1-\delta$, we obtain
-$\mathcal{R}_{T}$ $\leq\sqrt{2\log(2/\delta)\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})+q\right)}+\frac{\sqrt{2}}{3}\log(2/\delta)+qT+\sqrt{6\log(2/\delta)\sum_{t=1}^{T}\frac{KS_{\max}}{q}S_{t}(\bm{W}_{t})}$
-$\qquad+\frac{\sqrt{2}KS_{\max}}{3q}\log(2/\delta)+(1-a-\xi)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)}$
+$\mathcal{R}_{T}$ $\leq\sqrt{2\log(2/\delta)\sum_{t=1}^{T}\left((1-a)S_{t}(\bm{W}_{t})+q\right)}+\frac{\sqrt{2}}{3}\log(2/\delta)+qT$
+$\qquad+\sqrt{6\log(2/\delta)\sum_{t=1}^{T}\frac{KS_{\max}}{q}S_{t}(\bm{W}_{t})}+\frac{\sqrt{2}KS_{\max}}{3q}\log(2/\delta)$
+$\qquad+(1-a-\xi)\sum_{t=1}^{T}S_{t}(\bm{W}_{t})+\frac{B^{2}Kb}{2q(1-\xi)}$
 $\leq\frac{1}{2(a+\xi-1)}\left((1-a)+\frac{3KS_{\max}}{q}\right)\log(2/\delta)+\sqrt{2qT\log(2/\delta)}+\frac{\sqrt{2}}{3}\log(2/\delta)+qT$
@@
 where the first inequality follows from (4) and $\sum_{t=1}^{T}S_{t}(\bm{U})>\sum_{t=1}^{T}S_{t}(\bm{W}_{t})$, and the second inequality follows from Assumption 3.2, the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
-Substituting $q=\sqrt{\mathcal{C}/2}$ and $\delta^{\prime}=\delta/2$ and in the last inequality, we obtain
+Substituting $q=\sqrt{\mathcal{C}/T}$ and $\delta^{\prime}=\delta/2$ in the last inequality, we obtain
 $\mathcal{R}_{T}\leq\frac{(1-a)\log(2/\delta)}{2a}+\sqrt{2\log(2/\delta)}(\mathcal{C}T)^{1/4}+\frac{\sqrt{2}}{3}\log(2/\delta)+\sqrt{\mathcal{C}T}.$
@@
 D.3 Proof of Theorem 3.7
-Here, we provide the proof of Theorem 3.8.
+Here, we provide the formal version and the proof of Theorem 3.7.
+Theorem D.4 (Formal version of Theorem 3.7).
+The algorithm in Section 3.4 achieves
+$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2^{5/3}\omega^{1/3}\left(BC_{x}T\right)^{2/3}.$
 We recall that $\bm{P}_{t}=\mathbb{E}_{t}\left[{\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}}\right]$.
@@
 $\bm{A}=\bm{R}\begin{pmatrix}O&O\\ O&\bm{\Lambda}\end{pmatrix}\bm{R}^{\top}\quad\mbox{and}\quad\bm{B}=\bm{R}\begin{pmatrix}O&O\\ O&\bm{B}^{\prime}\end{pmatrix}\bm{R}^{\top}.$
-Then,
+Then, we have
 $\bm{A}^{+}=\bm{R}\begin{pmatrix}O&O\\ O&\bm{\Lambda}^{-1}\end{pmatrix}\bm{R}^{\top}\quad\mbox{and}\quad\bm{B}^{+}=\bm{R}\begin{pmatrix}O&O\\ O&{\bm{B}^{\prime}}^{-1}\end{pmatrix}\bm{R}^{\top}.$ (6)
@@
 ∎
-Using this lemma we prove a property of $\bm{P}_{t}$ and an upper bound of $\mathbb{E}_{t}\left[{\operatorname*{tr}\left(\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}\right)}\right]$.
+Using this lemma, we prove a property of $\bm{P}_{t}$ and an upper bound of $\mathbb{E}_{t}\left[{\operatorname*{tr}\left(\widehat{\bm{y}}_{t}\widehat{\bm{y}}_{t}^{\top}\right)}\right]$.
 In what follows, we use $\lambda_{\min}(\bm{A})$ to denote the minimum eigenvalue of a matrix $\bm{A}$.
 Lemma D.6.
-Suppose that $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where we recall $\mu$ is the uniform distribution over $\mathcal{Y}$.
+Suppose that $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$, where we recall that $\mu$ is the uniform distribution over $\mathcal{Y}$.
 Then, we have
@@
 $=\operatorname*{tr}\left({\bm{V}}^{-1}\bm{P}^{+}_{t}\left({\bm{V}}^{-1}\right)^{\top}\right),$
-where the first inequality follows from $\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle=L_{t}(\widehat{\bm{y}}_{t})-\langle\widehat{\bm{y}}_{t},\bm{b}\rangle-c(\bm{y}_{t})\leq L_{t}(\widehat{\bm{y}}_{t})\leq 1$ since $\bm{b}\geq 0$ and $c(\cdot)$ is non-negative and the last equality follows from $\bm{P}^{+}_{t}\bm{P}_{t}\bm{P}^{+}_{t}=\bm{P}^{+}_{t}$.
-Hence,
+where the first inequality follows from $-1\leq\langle\widehat{\bm{y}}_{t},\bm{V}\bm{y}_{t}\rangle\leq 1$ and the last equality follows from $\bm{P}^{+}_{t}\bm{P}_{t}\bm{P}^{+}_{t}=\bm{P}^{+}_{t}$.
+The right-hand side is bounded as follows:
 $\operatorname*{tr}\left({\bm{V}}^{-1}\bm{P}^{+}_{t}\left({\bm{V}}^{-1}\right)^{\top}\right)$ $=\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}{\bm{V}}^{-1}\bm{P}^{+}_{t}\left({\bm{V}}^{-1}\right)^{\top}\bm{\mathrm{e}}_{i}\leq\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}{\bm{V}}^{-1}\left(q\bm{Q}\right)^{+}\left({\bm{V}}^{-1}\right)^{\top}\bm{\mathrm{e}}_{i}$
@@
 $\leq 2\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2C_{x}^{2}\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}^{2},$
-where we recall $C_{x}=\mathrm{diam}(\mathcal{X})$. From this inequality,
-$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$ $\leq 2\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2C_{x}^{2}\mathbb{E}_{t}\left[{\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\leq 2bS_{t}(\bm{W}_{t})+2C_{x}^{2}\left(\|\bm{y}_{t}\|_{2}^{2}-2\bm{y}_{t}^{\top}\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]+\mathbb{E}_{t}\left[{\|\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\right)$
+where we recall $C_{x}=\mathrm{diam}(\mathcal{X})$. From this inequality, we obtain
+$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]$ $\leq 2\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2C_{x}^{2}\mathbb{E}_{t}\left[{\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]$
+$\leq 2bS_{t}(\bm{W}_{t})+2C_{x}^{2}\left(\|\bm{y}_{t}\|_{2}^{2}-2\bm{y}_{t}^{\top}\mathbb{E}_{t}\left[{\widetilde{\bm{y}}_{t}}\right]+\mathbb{E}_{t}\left[{\|\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\right)$
 $=2bS_{t}(\bm{W}_{t})+2C_{x}^{2}\left(\|\bm{y}_{t}\|_{2}^{2}-2\|\bm{y}_{t}\|_{2}^{2}+\mathbb{E}_{t}\left[{\|\widetilde{\bm{y}}_{t}\|_{2}^{2}}\right]\right)$
@@
 ∎
-Finally, we are ready to prove Theorem 3.8.
-Proof of Theorem 3.8.
+Finally, we are ready to prove Theorem D.4.
+Proof of Theorem D.4.
 From Assumption 3.2, we have
@@
 $\leq 2B\sqrt{b\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BC_{x}\sqrt{\omega/q},$
-where the first inequality follows from Lemma 3.3 and the last inequality follows from Lemma D.6 and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$.
-Therefore, by combining with the last inequality, we have
+where the first inequality follows from Lemma 3.3 and the last inequality follows from Lemma D.7 and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, we obtain
 $\mathbb{E}[\mathcal{R}_{T}]$ $\leq 2B\sqrt{b\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BC_{x}\sqrt{\omega/q}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$ (7)
@@
 D.4 Proof of Corollary 3.8
-Here, we derive the regret upper bounds provided by the algorithm established in Theorem 3.8 for online multiclass classification, online multilabel classification, and ranking.
+We derive the surrogate regret upper bounds provided by the algorithm established in Theorem D.4 for online multiclass classification, online multilabel classification, and ranking.
 Recall that we can achieve
-$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(\omega^{1/3}\left(BC_{x}T\right)^{2/3}\right),$ (11)
+$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2^{5/3}\omega^{1/3}\left(BC_{x}T\right)^{2/3},$ (8)
 where we recall that $\omega$ is defined as $\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)\leq\omega$ for $\bm{Q}=\mathbb{E}_{\bm{y}\sim\mu}[\bm{y}\bm{y}^{\top}]$.
-Note that when $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, then the matrix $\bm{Q}$ is invertible and $\lambda_{\min}(\bm{Q})>0$, and thus
-$\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)=\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\leq\frac{1}{\lambda_{\min}(\bm{Q})}\sum_{i=1}^{d}\|(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\|_{2}^{2}\leq\frac{1}{\lambda_{\min}(\bm{Q})}\|\bm{V}^{-1}\|_{\mathrm{F}}^{2}.$ (12)
-In each problem setting, this regret upper bound can be reduced to the following bounds:
+Note that when $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, then the matrix $\bm{Q}$ is invertible and $\lambda_{\min}(\bm{Q})>0$, and hence
+$\operatorname*{tr}\left(\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\right)$ $=\sum_{i=1}^{d}\bm{\mathrm{e}}_{i}^{\top}\bm{V}^{-1}\bm{Q}^{+}(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\leq\frac{1}{\lambda_{\min}(\bm{Q})}\sum_{i=1}^{d}\|(\bm{V}^{-1})^{\top}\bm{\mathrm{e}}_{i}\|_{2}^{2}$
+$\leq\frac{1}{\lambda_{\min}(\bm{Q})}\|\bm{V}^{-1}\|_{\mathrm{F}}^{2}.$ (9)
+Consequently, surrogate regret bounds for specific problems are obtained as follows:
 Multiclass classification with 0-1 loss
 We first consider multiclass classification with the 0-1 loss.
-Since $\bm{V}=\bm{1}\bm{1}^{\top}-\bm{I}$, we have $\|{\bm{V}}^{-1}\|_{\mathrm{F}}^{2}\leq d$ for $d\geq 2$.
+From $\bm{V}=\bm{1}\bm{1}^{\top}-\bm{I}$, we have $\|{\bm{V}}^{-1}\|_{\mathrm{F}}^{2}\leq d$ for $d\geq 2$.
 Recalling that $\mu$ is the uniform distribution over $\mathcal{Y}=\{\bm{\mathrm{e}}_{1},\ldots,\bm{\mathrm{e}}_{d}\}$, we have $\mathbb{E}_{\bm{y}\sim\mu}[(\bm{y}^{\top}\bm{x})^{2}]=\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}$ for any $\bm{x}\in\mathbb{R}^{d}$.
-Hence, $\lambda_{\min}(\bm{Q})=\min_{\|\bm{x}\|_{2}=1}\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]=\frac{1}{d}$, where the first equality is from Cesa-Bianchi & Lugosi (2012, Lemma 2).
-Since $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$ in this case, from (12) we can let $\omega=d/\lambda_{\min}(\bm{Q})=d^{2}$.
-Substituting these into our upper bound in (11), we obtain
-$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left((BC_{x}dT)^{2/3}\right).$
+Hence, we have $\lambda_{\min}(\bm{Q})=\min_{\|\bm{x}\|_{2}=1}\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]=\frac{1}{d}$, where the first equality follows from [8, Lemma 2].
+Since $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$ holds in this case, from (D.4), we can set $\omega=d/\lambda_{\min}(\bm{Q})=d^{2}$.
+Substituting these into our surrogate regret upper bound in (8), we obtain
+$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2^{5/3}\left(dBC_{x}T\right)^{2/3}.$
 Online multilabel classification with $m$ correct labels and the Hamming loss
@@
 $\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]=\frac{m}{d}\sum_{i=1}^{d}x_{i}^{2}+\frac{m^{2}}{d^{2}}\sum_{i\neq j}x_{i}x_{j}=\left(\frac{m}{d}\sum_{i=1}^{d}x_{i}\right)^{2}+\frac{m(d-m)}{d^{2}}\|\bm{x}\|_{2}^{2}\geq\frac{m(d-m)}{d^{2}}.$
-Hence, we have $\lambda_{\min}(\bm{Q})=\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]\geq\frac{m(d-m)}{d^{2}}$, where the equality is from Cesa-Bianchi & Lugosi (2012, Lemma 2).
-Since $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, from (12) we can choose $\omega=d^{3}/(4\lambda_{\min}(\bm{Q}))=\frac{d^{5}}{4m(d-m)}$.
-Therefore, our regret upper bound in (11) is reduced to
-$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(\left(\frac{B^{2}C_{x}^{2}d^{5}}{m(d-m)}\right)^{1/3}T^{2/3}\right).$
+Thus, $\lambda_{\min}(\bm{Q})=\min_{\|\bm{x}\|_{2}=1}\mathbb{E}_{\bm{y}\sim\mu}\left[(\bm{y}^{\top}\bm{x})^{2}\right]\geq\frac{m(d-m)}{d^{2}}$ holds, where the equality is from Cesa-Bianchi and Lugosi [8, Lemma 2].
+Since we have $\operatorname*{span}(\mathcal{Y})=\mathbb{R}^{d}$, from (D.4), we can set $\omega=\frac{d^{5}}{4m(d-m)}\geq\|\bm{V}^{-1}\|_{\mathrm{F}}^{2}/\lambda_{\min}(\bm{Q})$.
+Therefore, our surrogate regret upper bound in (8) is reduced to
+$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+2\left(\frac{d^{5}}{m(d-m)}\right)^{1/3}(BC_{x}T)^{2/3}.$
 Ranking with the Hamming loss and the number of items $m$
 We finally consider online ranking with the Hamming loss and the number of items $m$.
-From Cesa-Bianchi & Lugosi (2012, Proposition 4), the smallest positive eigenvalue is at least $1/m$.
+From Cesa-Bianchi and Lugosi [8, Proposition 4], the smallest positive eigenvalue is at least $1/m$.
 Hence, since $\bm{V}=-\frac{1}{m}\bm{I}$, we have
 $\operatorname*{tr}({\bm{V}}^{-1}\bm{Q}^{+}({\bm{V}}^{-1})^{\top})=m^{2}\operatorname*{tr}(\bm{Q}^{+})\leq m^{2}\sum_{i=1}^{\mathrm{rank}(\bm{Q}^{+})}m\leq m^{5},$
-where we used $\mathrm{rank}(\bm{Q}^{+})\leq d=m^{2}$, and this allows us to choose $\omega=m^{5}$. Substituting these values into our regret upper bound in (11) , we obtain
-$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+O\left(m^{5/3}(BC_{x}T)^{2/3}\right).$
-Appendix E Omitted Details of Section 4
+where we used $\mathrm{rank}(\bm{Q}^{+})\leq d=m^{2}$, and this allows us to choose $\omega=m^{5}$. Substituting these values into our surrogate regret upper bound in (8) , we obtain
+$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{bB^{2}}{a}+(2m)^{5/3}(BC_{x}T)^{2/3}.$
+Appendix E Details omitted from Section 4
 This section provides the proofs of the theorems in Section 4.
@@
 We provide a more detailed explanation of the Optimistic Delayed Adaptive FTRL (ODAFTRL) algorithm used for updating $\bm{W}_{t}$ in Section 4.
-We recall that ODAFTRL computes $\bm{W}_{t}$ by the following update rule:
+Recall that ODAFTRL computes $\bm{W}_{t}$ by the following update rule:
 $\bm{W}_{t+1}=\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}\left\{\sum_{i=1}^{t-D}\langle\bm{G}_{i},\bm{W}\rangle+\frac{\lambda_{t}\|\bm{W}\|_{\mathrm{F}}^{2}}{2}\right\},$ (10)
-where $\lambda_{t}\geq 0$ is a regularization parameter. The ODAFTRL algorithm, when using this parameter update called AdaHedgeD, satisfies the following lemma:
-Lemma E.1 (Flaspohler et al. 2021, Theorem 12).
-Fix $\alpha>0$. Let $S_{t}:\mathcal{W}\to\mathbb{R}$ be a convex function for each $t=1,\dots,T$. Suppose that we update $\lambda_{t}$ in (13) by the following AdaHedgeD update:
+where $\lambda_{t}\geq 0$ is the regularization parameter.
+Note that we use the notation $a_{1:t}=\sum_{i=1}^{t}a_{i}$ for simplicity in the following.
+The ODAFTRL algorithm, when using the parameter update called AdaHedgeD, satisfies the following lemma:
+Lemma E.1 ([17, Theorem 12]).
+Fix $\alpha>0$. Let $f_{t}\colon\mathcal{W}\to\mathbb{R}$ be a convex function for each $t=1,\dots,T$. Suppose that we update $\lambda_{t}$ in (10) by the following AdaHedgeD update:
 $\lambda_{t+1}=\frac{1}{\alpha}\sum_{s=1}^{t-D}\delta_{s},$
-$\delta_{t}$ $=\min\left\{F_{t+1}(\bm{W}_{t})-F_{t+1}(\bar{\bm{W}}_{t}),\langle\bm{G}_{t},\bm{W}_{t}-\bar{\bm{W}}_{t}\rangle,F_{t+1}(\widehat{\bm{W}}_{t})-F_{t+1}(\bar{\bm{W}}_{t})+\langle\bm{G}_{t},\bm{W}_{t}-\widehat{\bm{W}}_{t}\rangle\right\}_{+},$
+$\delta_{t}=\min\big{\{}F_{t+1}(\bm{W}_{t})-F_{t+1}(\bar{\bm{W}}_{t}),\langle\bm{G}_{t},\bm{W}_{t}-\bar{\bm{W}}_{t}\rangle,F_{t+1}(\widehat{\bm{W}}_{t})-F_{t+1}(\bar{\bm{W}}_{t})+\langle\bm{G}_{t},\bm{W}_{t}-\widehat{\bm{W}}_{t}\rangle\big{\}}_{+},$
 $\bar{\bm{W}}_{t}=\operatorname*{arg\,min}_{\bm{W}\in\mathcal{W}}F_{t+1}(\bm{W}),$
@@
 Then, for any $\bm{U}\in\mathcal{W}$, ODAFTRL achieves
-$\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\leq\sum_{t=1}^{T}{S_{t}(\bm{U})}+\left(\frac{B^{2}}{2\alpha}+1\right)\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+2\alpha b_{t}}\right),$
+$\sum_{t=1}^{T}{f_{t}(\bm{W}_{t})}-\sum_{t=1}^{T}{f_{t}(\bm{U})}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq\left(\frac{B^{2}}{2\alpha}+1\right)\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+2\alpha b_{t}}\right),$
 where
@@
 $b_{t}$ $=\operatorname{huber}(\|\bm{G}_{t-D:t}\|_{\mathrm{F}},\|\bm{G}_{t}\|_{\mathrm{F}}),\>\text{and}\>\operatorname{huber}(x,y)=\frac{1}{2}x^{2}-\frac{1}{2}(|x|-|y|)^{2}_{+}\leq\min\left\{\frac{1}{2}x^{2},|x||y|\right\}.$
-In the following, we let $\alpha=\max_{\bm{W}\in\mathcal{W}}\frac{\|\bm{W}\|_{\mathrm{F}}^{2}}{2}=\frac{B^{2}}{2}$ for simplicity. Then,
-$\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\leq\sum_{t=1}^{T}{S_{t}(\bm{U})}+2\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+B^{2}b_{t}}\right),$
+In the following, we let $\alpha=\frac{B^{2}}{2}$ for simplicity. Then, since $S_{t}$ is the convex function ,we have
+$\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}-\sum_{t=1}^{T}{S_{t}(\bm{U})}\leq\sum_{t=1}^{T}{\langle\bm{G}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq 2\left(2\max_{s\in[T]}a_{s-D:s-1}+\sqrt{\sum_{t=1}^{T}a_{t}^{2}+B^{2}b_{t}}\right).$
 E.2 Proof of Theorem 4.3
-We present Theorem 4.3 in a more detailed form and provide its proof.
+We present Theorem 4.3 in a more detailed form and provide its proof. In what follows, let
+$C_{y}=\max\left\{\max_{\bm{y}\in\operatorname{conv}(\mathcal{Y})}\|\bm{y}\|_{2},\max_{\bm{y},\bm{y}^{\prime}\in\operatorname{conv}(\mathcal{Y})}\|\bm{y}-\bm{y}^{\prime}\|_{2}\right\}$
+denote the maximum of the largest Euclidean norm of vectors in $\operatorname{conv}(\mathcal{Y})$ or the diameter of $\operatorname{conv}(\mathcal{Y})$.
 Theorem E.2 (Formal version of Theorem 4.3).
-Let $\alpha=\frac{B^{2}}{2}$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a delay of $D$ achieves
-$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BDL+\frac{2bB^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3}.$
+Let $\alpha=\frac{B^{2}}{2}$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a fixed delay of $D$ achieves
+$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BC_{x}C_{y}D+\frac{2bB^{2}}{a}+\min\left\{\frac{b(D+1)^{2}}{2},\frac{3}{2}\left(a^{-1}bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\right)^{1/3}\right\}.$
 Proof.
-From Lemma E.1 and the definition of $a_{t}$ and $b_{t}$, we have
-$\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}$ $\leq 2\left(2B\max_{s\in[T]}\sum_{i=s-D}^{s-1}\|\bm{G}_{i}\|_{\mathrm{F}}+\sqrt{\sum_{t=1}^{T}{\left(B^{2}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+B^{2}\|\bm{G}_{t-D:t}\|_{\mathrm{F}}\|\bm{G}_{t}\|_{\mathrm{F}}\right)}}\right)$
-$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right),$
-where we used the Cauchy–Schwarz inequality, $\|\bm{G}_{t}\|_{\mathrm{F}}\leq L$, $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$, and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$ in the last inequality.
-From this inequality and Assumption 4.1, we can evaluate surrogate regret as
+From the definition of $b_{t}$, it holds that
+$b_{t}$ $\leq\min\left\{\frac{1}{2}\|\bm{G}_{t-D:t}\|_{\mathrm{F}}^{2},\|\bm{G}_{t-D:t}\|_{\mathrm{F}}\|\bm{G}_{t}\|_{\mathrm{F}}\right\}$
+$\leq\min\left\{\frac{b(D+1)}{2}\sum_{s=t-D}^{t}S_{t}(\bm{W}_{t}),C_{x}C_{y}(D+1)\sqrt{bT\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}\right\},$
+where we used $\|\sum_{i=1}^{n}\bm{A}_{i}\|_{\mathrm{F}}^{2}\leq n\sum_{i=1}^{n}\|\bm{A}_{i}\|_{\mathrm{F}}^{2}$ for any matrix $\bm{A}_{i}$, the Cauchy–Schwarz inequality, $\|\bm{G}_{t}\|_{\mathrm{F}}\leq\|\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\bm{y}_{t}\|_{2}\|\bm{x}_{t}\|_{2}\leq C_{x}C_{y}$, and $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$.
+Combining this inequality with Lemma E.1 and the definition of $a_{t}$, we have
+$\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}$
+$\leq 4B\max_{s\in[T]}\sum_{i=s-D}^{s-1}\|\bm{G}_{i}\|_{\mathrm{F}}$
+$\qquad+2B\sqrt{\sum_{t=1}^{T}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+\min\left\{\frac{b(D+1)}{2}\sum_{t=1}^{T}\sum_{s=t-D}^{t}S_{t}(\bm{W}_{t}),C_{x}C_{y}(D+1)\sqrt{bT\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}\right\}}$
+$\leq 4BC_{x}C_{y}D+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
+$\qquad+2\min\left\{\sqrt{\frac{b(D+1)^{2}}{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}},\left(bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right\},$
+where we used $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$ and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$ in the last inequality.
+From this inequality and Assumption 4.1, we can the evaluate surrogate regret as
 $\mathbb{E}\left[\mathcal{R}_{T}\right]\leq\sum_{t=1}^{T}{\left((1-a)S_{t}(\bm{W}_{t})-S_{t}(\bm{U})\right)}$
-$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right)-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
-$\leq 4BDL+\frac{2bB^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3}.$
-We used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq({3}/{4})\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$ that hold for any $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$ in the last inequality.
+$\leq 4BC_{x}C_{y}D+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
+$+2\min\left\{\sqrt{\frac{b(D+1)^{2}}{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}},\left(bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right\}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
+$\leq 4BC_{x}C_{y}D+\frac{2bB^{2}}{a}+\min\left\{\frac{b(D+1)^{2}}{a},\frac{3}{2}\left(a^{-1}bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\right)^{1/3}\right\},$
+where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq({3}/{4})\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$, which hold for any $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$, in the last inequality.
 ∎
-E.3 Proof of Theorem 4.4
-We present Theorem 4.4 in a more detailed form and provide its proof.
+E.3 Proof of high probability bound
+We present the result of the high probability bound in a more detailed form and provide its proof.
 Theorem E.3.
-Let $\alpha=\frac{B^{2}}{2}$ and $\delta\in(0,1)$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a delay of $D$ achieves
-$\mathcal{R}_{T}\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3},$
+Let $\alpha=\frac{B^{2}}{2}$ and $\delta\in(0,1)$. Then, ODAFTRL with the AdaHedgeD update in online structured prediction with a fixed delay of $D$ achieves
+$\mathcal{R}_{T}\leq$ $4BC_{x}C_{y}D+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}$
+$+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}+\min\left\{\frac{b(D+1)^{2}}{a},\frac{3}{2}\left(a^{-1}bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\right)^{1/3}\right\},$
 with probability at least $1-\delta$.
@@
 Let $Z_{t}=L_{t}(\widehat{\bm{y}}_{t})-\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]$.
-Then, we have $|Z_{t}|\leq 1$ and $\mathbb{E}_{t}\left[{Z_{t}^{2}}\right]\leq(1-a)S_{t}(\bm{W}_{t})$ from Assumption 4.1.
+Then, we have $|Z_{t}|\leq 1$ and $\mathbb{E}_{t}\left[{Z_{t}^{2}}\right]\leq\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]\leq(1-a)S_{t}(\bm{W}_{t})$ from Assumption 4.1.
 Hence, from Lemma D.1, with probability at least $1-\delta$, the first term in (11) is upper bounded as
 $\sum_{t=1}^{T}{Z_{t}}\leq\sqrt{2(1-a)\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\log\frac{1}{\delta}}+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}.$ (12)
-From from Assumption 4.1 and Lemma E.1, the second term in (14) is also upper bounded as
+From Assumption 4.1 and Lemma E.1, the second term in (11) is also upper bounded as
 $\sum_{t=1}^{T}{(\mathbb{E}_{t}\left[{L_{t}(\widehat{\bm{y}}_{t})}\right]-S_{t}(\bm{U}))}$
-$\leq 2\left(2BDL+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\left(bB^{4}L^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right)-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})},$ (16)
-where we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, substituting (15) and (16) into (14) gives
-$\mathcal{R}_{T}$ $\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\left(\sqrt{2(1-a)\log\frac{1}{\delta}}+2\sqrt{bB^{2}}\right)\sqrt{\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
-$\quad+2\left(bB^{4}(D+1)^{2}L^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
-$\leq 4BDL+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}+\frac{3}{2}\left(a^{-1}bB^{4}L^{2}(D+1)^{2}T\right)^{1/3},$
+$\leq 4BC_{x}C_{y}D+\sqrt{bB^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
+$\quad+2\min\left\{\sqrt{\frac{b(D+1)^{2}}{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}},\left(bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right\}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})},$ (13)
+where we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. Therefore, substituting (12) and (E.3) into (11) gives
+$\mathcal{R}_{T}$ $\leq 4BC_{x}C_{y}D+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\left(\sqrt{2(1-a)\log\frac{1}{\delta}}+2\sqrt{bB^{2}}\right)\sqrt{\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}$
+$+\min\left\{\sqrt{2b(D+1)^{2}\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}},2\left(bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right)^{1/4}\right\}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
+$\leq 4BC_{x}C_{y}D+\frac{\sqrt{2}}{3}\log\frac{1}{\delta}+\frac{\left(\sqrt{(1-a)\log\frac{1}{\delta}}+\sqrt{2bB^{2}}\right)^{2}}{a}$
+$+\min\left\{\frac{b(D+1)^{2}}{a},\frac{3}{2}\left(a^{-1}bB^{4}C_{x}^{2}C_{y}^{2}(D+1)^{2}T\right)^{1/3}\right\},$
 where we used $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ and $c_{1}x-c_{2}x^{4}\leq({3}/{4})\left({c_{1}^{4}}/(4c_{2})\right)^{1/3}$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$ in the last inequality.
@@
 ∎
-Appendix F Omitted Details of Section 5
+Appendix F Details omitted from Section 5
 This section provides the omitted proofs of the theorems in Section 5.
 F.1 Common analysis
-We provide the analysis that is commonly used in the proofs of Theorems 5.1 and 5.2. We use ODAFTRL with the AdaHedgeD update in Section E.1 for $\mathsf{ALG}$. From Lemma E.1, it holds that
-$\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq 2\left(2\max_{t\in[T]}a_{t-D:t-1}+\sqrt{\sum_{t=1}^{T}{a_{t}^{2}}+B^{2}b_{t}}\right),$ (17)
+We provide the analysis that is commonly used in the proofs of Theorems 5.1 and 5.2.
+Although we use $\widehat{\bm{G}}_{t}$ as a notation for the estimator for convenience in this subsection, the same argument applies equally to $\widetilde{\bm{G}}_{t}$.
+We use ODAFTRL with the AdaHedgeD update in Section E.1 as $\mathsf{ALG}$.
+Here, we recall that $\mathbb{E}[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}]\leq\mathbb{E}[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}]$ from the convexity of $S_{t}$ and the unbiasness of $\widehat{\bm{G}}_{t}$.
+From Lemma E.1, it holds that
+$\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\leq 2\left(2\max_{t\in[T]}a_{t-D:t-1}+\sqrt{\sum_{t=1}^{T}{a_{t}^{2}}+B^{2}b_{t}}\right),$ (14)
 where
-$a_{t}=B\min\{\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}},\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\}\quad\mbox{and}\quad b_{t}\leq\min\left\{\frac{1}{2}\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}^{2},\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right\}.$
+$a_{t}=B\min\{\|\widehat{\bm{G}}_{t-D:t}\|_{\mathrm{F}},\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\}\quad\mbox{and}\quad b_{t}\leq\min\left\{\frac{1}{2}\|\widehat{\bm{G}}_{t-D:t}\|_{\mathrm{F}}^{2},\|\widehat{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\right\}.$
 By the definition of $a_{t}$, we have
-$\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]\leq B\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]\leq B\mathbb{E}\left[\sqrt{D\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}^{2}}\right]\leq B\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]},$ (18)
-from the Cauchy–Schwarz inequality and Jensen’s inequality. Hence, from (17) and (18), we have
-$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2\left(2\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{a_{t}^{2}}\right]}+B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
-$\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right),$ (19)
+$\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]\leq B\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}\right],$ (15)
+and thus
+$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2\left(2\mathbb{E}\left[\max_{t\in[T]}a_{t-D:t-1}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{a_{t}^{2}}\right]}+B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
+$\!\leq 2B\left(2\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right),$ (16)
 where we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$. The last term in the last inequality is further bounded as
-$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
-$=\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right],$ (20)
+$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t-D:t}\|_{\mathrm{F}}\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}}\right]\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
+$=\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}}\right],$ (17)
 where the second inequality follows from the triangle inequality and the equality follows from the law of total expectation.
@@
 The algorithm in Section 5.1 achieves
-$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq\frac{4bB}{a}\left(D^{1/4}+D^{-1/4}\right)^{2}\sqrt{KT}+2BL\sqrt{DT}+B\sqrt{DKT}=O(D\sqrt{KT}).$
+$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BC_{x}C_{y}D+\left(\frac{4bB}{a}+1\right)\sqrt{KT}+2BC_{x}C_{y}\sqrt{DT}=O(\sqrt{(K+D)T}).$
 Proof.
 First, we will upper bound $\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$. From (F.1), we have
-$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
-$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+L\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
-$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DL^{2}T,$ (21)
-where the second and third inequality follow from the inequality $\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]=\|\bm{G}_{t}\|_{\mathrm{F}}\leq L$.
-Hence, from (F.1) and (F.2), it holds that
-$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DL^{2}T}}\right)$
-$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2BL\sqrt{DT}$
-$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BL\sqrt{DT},$ (22)
+$\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
+$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+C_{x}C_{y}\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
+$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+C_{x}^{2}C_{y}^{2}DT,$ (18)
+where the second and third inequality follow from the inequality $\mathbb{E}_{t}\left[{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}}\right]=\|\bm{G}_{t}\|_{\mathrm{F}}\leq C_{x}C_{y}$.
+Hence, from $\mathbb{E}_{t}\left[{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}}\right]\leq C_{x}C_{y}$, (F.1) and (F.2), it holds that
+$\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$ $\leq 2B\left(2C_{x}C_{y}D+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+C_{x}^{2}C_{y}^{2}DT}}\right)$
+$\leq 4BC_{x}C_{y}D+4B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2BC_{x}C_{y}\sqrt{DT}$
+$\leq 4BC_{x}C_{y}D+4B\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BC_{x}C_{y}\sqrt{DT},$ (19)
 where in the second inequality we used the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$ and in the last inequality we used
-$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}).$
+$\mathbb{E}_{t}\left[{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]=\frac{\|\bm{G}_{t}\|_{\mathrm{F}}^{2}}{p_{t}(\bm{y}_{t})}\leq\frac{K}{q}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq\frac{bK}{q}S_{t}(\bm{W}_{t}).$
 Therefore, combining all the above arguments yields
 $\mathbb{E}\left[\mathcal{R}_{T}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
-$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
-$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BL\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
-$\leq\frac{4bB^{2}}{a}\left(\sqrt{D}+1\right)^{2}\frac{bK}{q}+2BL\sqrt{DT}+qT,$
-where the first inequality follows from Assumption 3.2, the second inequality follows from the convexity of $S_{t}$ and the unbiasedness of $\widetilde{\bm{G}}_{t}$, and the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
-Finally, by substituting $q=B\sqrt{DK/T}$, we obtain
-$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq\frac{4bB}{a}\left(D^{1/4}+D^{-1/4}\right)^{2}\sqrt{KT}+2BL\sqrt{DT}+B\sqrt{DKT},$
+$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widehat{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
+$\leq 4BC_{x}C_{y}D+4B\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+2BC_{x}C_{y}\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
+$\leq 4BC_{x}C_{y}D+\frac{4bB^{2}}{a}\frac{bK}{q}+2BC_{x}C_{y}\sqrt{DT}+qT,$
+where the first inequality follows from Assumption 3.2, the second inequality follows from the convexity of $S_{t}$ and the unbiasedness of $\widehat{\bm{G}}_{t}$, and the last inequality follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
+Finally, from $q=B\sqrt{K/T}$, we obtain
+$\mathbb{E}\left[\mathcal{R}_{T}\right]\leq 4BC_{x}C_{y}D+\left(\frac{4bB}{a}+1\right)\sqrt{KT}+2BC_{x}C_{y}\sqrt{DT},$
 which is the desired bound. ∎
@@
 The algorithm in Section 5.2 achieves
-$\mathbb{E}[\mathcal{R}_{T}]=\frac{8bB^{2}}{a}\Big{(}\sqrt{D}+1\Big{)}^{2}+2B(L+C_{x}C_{y})\sqrt{DT}+O\left(B^{2/3}D^{1/3}T^{2/3}\right).$
+$\mathbb{E}[\mathcal{R}_{T}]=4BC_{x}C_{y}(2D+\sqrt{DT})+\frac{8bB^{2}}{a}+O\left(\omega^{1/3}D^{1/3}T^{2/3}\right).$
 Proof.
 First, we will derive an upper bound of $\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$. We first observe that
-$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]$ $=\mathbb{E}_{t}\left[{\|(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}\|_{\mathrm{F}}}\right]\leq\mathbb{E}_{t}\left[{\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}}\right]\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\mathbb{E}_{t}\left[{\|\bm{y}_{t}\|_{2}+\|\widetilde{\bm{y}}_{t}\|_{2}}\right]$
+$\mathbb{E}_{t}\left[{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}}\right]$ $=\mathbb{E}_{t}\left[{\|(\widehat{\bm{y}}_{\Omega}(\bm{\theta}_{t})-\widetilde{\bm{y}}_{t})\bm{x}_{t}^{\top}\|_{\mathrm{F}}}\right]$
+$\leq\mathbb{E}_{t}\left[{\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\|\bm{y}_{t}-\widetilde{\bm{y}}_{t}\|_{2}}\right]\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}\mathbb{E}_{t}\left[{\|\bm{y}_{t}\|_{2}+\|\widetilde{\bm{y}}_{t}\|_{2}}\right]$
 $\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+C_{x}\mathbb{E}_{t}\left[{\sqrt{\operatorname*{tr}\left(\widetilde{\bm{y}}_{t}\widetilde{\bm{y}}_{t}^{\top}\right)}}\right]\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+C_{x}{\sqrt{\mathbb{E}_{t}\left[{\operatorname*{tr}\left(\widetilde{\bm{y}}_{t}\widetilde{\bm{y}}_{t}^{\top}\right)}\right]}}$
-$\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\leq L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q},$
-where the first inequality follows from $C_{x}=\mathrm{diam}(\mathcal{X})$, the third inequality follows from $C_{y}=\mathrm{diam}(\mathcal{Y})$, the fourth inequality follows from Jensen’s inequality, and the fifth inequality follows from Lemma D.5.
+$\leq\|\bm{G}_{t}\|_{\mathrm{F}}+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\leq 2C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q},$ (20)
+where the first inequality follows from $C_{x}\geq\|\bm{x}_{t}\|_{2}$, the third inequality follows from $C_{y}\geq\|\bm{y}_{t}\|_{2}$, the fourth inequality follows from Jensen’s inequality, and the fifth inequality follows from Lemma D.6.
 Thus, combining (F.1) with the last inequality, we have
 $\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\mathbb{E}\left[\sum_{t=1}^{T}{\mathbb{E}_{t}\left[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}\right]\sum_{s=t-D}^{t-1}\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}}\right]$
-$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\left(L+C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
-$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DT\left(L+C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)^{2}.$ (23)
-Hence, from (F.1) and (23), we have
+$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+\left(2C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)\mathbb{E}\left[\sum_{t=1}^{T}{\sum_{s=t-D}^{t-1}\mathbb{E}_{s}\left[\|\widetilde{\bm{G}}_{s}\|_{\mathrm{F}}\right]}\right]$
+$\leq\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+DT\left(2C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)^{2}.$ (21)
+Hence, from (F.1), (F.3),and (21), we have
 $\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]$
-$\leq 2B\left(2\sqrt{D\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
-$\leq 4B\left(\sqrt{D}+1\right)\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT}$
-$\leq 4B\left(\sqrt{D}+1\right)\sqrt{2\sum_{t=1}^{T}{\left(bS_{t}(\bm{W}_{t})+\frac{C_{x}^{2}\omega}{q}\right)}}+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT},$ (24)
-where the first inequality follows from (F.1), the second inequality follows from (23) and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$, and the last inequality follows from Lemma D.6.
+$\leq 2B\left(2\mathbb{E}\left[\max_{t\in[T]}\sum_{s=t-D}^{t-1}\|\widehat{\bm{G}}_{s}\|_{\mathrm{F}}\right]+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
+$\leq 2B\left(4C_{x}C_{y}D+2C_{x}D\sqrt{\omega/q}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{b_{t}}\right]}\right)$
+$\leq 8BC_{x}C_{y}D+4BC_{x}D\sqrt{\omega/q}+4B\sqrt{\mathbb{E}\left[\sum_{t=1}^{T}{\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]}+2B\left(2C_{x}C_{y}+C_{x}\sqrt{\omega/q}\right)\sqrt{DT}$
+$\leq 8BC_{x}C_{y}D+4BC_{x}D\sqrt{\omega/q}$
+$\qquad+4B\sqrt{2\sum_{t=1}^{T}{\left(bS_{t}(\bm{W}_{t})+\frac{C_{x}^{2}\omega}{q}\right)}}+2B\left(2C_{x}C_{y}+C_{x}\sqrt{\omega/q}\right)\sqrt{DT},$ (22)
+where the first inequality follows from (F.1), the second inequality follows from (F.3), the third inequality follows from (21) and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$, and the last inequality follows from Lemma D.7.
 Therefore, combining all the above arguments yields
@@
 $\leq\mathbb{E}\left[\sum_{t=1}^{T}{\langle\widetilde{\bm{G}}_{t},\bm{W}_{t}-\bm{U}\rangle}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
-$\leq 4B\left(\sqrt{D}+1\right)\left(\sqrt{2b\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}}+\sqrt{2C_{x}^{2}\omega T/q}\right)$
-$\qquad+2B\left(L+C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
-$\leq\frac{8bB^{2}}{a}\Big{(}\sqrt{D}+1\Big{)}^{2}+2B(L+C_{x}C_{y})\sqrt{DT}+2BC_{x}\big{(}(2\sqrt{2}+1)\sqrt{D}+2\sqrt{2}\big{)}\sqrt{\omega T/q}+qT,$
+$\leq 8BC_{x}C_{y}D+4B\left(\sqrt{2b\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+C_{x}\sqrt{2\omega T/q}\right)+4BC_{x}D\sqrt{\omega/q}$
+$\qquad+2B\left(2C_{x}C_{y}+C_{x}\sqrt{\omega/q}\right)\sqrt{DT}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
+$\leq 4BC_{x}C_{y}(2D+\sqrt{DT})+\frac{8bB^{2}}{a}+4BC_{x}D\sqrt{\omega/q}+2BC_{x}\big{(}\sqrt{D}+2\sqrt{2}\big{)}\sqrt{\omega T/q}+qT,$
 where the first inequality follows from Assumption 3.2, the third inequality follows from (F.3) and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$, and the last inequality follows from the definition of $\varepsilon$ and $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
@@
 ∎
-Appendix G Additional Experiments
-(a) $n^{\prime}=2,p=0.0$
-(b) $n^{\prime}=2,p=0.1$
-(c) $n^{\prime}=4,p=0.0$
-(d) $n^{\prime}=4,p=0.1$
-Figure 2: Results of the synthetic experiments in multiclass classification with bandit feedback.
-In all figures, the horizontal axis represents the number of classes $K$ and the vertical axis represents the cumulative prediction loss.
-This section provides the results of numerical experiments for online multiclass classification with bandit feedback on synthetic data, which could not be included in the main text due to space constraints.
-G.1 Setup
-Here, we describe the experimental setup.
-As Section 6, we compare three algorithms: Gaptron (Van der Hoeven, 2020) with logistic loss, Gappletron (Van der Hoeven et al., 2021) with hinge loss, and our algorithm in Section 5.1.
+Appendix G Variable delay
+This section provides the algorithms and analyses under the variable-delay setting, which is a natural extension of the fixed-delay setting.
+As the notation for the variable-delay setting, let $\tau_{t}$ denote the delay time of the feedback received at time $t$, and define $\tau_{\ast}=\max_{t}\tau_{t}$.
+Under this setting, by leveraging the Single-instance Online Learning In Delayed environments (SOLID) [23], we achieve a surrogate regret bound of $O(\sqrt{\tau_{1:T}}+\tau_{\ast})$ in the full-information setting (Theorem G.2), and bounds of $O(\sqrt{KT}+\sqrt{\tau_{1:T}}+T^{2/3}+\tau_{\ast})$ (Theorem G.3) and $O(T^{1/6}\sqrt{\tau_{1:T}}+\tau_{\ast})$ (Theorem G.4) in the bandit setting.
+G.1 Single-instance Online Learning In Delayed environments (SOLID)
+We provide a detail of SOLID algorithm used for updating $\bm{W}_{t}$ under the variable-delay setting.
+Algorithm 3 Single-instance Online Learning In Delayed environments (SOLID)
+1:BASE, the first prediction $\bm{W}$ of BASE
+2:for time step $t=1,2,\dots,T$ do
+3:     Set $\bm{W}_{t}\leftarrow\bm{W}$ as the prediction for the current time step.
+4:     Reieve the feedbacks $H_{t}$ that arrive at the end of time step $t$.
+5:     for all feedback in $H_{t}$ do
+6:         Update BASE with feedback
+7:         $\bm{W}\leftarrow$ the next prediction of BASE
+Consider any deterministic non-delayed online learning algorithm (call it BASE).
+SOLID is an algorithm that, regardless of the original arrival time of the feedback, provides the feedback to BASE in the order in which it is observed, and makes predictions based on the outputs of BASE (Algorithm 3).
+Below, let $\rho(t)$ denote the time step of the $t$th feedback from SOLID to BASE for any $t\in[T]$, as in [23].
+When we use OGD as BASE, SOLID achieves the following bound:
+Lemma G.1 ([23, Theorem 5]).
+Let BASE OGD with learning rate
+$\widetilde{\eta}_{t}=\sqrt{2}R\left(\sqrt{\sum_{s=1}^{t}(\|\bm{G}_{\rho(s)}\|_{\mathrm{F}}^{2}+2\|\bm{G}_{\rho(s)}\|_{\mathrm{F}}\sum_{i=t-\widetilde{\tau_{t}}}^{t-1}\|\bm{G}_{\rho(i)}\|_{\mathrm{F}})}+C_{x}^{2}C_{y}^{2}(\tau_{\ast}^{2}+\tau_{\ast})\right)^{-1},$
+where $R>0$ satisfies $\widetilde{\eta}_{T}\sum_{t=1}^{T}{\|\bm{U}-\bm{W}_{t}\|_{\mathrm{F}}^{2}}\leq 4R^{2}$. Then, SOLID achieves
+$\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}$
+$\leq 2\sqrt{2}R\sqrt{\sum_{t=1}^{T}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2\sum_{t=1}^{T}\|\bm{G}_{\rho(t)}\|_{\mathrm{F}}\sum_{s=t+1}^{T}\|\bm{G}_{\rho(s)}\|_{\mathrm{F}}\mathds{1}\{s-\widetilde{\tau}_{s}\leq t\}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}.$
+G.2 Full-inforamtion
+Here, we provide the algorithm for the variable-delay full-information setting. This subsection assumes Assumption 4.1, as in Section 4.
+Algorithm
+We use SOLID with OGD as $\mathsf{ALG}$ for updating $\bm{W}_{t}$.
+Regret bound and analysis
+The above algorithm achieves the following bound:
+Theorem G.2.
+SOLID with OGD update in online structured prediction with a delay of $\tau_{t}$ achieves
+$\mathbb{E}[\mathcal{R}_{T}]\leq\frac{2bR^{2}}{a}+4C_{x}C_{y}R\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}=O\left(\sqrt{\tau_{1:T}}+\tau_{\ast}\right).$
+Proof.
+Using SOLID with OGD (Lemma G.1), we have
+$\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}$
+$\leq 2\sqrt{2}R\sqrt{\sum_{t=1}^{T}\|\bm{G}_{t}\|_{\mathrm{F}}^{2}+2\sum_{t=1}^{T}\|\bm{G}_{\rho(t)}\|_{\mathrm{F}}\sum_{s=t+1}^{T}\|\bm{G}_{\rho(s)}\|_{\mathrm{F}}\mathds{1}\{s-\widetilde{\tau}_{s}\leq t\}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}$
+$\leq 2\sqrt{2}R\sqrt{b\sum_{t=1}^{T}S_{t}(\bm{W}_{t})}+4C_{x}C_{y}R\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})},$ (23)
+where we used $\|\bm{G}_{t}\|_{\mathrm{F}}^{2}\leq bS_{t}(\bm{W}_{t})$, $\|\bm{G}_{t}\|_{\mathrm{F}}\leq C_{x}C_{y}$, $\sum_{s=t+1}^{T}\mathds{1}\{s-\widetilde{\tau}_{s}\leq t\}=\widetilde{\tau}_{s}$, $\sum_{t=1}^{T}{\widetilde{\tau}_{t}}=\sum_{t=1}^{T}{\tau_{t}}$, and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$ in the last ineqaulity.
+Therefore, from this inequality, it holds that
+$\mathbb{E}[\mathcal{R}_{T}]$ $\leq\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
+$\leq 2\sqrt{2}R\sqrt{b\sum_{t=1}^{T}S_{t}(\bm{W}_{t})}+4C_{x}C_{y}R\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}-a\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}$
+$\leq\frac{2bR^{2}}{a}+4C_{x}C_{y}R\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})},$
+where the first inequality follows from Assumption 4.1, the second ineqaulity follows from (G.2), and the last ineqaulity follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
+This is the desired bound.
+∎
+This result is superior to the algorithm designed for the fixed-delay feedback in that it can handle variable-delay feedback.
+When $D\geq\tau_{\ast}$ is known and small, we may also use the algorithm developed for fixed-delay feedback to achieve the $O(D^{2})$ bound.
+G.3 Bandit feedback
+We provide algorithms for the variable-delay bandit setting.
+This subsection assumes Assumption 3.2, as in Section 3.
+Then, by using the inverse-weighted gradient estimator and the pseudo-inverse matrix estimator as gradient estimators, we can achieve surrogate regret upper bounds of $O(\sqrt{KT}+\sqrt{\tau_{1:T}}+\tau_{\ast})$ and $O(T^{1/6}\sqrt{\tau_{1:T}}+T^{2/3}+\tau_{\ast})$, respectively.
+Below, we provide details of these results.
+G.3.1 Algorithm based on inverse-weighted gradient estimator with $O(\sqrt{KT}+\sqrt{\tau_{1:T}}+\tau_{\ast})$ regret
+Here, we introduce an algorithm with a surrogate regret upper bound of $O(\sqrt{KT}+\sqrt{\tau_{1:T}}+\tau_{\ast})$.
+Algorithm
+We use RDUE with $q=R\sqrt{K/T}$ for decoding (assuming $T\geq R^{2}K$), the gradient estimator $\widehat{\bm{G}}_{t}$ as in Section 3.3, and SOLID with OGD as $\mathsf{ALG}$.
+Regret bound and analysis
+The above algorithm achieves the following surrogate regret bound:
+Theorem G.3.
+The above algorithm achieves
+$\mathbb{E}[\mathcal{R}_{T}]\leq\left(\frac{2b}{a}+1\right)R\sqrt{KT}+4C_{x}C_{y}R\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}=O(\sqrt{KT}+\sqrt{\tau_{1:T}}+\tau_{\ast}).$
+This result is an extension of the bound under the fixed-delay setting. In particular, if $\tau_{t}=D$ for any $t$, we obtain $\mathbb{E}[\mathcal{R}_{T}]=O(\sqrt{(K+D)T})$.
+Proof.
+First, we recall that $\mathbb{E}_{t}[\|\bm{G}_{t}\|_{\mathrm{F}}]\leq C_{x}C_{y}$ and $\mathbb{E}_{t}[\|\bm{G}_{t}\|_{\mathrm{F}}^{2}]\leq bKS_{t}(\bm{W}_{t})/q$ from the proof of Theorem 5.1.
+Using SOLID with OGD (Lemma G.1) and the subadditivity of $x\mapsto\sqrt{x}$ for $x\geq 0$, we have
+$\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]$
+$\leq 2\sqrt{2}R\mathbb{E}\left[\sqrt{\sum_{t=1}^{T}\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}+2\sum_{t=1}^{T}\sum_{s=t+1}^{T}\|\widehat{\bm{G}}_{\rho(t)}\|_{\mathrm{F}}\|\widehat{\bm{G}}_{\rho(s)}\|_{\mathrm{F}}\mathds{1}\{s-\widehat{\tau}_{s}\leq t\}}\right]$
+$\qquad+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}$
+$\leq 2\sqrt{2}R\mathbb{E}\left[\sqrt{\sum_{t=1}^{T}\|\widehat{\bm{G}}_{t}\|_{\mathrm{F}}^{2}}\right]+4R\mathbb{E}\left[\sqrt{\sum_{t=1}^{T}\sum_{s=t+1}^{T}\|\widehat{\bm{G}}_{\rho(t)}\|_{\mathrm{F}}\|\widehat{\bm{G}}_{\rho(s)}\|_{\mathrm{F}}\mathds{1}\{s-\widehat{\tau}_{s}\leq t\}}\right]$
+$\qquad+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}.$ (24)
+The second term is bounded as
+$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{s=t+1}^{T}\|\widehat{\bm{G}}_{\rho(t)}\|_{\mathrm{F}}\|\widehat{\bm{G}}_{\rho(s)}\|_{\mathrm{F}}\mathds{1}\{s-\widehat{\tau}_{s}\leq t\}\right]$
+$\leq\mathbb{E}\left[\sum_{t=1}^{T}\sum_{s=t+1}^{T}\mathbb{E}_{\rho_{\max}}\left[\|\widehat{\bm{G}}_{\rho_{\max}}\|_{\mathrm{F}}\right]\|\widehat{\bm{G}}_{\rho_{\min}}\|_{\mathrm{F}}\mathds{1}\{s-\widehat{\tau}_{s}\leq t\}\right]$
+$\leq C_{x}C_{y}\mathbb{E}\left[\sum_{t=1}^{T}\sum_{s=t+1}^{T}\mathbb{E}_{\rho_{\min}}\left[\|\widehat{\bm{G}}_{\rho_{\min}}\|\right]_{\mathrm{F}}\mathds{1}\{s-\widetilde{\tau}_{s}\leq t\}\right]\leq C_{x}^{2}C_{y}^{2}\sum_{t=1}^{T}\tau_{t},$ (25)
+where we assumed $\rho_{\max}=\max\{\rho(t),\rho(s)\}$ and $\rho_{\min}=\min\{\rho(t),\rho(s)\}$, used the tower property in the first and second ineqaulities, and used $\mathbb{E}_{t}[\|\bm{G}_{t}\|_{\mathrm{F}}]\leq C_{x}C_{y}$ in the second and last ineqaulities.
+Hence, it holds that
+$\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]$
+$\leq 2\sqrt{2}R\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+4C_{x}C_{y}R\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})},$ (26)
+where the ineqaulity follows from $\mathbb{E}_{t}[\|\bm{G}_{t}\|_{\mathrm{F}}^{2}]\leq bKS_{t}(\bm{W}_{t})/q$ and (G.3.1). Therefore, combining all the above arguments yields
+$\mathbb{E}[\mathcal{R}_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
+$\leq 2\sqrt{2}R\sqrt{\frac{bK}{q}\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]}+4C_{x}C_{y}R\sqrt{\tau_{1:T}}$
+$\qquad+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
+$\leq\frac{2bR^{2}K}{aq}+4C_{x}C_{y}R\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}+qT,$
+where the first ineqaulity follows from Assumption 3.2, the second ineqaulity follows from (G.3.1), and the last ineaaulity follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
+Finally, choosing $q=R\sqrt{K/T}$ gives the desired bound.
+∎
+G.3.2 Algorithm based on pseudo-inverse matrix estimator with $O(T^{1/6}\sqrt{\tau_{1:T}}+T^{2/3}+\tau_{\ast})$ regret
+Here, we provide an algorithm that achieves a surrogate regret upper bound of $O(T^{1/6}\sqrt{\tau_{1:T}}+T^{2/3}+\tau_{\ast})$. This subsection assumes Assumption 3.6 in addition to Assumption 3.2.
+Algorithm
+We use RDUE with $q=\left(\frac{\omega R^{2}C_{x}^{2}}{T}\right)^{1/3}$ for decoding (assuming $T\geq\omega R^{2}C_{x}^{2}$), the gradient estimator $\widetilde{\bm{G}}_{t}$ as in Section 3.4, and SOLID with OGD as $\mathsf{ALG}$.
+Regret bound and analysis
+The algorithm described above achieves the following surrogate regret bound:
+Theorem G.4.
+The above algorithm achieves
+$\mathbb{E}[\mathcal{R}_{T}]$ $\leq\frac{4bR^{2}K}{a}+8C_{x}C_{y}R\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}+O\left(\omega^{1/3}R^{2/3}T^{1/6}\sqrt{\tau_{1:T}}+T^{2/3}\right)$
+$=O(T^{1/6}\sqrt{\tau_{1:T}}+T^{2/3}+\tau_{\ast}).$
+Proof.
+First, we recall that $\mathbb{E}_{t}[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}]\leq 2C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}$ and $\mathbb{E}_{t}[\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}]\leq 2bKS_{t}(\bm{W}_{t})+\frac{2C_{x}^{2}\omega}{q}$ hold from (F.3) and Lemma D.7, respectively.
+Following the same steps as in the proof of Theorem G.3, we obtain
+$\mathbb{E}\left[\sum_{t=1}^{T}\sum_{s=t+1}^{T}\|\widetilde{\bm{G}}_{\rho(t)}\|_{\mathrm{F}}\|\widetilde{\bm{G}}_{\rho(s)}\|_{\mathrm{F}}\mathds{1}\{s-\widetilde{\tau}_{s}\leq t\}\right]\leq\left(2C_{x}C_{y}+\sqrt{C_{x}^{2}\omega/q}\right)^{2}\sum_{t=1}^{T}{\tau_{t}}.$ (27)
+Therefore, by using these inequalities and (G.3.1), we get
+$\mathbb{E}[\mathcal{R}_{T}]$ $\leq\mathbb{E}\left[\sum_{t=1}^{T}{(S_{t}(\bm{W}_{t})-S_{t}(\bm{U}))}\right]-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
+$\leq 4R\mathbb{E}\left[\sqrt{bK\mathbb{E}\left[\sum_{t=1}^{T}S_{t}(\bm{W}_{t})\right]+\frac{C_{x}^{2}\omega}{q}}\right]+4R\left(2C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)\sqrt{\tau_{1:T}}$
+$\qquad+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}-a\mathbb{E}\left[\sum_{t=1}^{T}{S_{t}(\bm{W}_{t})}\right]+qT$
+$\leq\frac{4bR^{2}K}{a}+4R\sqrt{\frac{C_{x}^{2}\omega}{q}}+4R\left(2C_{x}C_{y}+\sqrt{\frac{C_{x}^{2}\omega}{q}}\right)\sqrt{\tau_{1:T}}+C_{x}C_{y}R\sqrt{2(\tau_{\ast}^{2}+\tau_{\ast})}+qT,$
+where the first inequality follows from Assumption 3.2, the second ineqaulity follows from Lemma D.7, (G.3.1), and (27), the last ineqaulity follows from $c_{1}\sqrt{x}-c_{2}x\leq{c_{1}^{2}}/(4c_{2})$ for $x\geq 0$, $c_{1}\geq 0$, and $c_{2}>0$.
+Finally, by substituting $q=\left(\frac{\omega R^{2}C_{x}^{2}}{T}\right)^{1/3}$, we can obtain the desired bound.
+∎
+Appendix H Numerical experiments
+This section presents the results of numerical experiments for online multiclass classification and multilabel classification under bandit feedback on MNIST and synthetic data.
 All experiments were run on a system with 16GB of RAM, Apple M3 CPU, and in Python 3.11.7 on a macOS Sonoma 14.6.1.
+Figure 1: A box plot of error rates of the MNIST experiment for multiclass classification with bandit feedback.
+H.1 Multiclass classification
+Setup
+We describe the experimental setup.
+We compare four algorithms: Gaptron [38] with logistic loss, Gappletron [39] with logistic loss and hinge loss, and our algorithm in Section 5.1.
+Theoretically, these methods have their own advantages: ours enjoys a surrogate regret bound of $O(\sqrt{KT})$, which is better than $O(K\sqrt{T})$ of the others; however, Gaptron/Gappletron can work with a broader class of surrogate losses.
+This section aims to compare those methods from the empirical perspective.
+Details of algorithms
+As the algorithm $\mathsf{ALG}$ to update the linear estimator, we employ the OGD in Section 3.2.
+Following [39], we use the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\left(10^{-8}+\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}\right)}}$ and no projection is conducted in OGD.
+Here, the addition of $10^{-8}$ to the denominator is to prevent division by zero.
+Although $B=\mathrm{diam}(\mathcal{W})$ is unknown, we fixed $B=10$ for the experiments, regardless of whether this value represents the actual diameter.
+All other parameters are set according to theoretical values.
+Under these parameter settings, we repeat experiments $20$ times.
+H.1.1 MNIST
+We use the MNIST dataset [26], a widely used dataset of digit images.
+Result
+The box plot in Figure 1 summarizes the resulting misclassification rates.
+It indicates that our method achieves the lowest misclassification rate.
+Even though our method is not specialized for multiclass classification, it outperforms existing algorithms designed for multiclass classification.
+H.1.2 Synthetic data
+We also run experiments on synthetic data to facilitate comparisons across different values of $K$.
 Data generation
@@
 The synthetic data was generated by using the same method as van der Hoeven et al.
-(2021).
+[39].
 The input vector consists of a binary vector with entries of 00 and $1$, and is composed of two parts.
 The first part corresponds to a unique feature vector associated with the label, and the second part is randomly selected and unrelated to the label.
-Specifically, the data are generated as follows: Let $n^{\prime}$, $K\in\mathbb{N}$, $r\in[0,1]$, and $T>0$.
-We generate $K$ unique feature vectors of length $10n^{\prime}$ using the following as follows.
+Specifically, the data is generated as follows: Let $n^{\prime}$, $K\in\mathbb{N}$, $r\in[0,1]$, and $T>0$.
+We generate $K$ unique feature vectors of length $10n^{\prime}$ as follows.
 We first randomly select an integer $s$ uniformly from the range $n^{\prime}$ to $5n^{\prime}$.
@@
 Additionally, as noise, with probability $r$, the corresponding class is replaced with a randomly selected class instead of the original one.
-The length of the generated input vector $n$ is $40n^{\prime}$.
+The length, $n$, of the generated input vector is $40n^{\prime}$.
 These input vectors are generated for $T$.
 Based on the above procedure, we create data for $n^{\prime}\in\{2,4\}$ and $r\in\{0.0,0.1\}$.
+Results
+The results on synthetic data are provided in Figure 2.
+Our algorithm achieves comparable or better performance than the existing methods for datasets with $K\leq 24$.
+By contrast, when $K=48,96$, the cumulative losses of our algorithm are larger compared to the other two algorithms, Gaptron with the hinge loss and Gappletron with the logistic loss.
+Note that these experimental results do not disagree with the theoretical claims since, for larger values of $K$, the upper bound on the cumulative 0-1 loss of Gappletron can be better than ours due to the difference in the surrogate loss functions (see the discussion in Appendix C for details).
+Nevertheless, in this experiment, our method for structured prediction does not demonstrate its full ability, as the setup is favorable for the existing algorithms for multiclass classification.
+(a) $n^{\prime}=2,r=0.0$
+(b) $n^{\prime}=2,r=0.1$
+(c) $n^{\prime}=4,r=0.0$
+(d) $n^{\prime}=4,r=0.1$
+Figure 2: Results of the synthetic experiments in multiclass classification with bandit feedback.
+In all figures, the horizontal axis represents the number of classes $K$, and the vertical axis represents the cumulative target loss.
+Figure 3: Results of the synsetic experiments in multilabel classification with bandit feedback.
+The horizontal axis shows the number of labels, and the vertical axis indicates the cumulative target loss.
+H.2 Multilabel classification
+In the results presented in Section 3, the algorithm based on the pseudo-inverse matrix estimator achieves a tighter upper bound in terms of its dependence on $K$ compared to the one based on the inverse-weighted gradient estimator.
+Here, to examine whether this theoretical result can also be observed empirically, we conduct experiments on multilabel classification with a fixed number of correct labels.
+Setup
+Here, we compare two algorithms: the algorithm based on the inverse-weighted gradient estimator (Section 3.3) and that based on the pseudo-inverse matrix estimator (Section 3.4).
+Data generation
+We generate synthetic data by using the multilabel classification data generation function provided in scikit-learn [32].
+The data are generated as follows: Let $n,d,m,T\in\mathbb{N}$.
+Using the make_multilabel_classification function in scikit-learn, we generate $T$ multilabel classification samples with feature dimension $n$, label dimension $d$, and an average number, $m$, of correct labels per sample.
+After that, we extract only those with exactly $m$ correct labels from the generated data.
+Finally, we repeat this process until we obtain $T$ samples with $m$ correct labels.
+Based on the above procedure, we create data for $n=50$, $d\in\{10,12,16,20,24\}$, $m=5$, $T=10^{4}$.
 Details of algorithms
-As the algorithm $\mathsf{ALG}$ to update the linear estimator, we employ the adaptive OGD in Section 3.2.
-Following Van der Hoeven et al.
-(2021), we use the learning rate of $\eta_{t}=\frac{B}{\sqrt{2\left(10^{-8}+\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}\right)}}$ and no orthogonal projection in conducted in OGD.
-Since $B=\mathrm{diam}(\mathcal{W})$ is unknown, we fixed $B=10$ for the experiments, regardless of whether this value represents the actual diameter.
+As the algorithm $\mathsf{ALG}$ to update the linear estimator, we employ the OGD in Section 3.2 with learning rate $\eta_{t}=\frac{B}{\sqrt{2\left(10^{-8}+\sum_{i=1}^{t}\|\widetilde{\bm{G}}_{t}\|_{\mathrm{F}}^{2}\right)}}$ and orthogonal projection onto a ball of diameter 50.
+Here, the addition of $10^{-8}$ to the denominator is to prevent division by zero.
+We fixed $B=50$ for the experiments.
 All other parameters were set according to theoretical values.
-Under these parameter settings, we repeated experiments for $20$ times.
-G.2 Results
-The results on synthetic data are provided in Figure 2.
-Our algorithm achieves comparable or better results than the comparison methods for datasets when $K\leq 24$.
-By contrast, when $K=48,96$, the cumulative losses of our algorithm are larger compared to the other two algorithms, Gaptron with the hinge loss and Gappletron with the logistic loss.
-Note that these experimental results are consistent with existing theoretical results since for larger values of $K$ the upper bound on the cumulative loss of Gappletron is better than our algorithm due to differences in the surrogate loss function (see the discussion in Appendix C for details).
-Nevertheless, in this experiment, the advantages of structural prediction algorithms are not fully exploited, making the setup more favorable for the algorithms in previous studies.
-It is also worth noting that by using the same decoding function as theirs, our approach can achieve the same order of the cumulative losses in online multiclass classification with bandit feedback.
+Under these parameter settings, we repeated experiments $10$ times.
+Results
+We provide the results in Figure 3.
+When $d$ is small, the loss incurred by the algorithm based on the inverse-weighted gradient estimator is relatively low, whereas when $d$ is large, the algorithm based on the pseudo-inverse matrix estimator yields a smaller loss.
+The superiority of the algorithm based on the inverse-weighted gradient estimator for small $d$ is consistent with the theoretical result that this algorithm has a better dependence on $T$.
+Similarly, the better performance of the algorithm based on the pseudo-inverse matrix estimator for large $d$ matches the theoretical result that this algorithm does not depend explicitly on $K$.
+Thus, these experimental findings provide empirical validation of our theoretical results.